A Complete Bibliography of the Journal of Econometrics (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

ARMA(p, q) [ZWG91]. C_{L} [And91]. C_{p} [KS90]. F [AS96]. I(2) [HS98a]. K [WY97, Kin91]. MA(q) [RM92]. p [AKS92]. Q [BBDS92, Run97a, Run97b]. R [CW97, Sne99]. r − 1 [Sne99]. R^{2} [Lai93]. t [KM97]. VEC(1) [BOYF94].

-class [Kin91]. -cones [WY97]. -squared [CW97]. -statistic [Run97a, Run97b]. -tests [AS96]. -values [AKS92, KM97].

1-3 [BCL93]. 11 [Ano91o]. 1950-1973 [KP93b]. 1990 [Ano90f, Ano90x, Ano90q, Ano90v, Ano90u, Ano90t, Ano90s, Ano90w, Ano90r]. 1991 [Ano91p, Ano91s, Ano91r, Ano91q, Ano91n, Ano91t, Ano91o, Ano91u]. 1992 [Ano92n, Ano92r, Ano92m, Ano92p, Ano92q, Ano92o]. 1993 [Ano93p, Ano93q, Ano93l, Ano93m, Ano93k, Ano93n, Ano93o, Ano93j]. 1994 [Ano94s, Ano94t, Ano94o, Ano94u, Ano94p, Ano94n, Ano94r, Ano94q]. 1995 [Ano95p, Ano95q, Ano95m, Ano95o, Ano95l, Ano95k, Ano95n, Ano95r]. 1996 [Ano96c, Ano96y, Ano96w, Ano96x, Ano96r, Ano96s, Ano96u, Ano96t, Ano96q, Ano96v, Ano96p]. 1997 [Ano97-29, Ano97z, Ano97-28, Ano97x, Ano97y, Ano97r, Ano97v, Ano97u, Ano97-27, Ano97s, Ano97w, Ano97t]. 1998 [Ano98s, Ano98u, Ano98r, Ano98t, Ano98l, Ano98m, Ano98o, Ano98n, Ano98p, Ano98q, Ano98k]. 1999 [Ano99n, Ano99o, Ano99p, Ano99j, Ano99l, Ano99i, Ano99g, Ano99k, Ano99m, Ano99h].

2/3 [Ano92b]. 223-250 [Hal94b]. 26 [Ano98p]. 267 [Ano92b]. 267-272 [Ano92b]. 272 [Ano92b]. 291-308 [KO93].

3 [Ano91r, Che97]. 377-398 [Gil93].

47 [Ano92b].

75 [RU05].

95-120 [BCL93].

ability [WC95]. absence [Yan98]. absolute [Wei90]. accommodate [FLW98]. account [Blo96]. Acknowledgement [Ano90a, Ano90b, Ano91a, Ano92a, Ano92b, Ano94a, Ano96a, Ano97a]. Acknowledgement/Erratum [Ano92b]. across [HMS91, KP97b, KP97c]. across-regime [KP97b, KP97c]. Adaptive [Hod98, Ste92a, Ste95, JHB90, Phi94, Ste92b, Pöt95]. addiction [Lab99]. Additional [ST98, Kem92]. additive [GMP99, MMF93, mWK96]. additively [CM97]. adequacy [ET96, Per96, Spa90]. adjusted [BG98b, Naw90b, Naw90a, Sim93]. adjustment [Att95, Die93, GP93, GS95, Law90, MNS94, Mou95, SW94]. adjustment-costs [Law90]. advanced [HS98c]. after [Gil92]. against [BL95a, Bie97b, Cha93, KPSS92, LS96a, LT94, LT99, PS91]. aged [MWS91]. agent [PM99]. aggregate [BM93a, Dic93, Dur93, Hus92, Lew90]. Aggregation [HS90, BS90a, DFS94, GS95, NS96, PPL93, PS95c]. agricultural [BFR97, CFJL98, DL97b]. agriculture [Kum96]. AIDS [BM97]. airline [DGGV97]. Aitken [Cra92]. algebra [EG93]. algorithm [KP96, Ruu91]. allocation [Bha97b]. allocative [Kum97]. along [RS93]. Alternative [AS97a, Bro93, HS96, MS95, PS90a, BC94, Cam92, Cow91, Das97b, Gon94, KPSS92, LWG93, NF98, She94]. alternatives [Cha93, LS96a, PS91]. America [Ano90f]. among [BGH^{+}97, KNI96]. analogue [CEH90]. Analysis [Bha97a, CR94b, Ham90, JM91, LC97, Qui98a, AM93, Ana98, BM95, Bie97a, BH97, CP98, CG95, Chi95, Con91, CD98, Das91, DKF98, Era97, FH90, FHP97a, FHP97b, FKL98, GS92, Goz97, GU90, HM93, HMS91, HS98c, Joh92, JJ92, JS99a, KVC98, KP93a, KPP96, Koo96, KLOS97, KOS97, KP93b, Kwa99, LR94, LR96a, LM90, LS97b, Mai90, MR94, NVvS91, Pai97, PPL93, Poi96, Ryu93b, SvD91, ST90, Ste91, SR91, TZL96, TLL^{+}90, WWC97, Weg96, WY97, Zel96]. analysis-of-variance [NVvS91]. analytic [DC96]. analytical [Ste91, TP93]. Analyzing [Böc98, WY97, DL95, HS94b]. Andersen [Rui97]. Annals [Ano97d]. annealing [GFR94]. Announcement [Ano90c, Ano90d, Ano90e, Ano91b, Ano91c, Ano91d, Ano94b, Ano96b, Ano97b, Ano98a, BGK92]. Announcements [Ano97c]. ante [PJ96]. anticipation [BM99b]. anticipatory [MD98]. application [AH93, Att91, Bie97b, CCHS90, CR97, CN96, CKT97, Dav98, DR97, Gön98, HAD93, JJ94, LW98a, Lew95, Luc95, Mou91, NF98, Paa97a, Paa97b, PS92, PS98a, PN96, Rei90, Run97a, Run97b, SR91, TLL^{+}90, Woo91a]. applications [BS90a, Goz97, Joh95, MMF93, MX95a, MX95b, kMZ93, Pal95, Phi95a, Ric95, Woo90]. applied [CD97b]. Applying [DR97]. approach [AC93, Cha93, Chi93, DS98, DT94, DG96a, DH92, DSK99, EW98, FBA91, Fra94, GMP99, HS94b, Hol95, Hos96b, Jen94, McC94, MN97, MR91, Mou91, OP97, PH91, PSS98, PS90b, PP93, PM99, Phi91, PW91, PHH91, Qui98a, Sch96, Sch97, ST90, SR92, SK97, Stu95, mWK96, Woo90, XTZ10]. approaches [BYN98, HM93, Ind93, RS96, Ste91, Zel98]. Approximate [MS98, Sma94b, CS90, Thu92, Zab93]. approximating [BHSJ90, RS96]. approximation [GH90b, HS93]. approximations [MNS94, Mro99, Nel90, Per96, XP98]. April [Ano95n, Ano96u, Ano98o, Ano90q, Ano91p, Ano92n, Ano93p, Ano94s, Ano97-29]. arbitrage [MR91]. arbitrary [BCL92a]. Arch [HRS92, ENR90, BCK92, EM92, GM92, fL99, LN99, Nel90, Nel92, NF95, Nel96]. ARFIMA [KLOS97, MW99]. ARIMA [GMP99]. ARMA [BMR99, BDR94, CG94, GZW95, Hod98, vdL94]. arrival [AN98]. Ars [Sha96]. artificial [FD97]. asbestos [LO91]. aspects [HLvD95, Wei93]. Assessing [Fre95, Sch97, MWS91, ORS96]. assessment [KP94]. Asset [ENR90, AS98, Dan94, GHT90, GH90a, Sha90, Smi99, Stu95]. assistance [Era97]. Association [vdB97]. assumptions [AS97a, Cow91, Woo95]. Asymmetries [BEP96, Tho94]. asymmetry [Hus92, PP98]. Asymptotic [AK94, And91, CP92, Eas91, Hos96a, Kwa99, LS94, Nel96, AS97b, Buc95a, FG96, GS91, JM91, MMF99, MNS94, NP94, New97, Per96, Run97a, Run97b, ST98, WL99]. asymptotically [BGW91]. asymptotics [Hal94a, KW91, Phi98]. attitudinal [HK98]. auctions [Paa92]. Augmented [Dua97, GZW99]. August [Ano90t, Ano91q, Ano93q, Ano94t, Ano95p, Ano96y, Ano97z, Ano98s, Ano99n]. Author [Ano94c, Ano98b, AC94, AC97, Ano92t]. autocorrelated [Dag94, KS91, Oht90]. Autocorrelation [KM97, AS93, BL91a, DT94, DK91, Eva92, God98b, KZ90, Miz95, SK93, Sma94a, Wes97]. Autocorrelation- [KM97]. autocorrelation-consistent [Wes97]. autocorrelations [Bie93, Hos96a]. autocovariance [Kar99]. autocovariances [Hos96a]. automobiles [AH93]. autoregression [BFH97, HT97, ZHkM91]. autoregressions [BM93b, TP93, TY95, Zha99]. Autoregressive [HS94a, AR94, BBM96, BKS96, CP99, Chi93, CKT97, DNSW92, ET96, JS99b, KD97, LS97b, PH91, Per96, PS98b, RS97, SL97, Shi93, Tur98, YT98]. availability [EKS98]. average [CS90, GZW99, HT93, PH91]. Averaged [LR96b]. averages [KT96, PS96b]. averaging [CEH90]. aversion [CEK92]. axiom [RD99].

B. [Ste95]. background [Poi97a, Poi97b]. balanced [Oht99]. bands [KB99]. bandwidth [PS96b]. Bangladesh [MS97]. banking [FL90]. banks [CCHS90]. Bartlett [Att95]. base [Pen99]. base-independent [Pen99]. based [BG98a, CT90, Cho94, DM92, Dur91, EW98, GH90a, GH90b, GRD99, GM93a, GH96, Hal92, Hal94b, Hor94, Kao99, Kle93, KM90, Kwa99, Lan97, Naw90b, Naw90a, PK96, RK97, Vin98, WWC97, Woo91a]. Basel [Sti96]. basic [PP94]. Bayes [Chi92, Chi93, CG94, CGW98, HS94b, JHB90, KP99, Lan97, YP96]. Bayesian [Kar17, Ana98, BGW91, BKS96, BL95b, BMR99, DKF98, DH92, Dor95, FHP97a, FHP97b, Gew96, HLvD95, Kai98, KNI96, KS91, Koo91, KP93a, KOS95, KLOS97, KOS97, Kwa99, Lub95, MR91, Mil92, kMZ93, Mou91, Osi91, OS93a, OS93b, Pai97, Pal95, Phi95a, Phi95b, Poi96, Ric95, Rug94, SvD91, Sch96, SK97, SK96, Ste91, SR91, Stu95, mWK96, ZHkM91, Zel96, Zel98, vdBKOS94]. beef [NF98]. before [LR96a]. behavior [AR94, BW93, GW93, Hal93, Kra93, SV99, Squ94, Ste92b]. behaviour [God98a]. belief [KDM95]. benefits [Sch97]. Bernoulli [Sha96]. Bernoullis [Sti96]. beta [MX95a, MX95b, Wil96]. beta-II [Wil96]. between [BDR94, Dag90, HM95, LY96, Paa92, Poi97a, Poi97b, Sto94, WW96, dV91]. Bias [KP94, EKS98, HS98c, Kiv95, KPS95, MS98, MS97, RSU96, RU05, Sch94, Sto96, VV99]. biased [KS97]. Bierens [dJ96]. bills [ENR90]. binary [AM93, Ahn95, Bol96, CR97, IT98, Kle93]. binding [CL92, VK90]. birth [RW91]. bivariate [Koo91, vdB97]. Black [BSTY91, CL96]. Black-white [BSTY91, CL96]. Blaschke [LR94]. blind [BGH^{+}97]. Block [Zha99]. blocks [Fit98]. Board [Ano90g, Ano90h, Ano90i, Ano91e, Ano91f, Ano91h, Ano92d, Ano92f, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano94d, Ano94e, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano98c, Ano98d, Ano98e, Ano90j, Ano91g, Ano92e, Ano94f]. bond [Sch96]. Bootstrap [Hor94, KB99, DC96, Fit98, God98a, Goz97, LW98b, RCCA93, Vin95, WWC97]. Bootstrap-based [Hor94]. Bootstrapping [Bro90, LM97]. both [NV90]. bound [Ai94]. Bounded [Per90, Per91]. Bounded-influence [Per90, Per91]. Bounding [Bol96, Iwa96]. Bounds [Fri90, CFJL98, Cow91, GU90, Hol95, NP93, Ril93, TLL^{+}90, Tho93]. Box [BOYF94, Buc95b, KCH95, Run97a, Run97b]. Brazil [TS97]. break [Ino99, LMN98]. breaking [Per97]. breakpoint [DC96, GGH98]. breaks [AKS97, CEH96, GNW96]. British [KP93b, Paa97a, Paa97b]. budget [FFM96, Sim98]. Business [CD98, FG98, Bou95, FH90, Ghy97, Hus92, NRW93]. BVAR [RS93, Zel93]. cable [WW96]. Calculating [Chi96, Ril93, PHH91]. calorie [BFR97]. calorie-income [BFR97]. Canada [LO91]. Canadian [HAD93]. canonical [BOYF94, LS94]. capacity [Das97a, HY92, SS90]. capital [BM95, GH90a, PN96, Sch97]. cardinal [vP91]. care [Phi97]. Carl [Kle98]. Carlo [ACS99, BCL92a, BCL92b, BCL93, BKS96, Buc95a, CCN98, DM92, GLN94, Hau96, Lee97c, Lee97b, Lee97a, SK98, SV99, TM98, Tsi99, Vij97]. Case [LI96, BHHS94, CFJL98, CK97, EC96, HU94, LO91, MS95, MM94, McC98, ÖJ91]. Case-control [LI96]. categorical [BM98]. category [EW98]. cattle [NF98]. causality [CN96, FK93, Gri96, KO90, KO93, Tho94]. causality-in-variance [CN96]. cement [Das91]. censored [AP93, BS94a, Chi92, GJP97, GH99, HP94, Naw90a, VV99, Wan98, Wei93]. Censoring [HM98]. center [Ano92c]. centered [HU94]. CERGE [Ano92c]. CERGE-the [Ano92c]. chain [BKS96, CCN98]. change [Chi98, DG96b, DK96, GP96, GGH98, HR96, Hod92, LT94, PK96]. change-point [Chi98]. changepoint [ALP96]. changes [Bai99, Ham90, HS94a, Kua98]. changing [FD97, HS98b]. chaos [BGH^{+}97]. characteristics [FGM90]. chi [HS94c, VW93]. chi-square [HS94c, VW93]. child [fLRP97]. choice [AM93, Ahn95, DKF98, EW98, EKS98, Gön98, HK98, Hor93, IT98, Kle93, Lan97, fL95, Lee97a, LY96, Man93, Mat93, NV92, PS98a, PS96b, Sma94b, Tho93, WWC97]. choice-based [EW98, Lan97, WWC97]. choosing [Poi97a, Poi97b]. Christ [Kle98]. circumvent [BL91a]. class [Kin91, PO95]. classes [MMF99]. Classical [HLvD95, BL95b, KS91, Mag90]. classification [Ano90f, DFS94, Ano90f]. closed [vdL94]. Closing [DW96]. Cluster [HMS91]. clustering [Ano90f]. Co [DeJ92]. Co-integration [DeJ92]. coarseness [Cam92, Cau92]. Codependent [VE97]. coefficient [Bro93, Che99, DK91, FBA91, GKS91, LH96, SK97]. coefficients [IT98, OG93, Osb91, Soo90]. Coherency [BS94a, VK90, vSKK93]. cointegrated [Cho94, GNW96, HS98a, JS99b, LC97, LS97b, NP97, Urb95]. cointegrating [BOYF94, Han92a, Hod98, Ino99, LM97, Sne99, Wic96, Yan98, Yik98]. Cointegration [CEH96, Joh92, Koo91, LT96, PS96a, Bie97a, Bos95b, Can98, Cho94, Com99, Dav98, Dor95, EC96, GS95, GL94, GH96, Hal94a, Han92b, Hau96, HEGY90, JJ92, Joh95, JS99a, Kao99, KvD94b, Lau99, Lee92, OP97, Osb93, PPL93, Qui98a, RKJ99, SL97, Seo98]. collinearity [Soo90]. Columbian [Paa97a, Paa97b]. combination [HK98]. Combining [HLS98, kMZ93, Ste91]. Comment [Pöt95, KR94, Ste95]. Comments [AS97b, GKW95, Pal95]. commercial [CCHS90]. common [AV98, CW97, Can93, Gew93, McM92, Paa92, Wic96]. company [BBDS92]. comparative [BC94, Cha90]. Comparing [KS91, Poi97a, Poi97b, FK94]. Comparison [BOYF94, BG98a, BHF90, Chi98, DC96, Hau96, Ind93, LWG93, MS92, Thu92, TW91, Zab93]. comparisons [GRD99, HKU97, Pol91, SR92]. compatibility [BS90b, KR94]. competing [Sue92]. competition [BGH^{+}97, CD97b, EW98]. competitive [KVC98]. compliance [Era97]. component [BL91a, BCL92b, BCL93, BC94, BL95a, CKT97, HRS92, MN95, MP99]. components [AV98, AB95, DQ93, Fin93, GZW95, GZW99]. compound [Pai97]. computation [KWvD97, Pin93]. computational [HMR96]. computers [BGR95]. Computing [AKS92, PP93]. concept [VB90, vP91]. concepts [PP94]. concerning [GKS91, HU94]. condition [Hah99, KB95, dL97a]. Conditional [Bos95a, Eri95, Ahn95, BB92, BBM96, Bos94, Das97b, GHT90, HS94a, KS93, LT96, LN99, MS92, PS90a, Rob91, Seo99, SV99, Woo90, Woo91a, Woo95]. conditions [AB91, BB98, BDR94, BQSW99, CKT97, GH90b, Hon93, KM99, LB97a, LB97b]. Conduct [Cor99]. Cone [CCHS90]. Cone-Ratio [CCHS90]. cones [WY97]. conference [Van92]. Confidence [HU94, KB99, RCCA93]. Conjectandi [Sha96]. conjugate [KP93a]. consideration [CCN98]. consistency [Eas91]. Consistent [CF99, Lew95, Li99, KM97, LW98b, Pin98, Wes97, Zhe96]. constancy [LT94, LT99]. constrained [Phi91, XTZ10]. constraint [Sim98]. constraints [AF92, CL92, DR97, VK90, vSKK93]. construction [SR92]. consumer [AR98, CL92, DC98, HS90, KB95]. consumption [BW93, EGHL93, GH90a, GW93]. consumption-based [GH90a]. contaminated [LI96]. contemporaneous [NS96]. content [BL91b, DL92]. context [BG97a]. continuity [dJ98]. Continuous [DW96, AL97, Cha93, CL92, CR96, Com99, KP93b, LT94, Mro99]. continuous-time [AL97, KP93b]. Contracting [PS98a]. control [DM92, LI96]. controlled [BGH^{+}97]. controls [LI96, NV92]. Convenient [BL98]. Convergence [New97, BD96, LR96a, PS96a, PP94]. CORNER [Vin98]. Correcting [EKS98]. correction [Bos94, Bos95a, Bos95b, Eri95, KP94, KvD94b, MS98, Qui98b, Sch94, Seo98]. corrections [MS95, Woo95]. correctors [Miz95]. correlated [Are93, CG95, Iwa92b, Mic99]. Correlation [LS97a, Fre95, KP97b, KP97c, Kyr98, LH98, Rob91, Wei90]. Corrigendum [Kar17, RU05]. cost [BFR95, BW95, FL90, KOS97, Kum97, MN97, Men94, Mou95, PJ96, PHH91]. cost-effectiveness [MN97]. cost-minimizing [BW95]. cost-of-living [PHH91]. costs [Law90]. count [Böc98, CGW98, CD97b, GV97, GM97, GRS99, Ter98]. counterparts [BG97a]. countrie [ZkM99]. countries [HMS91, Tyb92]. country [KO90, KO93]. coupon [EKS98]. covariance [Ana98, BG98b, Buc95a, ENR90, GJ92, RR90, Wes97, vdL94]. covariances [MMF99]. covariates [LW90, Seo98, Sue92]. Cox [Buc95b, KCH95, Mic99, PP93]. Cramer [Ano92b]. criteria [GKW95, SW96, TW91]. criterion [KS90, PW91]. critical [Hor94, ST98]. criticism [Iwa96]. critique [HY92, ORS96]. cross [AL96, And91, Col97, Con99, CSS90, Fre95, Mof93, NV90, PW95, VN93]. cross-section-time-series [AL96]. cross-sectional [CSS90, Fre95]. cross-sections [Col97, Mof93, NV90, PW95, VN93]. cross-validation [And91]. currency [BGHT95, MT90]. current [Ray91, PW95]. curve [BHSJ90, RS96]. curves [HNP95, Hol93, SCS99]. customer [HAD93]. customer-dialed [HAD93]. cycle [Bou95, CD98, FG98, FH90, Ghy97]. cycles [AB98, Hus92, VE97].

D [Pöt95]. dagger [KP99]. data [AS95, AL96, AS97a, Ang91, Are93, AC93, BG98a, BCL92a, Bal95, BHL96, BGHT95, Bar97, BM93a, BW93, Bha97a, BR97, BBDS92, BB98, Böc98, BFR95, BM98, Cam92, CFJL98, CGW98, CSW92, CD97b, DL95, DKF98, Dea90, Dic93, DGV98, EKS98, FOS97, FKL98, Fre95, FH91, GJP97, GS92, Gön98, GM93a, GV97, GM97, GW93, Hah97, Hal92, Hal94b, HK98, HLS98, HAD93, IASW99, JM91, Kao99, Kiv95, KS97, KM97, Kum90, LS96b, LH98, Luc95, MN95, MRB98, Mai90, Met96, Naw90b, Naw90a, NRW93, Rom97a, Rom97b, Ryu93b, Saï91, Sim93, Spa95, Ter98, TT93, Tyb92, VV99, WY97, Woo95, Woo99, dJ96]. DEA [CFJL98, CCHS90, ST90, Sen90]. Debt [DGGV97]. December [Ano92o, Ano90v, Ano91s, Ano96w, Ano97-28, Ano98u, Ano99o]. Deciding [Paa92, Sto94]. decision [Era97, Mai90]. decision-making [Mai90]. Decomposing [BR97, Men94]. Decomposition [BHHS94, OR94]. decompositions [BM93b]. degree [BK98]. Demand [AB91, HW96, Att91, BG97a, Cha90, CL92, Des98, DR97, FFM96, Goz97, HK95, HS90, HS98c, KB95, KS97, Lab99, Law90, MR92, PN96, SR91, Ste96, VK90, vSKK93]. demographic [Goz97]. Dennis [Zel91]. densities [Had96, Smi94]. density [FG96, Iwa92a, PS96b, Ryu93a]. density-weighted [PS96b]. dependence [CR94a, Con99, Hos96b, JT94, dJ96]. dependencies [Böc98]. dependency [BS90a]. dependent [Abr99, BB92, BS94a, BSH93, DT94, God98b, GH99, HASM98, Hon93, Kai98, Lai93, NP94, NV90, Per96, PS92, Pin93, VB90]. derivatives [HMR96, HT93]. Deriving [Paa97a, Paa97b]. designs [BYN98, NVvS91]. detect [MS97]. Detecting [AKS97, Rei90, SHHS94]. detection [BCdL98]. determination [Par96, PS92]. Deterministic [Abe94, BHF90, Han92a, KO90, KO93]. detrending [HS96]. developing [Tyb92]. development [CD97b]. developments [Bau90, DG96b, ST90]. deviations [AS97b]. device [HS98c]. diagnosis [Stu95]. diagnostic [BS90a, HK91, Hil91]. diagnostics [GHT97, McC94, Woo91a]. dialed [HAD93]. Dickey [GZW99, Hal96, LMN98]. Diewert [Men94]. difference [Cha98, HP94]. differencing [BK96, CD94]. different [Ind93, MZ90, Sch90, Woo96]. differential [Cha98, LW98a]. differential-difference [Cha98]. differentials [CL96, OR94]. differently [FH91]. diffusion [Dua97, Nel90]. diffusions [HST98]. dimension [HT99a]. dimensional [DL95]. dimensions [vP91]. Direct [KvD94b]. disability [Ste96]. disaggregated [PPL93]. discovery [Wes96]. Discrete [CL92, Com99, Mro99, BH97, BR91, Cha93, Gön98, HM93, HASM98, jL92, Lee97a, LW90, PS98a, RCCA93, Sma94b, Tho93]. discrete-choice [PS98a]. discrete-response [HASM98]. Discrete/continuous [CL92]. discrimination [CR94b, Jen94, NS94, OR94, SHHS94]. Discussion [Die93, Hin97, Zel93, Ric95]. disequilibrium [Ai94, LS94, Lee97c, Lee97b, Lee97d, Lee97e, MD98]. dispersion [BS90a, CL96]. dispersion-dependency [BS90a]. distance [HAD93, TSB96, Yit94]. distributed [AF92, Gre90b, KD97]. Distribution [AM93, Che99, Seo99, Woo99, BS90a, CP92, DSW90, HSPHS90, HM94, IT98, Lew90, MX95a, MX95b, NP94, Run97a, Run97b, WL99, Wol90]. Distribution-free [Che99, Woo99]. Distributional [PS91, And94, BR97, Jen94]. distributions [CP98, CA91, Chi96, EMS99, GZW99, Hal96, Hos96a, Mic99, Pai97, Wil96, Yit94]. disturbance [KW91]. disturbances [BCL92a, BL95a, Dag94, FMB92, GZW95, Gil91, Gil92, Gil93, HRS92, OG93, RK97, Shi93, Sma94a, Smi94, SM95, Tsi99, Tur98]. dividends [Koo91]. Do [Jus95]. Does [BM97]. domain [CP93, RS93, XP98]. dominance [CFJL98, OK96, PW91]. Don’t [Miz95]. Double [Vin95, DC98, Lab99]. double-hurdle [Lab99]. drift [Bie97b]. drifts [Ent97]. dual [AC93]. due [HM98, Rom97a, Rom97b]. dummy [Che99, Mro99]. Duration [GM97, Blo97, FGM90, Gri93, Lil93, McC94, Rom97a, Rom97b, Ryu93b, TT93]. durations [FG98, Ghy97, vdB97]. Durbin [AS93, AKS92, Bar92, HM94, KW91, PL91]. Dynamic [AS98, DQ93, Fin93, Kim94, Man93, AS95, AS97a, AH98, BB92, BG97a, BB98, BEP96, Col97, Des98, DK96, FLW98, GP96, Gön98, GHR94, Hah99, Han99, HT99a, KW91, Kiv95, KPS95, Lee97c, Lee97b, Lee97d, Lee97e, Lee97a, fL99, Lüt97, Mof93, NER92, Now93, PS95a, PN96, Rob91, TZL96, Vog99, Wei91]. dynamics [BFR97, BM99b, KvD94a, PP98, PST96].

Earnings [Jen94, HS94b, MN95, PK94]. Econometric [AH93, BGR95, EW98, FL90, Sim98, Bau90, BL95b, CF99, Chi95, DR97, Ghy93, Li99, LM90, PS90b, SR92, Tsi99, US94, Van92]. Econometrics [Ano92b, Ano97b, Ano97d, Bar97, RU05, Ano98a, Bai96, CPW93, DG96b, Gew96, Kar17, KM95a, KM95b, KvD94a, Kle98, MS98, Mir95, NS94, PS95b, Spa95, WW98, Ano96c, AC94, AC97]. Economic [Yit94, Ano92c, BM95, BM97, BEP96, Gra92, GKW95, KP99, KPSS92, SS90, Slo90]. economics [BS90a, Ind93, Pel96, Phi97]. economy [PPL93]. Editor [WW98, Aig90, Bal95, BPvD96, Bha97a, BEP96, Ghy93, Kle98, LL90, Maa91, Poi91, Slo90]. Editorial [Aig91, Aig96, Ano90g, Ano90h, Ano90i, Ano90j, Ano91e, Ano91f, Ano91g, Ano91h, Ano92d, Ano92e, Ano92f, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano98c, Ano98d, Ano98e]. Editors [Ano97l, BK96, BL95b, CM90, CPW93, DG96b, FP95, HK91, HR93, KM95a]. education [Ano92c]. effect [BW93, Con91, Die93, Ent97, GP93, GP96, GW93, MN95, TP93]. effectiveness [MN97]. Effects [Soo90, Abr99, Are93, BL92, BS94b, BM93b, DHSR98, DS98, EGW90, FLW98, FKL98, Goz97, GM93b, Hah99, Han99, Hon93, KOS97, KP93b, fLRP97, Ste96, Sto96, Ter98, VN93, WW96]. efficiencies [Mai90]. Efficiency [GRD99, NP93, SM95, Ai94, AC93, BHF90, CSS90, FG95, FL90, GT99, Hol95, Joh92, Kem92, Kiv95, KOS97, Kum96, MMF99, Men94, NVvS91, Ril93, Szr94, TLL^{+}90, Tho93]. Efficient [AS95, AS97a, ACS99, Bos95b, DK97, Hah97, Han92a, IASW99, IL96, Tur98, Mai90, MS93]. elasticities [Dea90]. elasticity [HW96]. electrical [PN96]. electronic [HS98c]. elliptical [Osi91, OS93a, OS93b]. Empirical [BR91, ENR90, KDM95, PJ96, Sha90, BCK92, DGGV97, HS94b, JHB90, KVC98, McA95, ORS96, Paa92, Pal95, Phi95a, Ric95, Sch96, Slo90, Urb95]. employment [CR94b, Gri93, Hus92]. Encompassing [GHR94, MM94, Woo90]. end [AF92, FBA91]. end-use [FBA91]. endogeneity [NN98]. Endogenous [PN96, Att91, Che99, DC96, Iwa94, Lau99, Mro99, Ter98, VV99]. Engel [HNP95]. Engines [BT95]. entropy [EMS99, Hol93, MZ90, Ryu93a]. envelopment [GS92, Mai90, WY97]. environment [HR96]. epidemic [BM97]. Epstein [Smi99]. Epstein-Zin [Smi99]. equality [vSKK93]. equation [AK92, AV98, Hal94a, Hod92, Joh92, Kai98, KD97, Lee94, MMF93, Mro99, PS90b, Ste91, SR91, TM93, Zel98]. equations [Att95, Cha98, CP98, CP92, CSW92, Dhr94, Fri90, FFM96, Joh95, KR97a, KR97b, Kum97, Lil93, ORS96, RR90, Sch90, ST95, Smi94, Spa90, Tur98, Woo96]. equicontinuity [PP94]. equilibrium [AS98, AH98, CK97, Gon94, PS96a]. equity [BM99b]. equivalence [BL91b, MMF99, OS93a, Pen99]. Erratum [Ano90k, Ano90l, Ano91i, Ano92b, Ano99a]. error [AB95, BYN98, BL91a, BCL92b, BCL93, BC94, BL95a, BS90a, BW93, Bos94, Bos95a, Bos95b, CKT97, DT94, Eri95, GZW95, Gil92, GW93, GJ92, KP94, KvD94b, MP99, Phi98, Qui98b, RSU96, RU05, Seo98, SC93, WWC97, Weg96, Wei90, Wei93]. error-component [BL91a]. error-components [AB95, GZW95]. error-correction [KP94]. errors [Ame90, And91, Can98, Chi93, CG94, CG95, Dag94, DD97, DNSW92, DFS94, DK91, Era97, Fri90, HS93, HNIP91, HNP95, Hod98, Iwa92b, Kar99, KS91, NP94, Now93, Oht90, Osi91, Per96, PM99, PL91, SS97a, SS97b, TW91, Wan98, ZWG91, vdL94]. errors-in-variables [Ame90, Fri90, HNIP91, Iwa92b, Now93, Wan98]. estimate [Paa97a, Paa97b]. estimates [BGR95, Chi96, DL97b, DR93, ENR90, GS91, Mag90, Pen99, Pin93, PK94, Ste96, WL99, Zel98]. Estimating [AL97, Buc95a, Chu96, Col97, Gön98, Ind93, KS97, Lab99, LS94, Mro99, PS92, PS95a, Ter98, Tyb92, Woo96, AH93, Gon94, GJ92, KNI96, MRB98, OP97, Oht90, Phi91, XTZ10]. Estimation [AR94, AM97, AF92, Att91, Blo96, Chi98, FH91, GHT97, Gil92, GJP97, GM93b, Han99, HNP95, HKU97, JM91, fL99, MP99, NP97, NV90, SK98, WWC97, Wan98, ZWG91, Abr99, AM93, AP93, AS95, Ahn95, AS97a, ACS99, Ang91, AB95, BGHT95, BGW91, BKS96, Bau90, BMR99, BHF90, BS94a, BSH93, BCdL98, BM98, Cha98, Che97, CR97, Che99, CR94a, CD94, Con99, Cra92, Dag94, Dan94, Des98, Don95, DK97, Ent97, FBA91, GZW95, GH99, GL94, GRS99, Hah97, Han92a, HNIP91, Hod98, Hor93, HM98, Ich93, IT98, IASW99, IL96, Iwa92b, Iwa94, Kem92, KCH95, KP97a, KM90, LI91, Lee94, fL94, fL95, Lee97c, Lee97b, Lee97a, LW90, LS96b, LR96b, MW99, Mat93, MD98, MR92, Mof93, Mou91, Naw90b, Naw90a, PS90b, Phi94, Pöt95, RM92, RT99, Rug94]. estimation [Rui94, Ruu91, Ryu93a, SC93, Sow92, SS97a, SS97b, Ste92a, Ste95, Sue92, Tur98, VV99, VN93, VW93, Wei90, Wei91, Wol90, Woo91b, Woo99, Zhe96]. estimator [BS94b, CS90, HS94c, Iwa92a, JHB90, Kwa99, Lob99, McC98, MG91, MS93, Oht99, ÖJ91, PS98b, ST95, TSB96, Wes97]. estimators [Ame90, BC94, BG97a, BL98, BOYF94, Bro90, BG98b, CS98, CP92, DD97, Eas91, FG96, FMB92, Fri90, GT99, GRD99, Hal92, Hal94b, HT97, Hod92, HP94, HU94, KB99, KS91, Kin91, Kiv95, KPS95, KR97a, KR97b, Mar93, MS92, New97, OK96, PO95, Per90, Per91, QS99, RSU96, RU05, Rui97, Smi94, SM95, Ste92b, Szr94, Vin95, Yan98, Yik98, Zel98]. ethnic [NvPV94]. Euler [GH90b, ORS96]. Evaluating [Gra92]. evaluation [Phi97, Vij97]. events [KP93b]. Evidence [BFR97, BBDS92, BFR95, DeJ92, KP99, NRW93, TS97, BCL92a, BCK92, DGGV97, FL90, FK93, GS92, GS95, Hus92, Per97, YT98, PW95]. ex [PJ96]. Exact [DK96, KD97, Smi94, Szr94, Fri90, JS99b, Kum97, MR94, OK96, Sma94a, Thu92]. examination [HK95]. example [Jus95]. examples [US94]. excess [CEK92]. exchange [PS92, SvD91, WC95]. existence [Kin91]. existing [BCL92b, BCL93]. Exogeneity [TM93, PR99, SR91]. exogenous [IASW99]. expectations [AM93, Cro98, Gön98, HS93, JS99b, NF98, PS92, Ray91, SS92, Sim93]. expenditures [GS97a]. experiences [Man93]. Experimental [DR93, CFJL98, NV92]. experiments [Car95, DM92]. explanatory [Att91, IASW99, Iwa94]. exploration [PP94]. exploratory [BM95]. Exploring [HM95, DC98]. exponent [WL99]. exponentially [ZHkM91]. export [Law90]. extended [Luc95, Vij93]. extension [BYN98]. Extensions [Dea90, Ruu91, CG95, GLN94, KVC98]. extreme [CM97, GU90, Sma94b]. extreme-bounds [GU90].

F [Eas91, Kle98]. facet [BYN98]. factor [ENR90, KS91, Mro99, NER92, PP98, PN96, Weg96]. factor-arch [ENR90]. factors [CGW98, KP99, YP96]. failure [dL97a]. falsification [KB95]. family [HT99b, SCS99]. farm [Kum96]. farm-level [Kum96]. farming [TLL^{+}90]. Farrell [FG95]. feasible [SM95]. February [Ano90s, Ano91r, Ano92m, Ano93m, Ano94o, Ano97x, Ano95q, Ano96x, Ano98r, Ano99p]. Fellow [Ano92h, Bar97, Gra92, Maa92]. Fellows [Ano96c, Ano97b, Ano98a, Vin98]. fertility [Lil93]. Fighting [KS91]. filtered [Yik98]. Filtering [Nel92, NF95, Nel96]. filters [Die93, GP93, GP96]. FIML [McC98]. finance [BCK92]. financial [AS98, Ano92s, CN96]. Finetti [Pel96]. Finite [GS92, Hod92, KZ90, McC98, PS98b, Smi99, CA91, CP92, LS94, MG91, Ste92b, Szr94, YT98, Zel98]. Finite-sample [Hod92, PS98b, CA91, MG91, Szr94]. Firm [GR95, Squ94, BFR95]. firms [BW95, HM95]. first [CS90, DK96]. first-order [CS90, DK96]. fish [HK95]. fit [CW97, FG95, Var90]. Five [Gon94]. fixed [Abr99, Ent97, Hah99, HT99a, Hon93, VN93]. fixed-effect [Ent97]. fixed-effects [Abr99]. Flexible [DW95, BT98, GS91, LH96, RS96]. flow [LI95]. fluctuations [Dur93, PPL93]. force [Ste96]. forecast [NF95, Phi98, TW91]. Forecasting [BT98, Can93, Cha90, Gew93, ZHkM91, ZkM99, kMZ93, Nel92, NF95, Wei91]. forecasts [DL95, MS92, kMZ93]. foreign [MT90]. form [And94, BHSJ90, Cra92, GS91, HS98a, McC98, RS96, Zhe96, vdL94]. formation [BM95]. forming [CFJL98]. forms [DW95, US94]. formulae [OK96]. Foundations [Vin98]. Fourier [GS91]. Fractional [BK96, Bai96]. Fractionally [BBM96, CD94, LS96a, Mar98, Sow92]. fractionally-integrated [CD94, LS96a]. framework [DL95, HAD93, Lub95]. free [Che99, HK95, Woo99]. freedom [Sen91]. French [HM95]. Frequency [MR92, BGHT95, CP93, Gri93, XP98]. Friedman [Pel96, ZkM99]. Frisch [Bou95]. frontier [FOS97, Gre90b, KM90, ST90, vdBKOS94]. frontiers [Bau90, CSS90, GS92, KOS97, Kum90, PSS98]. Full [Des98, Urb95]. Fuller [GZW99, Hal96, LMN98]. Fully [KP97a, NF98]. function [Bar92, CM97, DeJ92, EGHL93, HT97, Kar99, Kum97, LW98b, Oht99, SC93]. functional [BHSJ90, DW95, Hol95, RS96, Zhe96]. functions [AH93, Blo96, Dag90, GFR94, Goz97, HS94b, Per97, PJ96, PP94, RCCA93, Ryu93a, Vin98]. Further [Per97, God98b].

G [Ano92b, Pöt95]. G. [Hin97]. gains [Kem92]. games [BR91]. Gamma [Gre90b]. Gamma-distributed [Gre90b]. gap [DW96, PK94]. GARCH [BP92, DS98, DW96, DK97, Dua97, GRD99, HT99b, Kar99, NS96, She94, dV91]. gasoline [BG97a]. Gaussian [RM92, TM98]. gender [PK94]. General [BT95, BHSJ90, GZW95, HAD93, Hyl95, OK96]. generalization [MX95a, MX95b]. generalizations [EG93]. Generalized [CM97, MZ90, And91, AKS92, BBM96, Chu96, LH96, QS99, SR92, Sma94b, SM95, WY97, Wil96]. generated [Ril93]. generator [CM97]. Generic [PP94]. German [Run97a, Run97b]. Getting [Nel92]. Gibbs [Chi93]. GIVE [KP97a]. Glejser [God96]. Global [GFR94]. globally [Kem91]. GLS [BL92, KPS95, MS95]. GLSEM [Dhr98]. GMM [AS97b, Con99, KP97a, KM99, Qui98a, Rui97]. Goodness [Var90, CW97, FG95]. Goodness-of-fit [Var90, FG95]. government [Sim98]. graduate [Ano92c]. group [BS94b, SHHS94]. Grouped [Ang91, Cam92, FH91, Naw90b, Naw90a, SG91]. grouped-adjusted [Naw90b, Naw90a]. Grouped-data [Ang91, Cam92]. Grouping [Cow91, Cam92, Cau92]. groups [CR94b, NvPV94]. growth [BM97, BT95, Lau99, kMZ93, ZHkM91, ZkM99].

H. [Hin97]. Hadamard [NPL95]. Hausman [AL96, God98b]. hazard [DGGV97, LW90, vdB97]. hazards [Lil93, Rug94, Sue92]. Health [TS97, Bha97a, GS97a, HK98, KWvD97, fLRP97, MS97, Phi97, Sch97]. hedging [BH97]. hedonic [AH93, Mou95]. Heikes [CS90]. heterogeneity [AR98, Böc98, CCN98, CD97b, DC98, GV97, JT94, Lab99]. Heterogeneous [AN98, BG97a, PS95a]. heteroscedastic [Sma94a]. heteroscedasticity [Ana98, Eva92, MMF93, SG91, Szr94, Thu92]. Heteroskedastic [Han92b, And91, Don95, NPL95, PRW97]. heteroskedasticity [BBM96, CA91, Cra92, Das97b, God96, HS94a, KS93, KM97, LT96, ÖJ91, Rob91, Seo99, Wes97]. heteroskedasticity- [Wes97]. heteroskedasticity-consistent [KM97]. Hierarchical [CG95]. High [BM95, BGHT95, fLRP97]. high-frequency [BGHT95]. high-mortality [fLRP97]. High-tech [BM95]. Higher [Bie93, DD97, XP98]. Higher-order [Bie93, XP98]. Highest [Iwa92a]. history [KB95]. hoarding [Chi95]. hold [Jus95]. homogeneity [Gil92, KB95]. homogeneous [CR94b, DW95]. honor [Kle98]. Hospital [KOS97]. household [Wol90]. households [Bha97b]. human [Sch97]. hurdle [DC98, Lab99]. hypotheses [Das97b, GAR97, JJ92, Woo90]. Hypothesis [And98, BD96, Bie93, Bie97b, CF99, Dur91, EC96, FK94, KDM95, KPSS92, Lew95].

ideal [BGW91]. identifiability [Weg96]. Identification [Dhr98, HNIP91, HM98, JJ94, Mof93, RM92, Dav98, HS90, Mat93, McM92, Now93, Ray91, Spa90]. identified [CP92]. Identifying [Joh95, HST98]. idling [Das91]. if [Maa90]. II [NF95, Wil96]. illusion [Lew90]. illustration [Urb95]. illustrative [CCHS90]. image [Hal96]. impact [Cam92, DFS94, DR93, Gri93, KO90, KO93, Mro99]. implementation [PJ96]. implications [Era97, FH90, Men94, Osb91, PST96, Sim98]. Implied [EM92, MS92]. import [Law90]. importance [Duf93, NRW93]. improper [FOS97]. Improved [QS99, SS97a, SS97b, Bro90, fLRP97]. Impulse [KPP96, LS97b, Phi98, BM93b, Koo96]. imputations [HM98]. inaccuracies [TT93]. Inadmissibility [Oht99]. INAR [Böc98]. Income [Lew90, BS90a, BFR97, Dag90, DSW90, FK93, HSPHS90, MZ90, Tho94, VB90, Wil96]. income-dependent [VB90]. incomes [KP93b]. Incomplete [BC94, DGV98, SW94]. inconsistencies [Rom97a, Rom97b]. inconsistency [Kiv95]. independence [Pin98, Woo95]. independent [Col97, Pen99]. Index [Ano90m, Ano90n, Ano90o, Ano90p, Ano91j, Ano91k, Ano91l, Ano91m, Ano92i, Ano92j, Ano92k, Ano92l, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano94i, Ano94j, Ano94k, Ano94l, Ano94m, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano96o, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, AC94, AC97, Ano92t, Ano94c, Ano94v, Ano98b, CS98, CEK92, DL92, Ich93, Kle93, Mou91, RD99, Sle91]. indexes [BGR95, Mou95]. Indian [Kum96]. indicators [BR97]. indices [Fis95, Par96, PHH91, SR92]. Indirect [MW99, VK90]. Individual [PK94, BFR95, KOS97, SHHS94]. Individual-specific [PK94]. individuals [MWS91]. Indonesian [Dea90]. Induced [Hal93, Kra93]. industries [BM95, HS94b, SS90]. industry [BFR95, GR95, PN96]. inefficiency [Kum90, Kum97]. inequalities [BSTY91, KWvD97]. Inequality [JS90, RW91, Cow91, Dag90, Sle90, Slo90, vSKK93]. infer [GHT90]. Inference [CK97, HT99a, RS97, Wes96, YT98, Bos95b, Chi92, CG94, Fit98, GKR97, GJP97, GM93a, Han99, HLvD95, Hos96b, Jus95, Kai98, KWvD97, KM99, Lan97, Lee92, NP97, Osi91, OS93b, Seo98, TY95, Tsi99, Van92, Vin98]. Inferring [CD97a, DGV98, HK98]. Infinite [KT96, Can98, LS97b, PL91, Run97a, Run97b, SL97]. infinite-variance [PL91]. Inflation [Sle90, Cro98]. influence [BK98, Per90, Per91]. Information [SW96, AN98, BL91b, CP98, DL92, Dhr98, Hor94, Kwa99, Orm90, Rom97a, Rom97b, Soo90, Yan98, Zel96]. information-matrix [Orm90]. informational [Cow91]. informative [Hah99]. Infrastructure [DSK99]. Initial [BB98, BCL92a, Hah99]. innovation [CD97b]. innovations [RS97]. input [Fis95, Squ94]. instability [Kuo98, Tho94]. Instrumental [Iwa92b, Ame90, AB95, Hal92, Hal94b, Lee94, Mag90, MN97, Met96, PH91, QS99]. instrumental-variables [MN97]. instruments [Iwa92b, KP97a, Sch90, Woo96]. insurance [FGM90, GS97a, HKT90]. integrated [BBM96, CD94, HM94, LS96a, Mar98, Per96, RS97, Sch97, She94, Sow92, TY95, YT98]. integration [Bai96, DeJ92, HEGY90, Lau99, Par96, WW96, vP91]. interactions [KVC98]. interactive [GKS91]. Interarea [Mou95]. intercept [PK94]. interest [AL97, Bie97b, CKT97, Cro98, Jus95, PST96]. interior [KP96]. International [Van92, HW96, kMZ93, ZHkM91]. Interpreting [BD96, Wic96]. Intertemporal [Sha90, BMN93, GHT90, GH90a]. interval [LS97a]. intervals [RCCA93]. intervention [FH90, RW91]. Introduction [GM97, Aig90, Ano97l, BK96, Bal95, BL95b, BPvD96, Bha97a, BEP96, CM90, CPW93, DG96b, FP95, Ghy93, GS97b, HK91, HR93, KM95a, Kle98, LL90, Maa91, Poi91, Slo90, WW98]. invariance [EH93, PS96c]. invariant [AKS92, DK91]. inventory [Duf93, Hal93, Kra93, NRW93]. Investigation [Sha90, GLN94, Ghy97, KO90, KO93, SV99]. investigator [PM99]. investigator-agent [PM99]. Investment [BBDS92, ORS96]. ISLM [JJ94]. Israeli [BFR95, GR95]. issues [AH93, MWS91, Wol90]. iterative [BL92, LR96a]. IV [CP92, KP97a].

J [Kar17, Mic99]. J. [Ano92b]. J.Aigner [Zel91]. jackknife [SG91]. Jacob [Sha96]. Jamaica [GS97a]. James [HU94]. January [Ano90s, Ano91n, Ano92m, Ano93m, Ano94o, Ano95m, Ano96r, Ano97x, Ano99j]. Japanese [EGHL93]. Jeffreys [CP98, Poi94]. Job [Blo97, CR94b, TT93]. job-search [TT93]. Johansen [BOYF94]. Joint [JT94, Cha98, EC96]. Journal [Ano92b, Ano97b, RU05, Ano98a, AC94, AC97, Ano96c, Ano97d]. July [Ano90t, Ano91q, Ano92p, Ano93l, Ano94p, Ano95l, Ano96s, Ano97r, Ano98l, Ano99i]. June [Ano90f, Ano93n, Ano96t, Ano90u, Ano91t, Ano92r, Ano94u, Ano95o, Ano97y, Ano98t, Ano99l].

Kansas [TLL^{+}90]. kilns [Das91]. Koenker [God96]. Kopp [Men94]. Korea [EC96]. KPSS [LS96a]. Kullback [Dhr98].

labor [Ang91, Blo96, Chi95, Kum96, NS94, Phi97, SHHS94, Ste96]. labor-supply [Ang91]. Labour [BMN93, MW97, Blo97]. lag [BK98, KD97]. lagged [DT94, God98b, Hon93]. lags [AF92]. large [CCHS90, KM99, dJ98]. Largest [Mai90]. latent [BM98, DC98]. laws [dJ98]. leads [AF92]. Leapfrog [Abr99]. Learning [KP97b, KP97c, Man93]. Least [Wei90, FMB92, Fit98, Ich93, JM91, LH96, MS93, Sch90, SM95, Szr94, Yik98]. Lending [AB98]. level [Bie97b, HS98b, Kum96]. levels [CSS90, MZ90]. leverage [TZL96]. Li [Hin97]. life [HMS91]. Likelihood [Bai99, JS99a, Att95, BSH93, BG98b, CD94, CD97b, Dan94, DeJ92, Des98, GRD99, Gre90a, Hos96b, IT98, Iwa93, Jus95, Lee92, fL95, Lee97d, Lee97e, Lee97a, MR94, Phi91, PW91, PS98b, RK97, RT99, Rui94, SK98, SG91, Sow92, Woo91b, XTZ10]. likelihood-ratio [SG91]. limit [Dua97, PO95]. limited [BSH93, CP98, Kai98, Kwa99, Lai93, PS92, Pin93]. limited-dependent [PS92]. Limiting [NT90]. line [KVC98]. linear [ALP96, Bro90, Cau92, CK97, DD97, DQ93, DR97, DK91, FL90, Fin93, Fit98, GP96, Gil91, Gil93, GHR94, Hil91, Hol95, Joh95, Kim94, KP94, KZ90, LS96b, LT99, Mar93, McC98, Naw90b, NPL95, NV90, Now93, Oht90, OG93, PS90b, Saï91, SS92, SK93, SS97a, SS97b, Szp97a, Szp97b, TW91, Tur98, Wan98, ZWG91]. linearity [SKA94]. Lipschitz [dJ98]. Lipschitz-continuity [dJ98]. live [Maa90]. living [JS90, PHH91]. load [FBA91]. Local [BH97, HT97, NP94, She94, Weg96, Ril93]. locally [Kem91]. location [GJ92]. log [Gre90a, Iwa93, Vel99]. log-likelihood [Gre90a]. log-periodogram [Vel99]. Logan [Ano90f]. logit [Ano92b, BS90b, CR91, EGW90, HS94c, KR94, KP93a, Poi94, Poi96]. Long [Bai96, BM99b, CR96, DHSR98, BK96, BM96, BCdL98, CR94a, Chu96, FKL98, Gon94, GL98, GD96, HR96, Hos96a, Hos96b, HAD93, Ind93, JJ94, Jus95, KT96, KOS95, KLOS97, LR96b, Lob99, PS95a, TSB96]. long-memory [HR96, Hos96a, TSB96]. long-range [CR94a, Hos96b]. Long-run [DHSR98, FKL98, Gon94, Ind93, JJ94, KOS95, PS95a]. Long-term [BM99b]. longitudinal [Böc98, Rom97a, Rom97b]. look [AB95, CD98, FK93, Koo91, Ray91]. Lorenz [BHSJ90, Hol93, RS96, SCS99, Wil96]. loss [JHB90, Oht99, Pai97, Rom97a, Rom97b]. Low [Yik98]. Low-pass [Yik98]. Lucas [ORS96]. Lyapunov [WL99].

M [Ste95]. MA [BL95a]. machinery [PN96]. macroeconomic [DeJ92, GAR97, Per97]. Maddala [Hin97]. magnitude [KP93b]. maintained [Hal96]. Making [NF95, Tyb92, HKT90, Mai90]. male [Blo96]. Mallows [KS90]. management [Ryu93b]. manufacturing [BM95, HM95]. mapping [EW98]. March [Ano90w, Ano92q, Ano93k, Ano94n, Ano95n, Ano96u, Ano97v, Ano98o].

Marginal [RK97, GHT90, MN97, OS93a, Sto96]. Marginal-likelihood [RK97]. Marginalization [NS96]. market [AH93, BGHT95, BM96, BM99b, CN96, Cor99, DL92, Gön98, HK95, KVC98, MW97, NS94, SHHS94]. Marketing [AR98, WW98, BYN98, FKL98]. markets [AS98, DHSR98]. Markov [BKS96, CCN98, Chi96, Ham96, Kim94, LR96a, PS98b]. Markov-Normal [LR96a]. Markov-switching [Ham96, Kim94]. Marriage [Lil93]. Marseille [HK95]. martingale [Dur91, McC94]. matched [PW95]. mathematical [ST90]. matrices [LR94]. matrix [BM99a, Buc95a, CD97a, GJ92, Hor94, Orm90, Wes97, Yan98, vdL94]. maximization [BS90b, KR94, SW94]. Maximum [Hol93, IT98, Lee92, MD98, Ryu93a, Sow92, BSH93, BG98b, CD94, CD97b, Dan94, Des98, GRD99, fL95, Lee97a, Phi91, PS98b, Rui94, SK98, Woo91b, XTZ10]. maximum-likelihood [Phi91]. maximum-likelihood-based [GRD99]. May [Ano90q, Ano91p, Ano92n, Ano93n, Ano94r, Ano95k, Ano96t, Ano98m, Ano99g]. McMillan [HS98a]. Mean [TW91, Bol96, CD94, FHP97a, FHP97b, Hos96a, JHB90, LN99, MS92, RSU96, RU05, Vog99, Woo90, Woo95]. means [Iwa96, Woo91a]. measure [CW97, HY92, Lai93]. measured [LS97a]. Measurement [KWvD97, PM99, AB91, BYN98, Cor99, Era97, Fis95, Jen94, LO91, MWS91, PN96, SS90, SC93, SS97a, SS97b, Sto96, Tin91, WWC97, Weg96, Wei93]. measures [Cow91, Dag90, MZ90, vdB97]. Measuring [AC93, CEK92, Cre90, DFS94, FL90, FK94, Rom97a, Rom97b, Vij93, HMS91]. mechanism [AP93]. Median [jL92]. medical [MN97]. medicine [Gön98]. meeting [Ano90f]. memory [BK96, Bai96, BM96, BCdL98, Chu96, CR96, GD96, HR96, Hos96a, KT96, KLOS97, LR96b, Lob99, TSB96]. men [MW97, TS97]. message [Miz95]. method [ACS99, BG98b, GT99, KP93b, Men94, QS99, SHHS94, Zab93]. Methodological [Wol90]. methodology [Pel96]. methods [Blo96, CD97b, DM92, GH90b, Gon94, HST98, HS96, LW98a, LWG93, kMZ93, PHH91, Ruu91, Sch94]. microeconometric [Era97]. Milton [ZkM99]. minimizing [BW95]. Minimum [VN93, VW93, HS94c, TSB96]. minority [NvPV94]. Mirror [Hal96]. Misclassification [HASM98]. mismeasured [Bol96]. Missing [EKS98, GMP99, CK97, Shi93]. misspecification [Maa90]. Misspecifications [BM93b]. misspecified [Das97b, LN99, Nel92, NF95, SW96, Zab93]. Mixed [Böc98]. mixture [Chi96, MS97, Phi94]. mobility [MZ90]. modal [Chi96]. mode [Hor93, jL93]. Model [CP99, NN98, Abr99, Ahn95, Ai94, Ame90, ACS99, Ano92b, BL91a, BCL92b, BL92, BCL93, BC94, BL95a, BGW91, BKS96, BL98, Buc95b, CT90, CEH90, CP98, Che97, Che99, Chi92, CM97, CK97, CR91, Cre90, DC98, DSW90, EGW90, EKS98, Fri90, FH91, GZW95, GKR97, Gil91, Gil92, Gil93, GH99, GV97, GKW95, Gre90b, GM93b, Hah99, Hal92, Hal94b, HK95, HK98, Hil91, Hor93, IT98, Iwa96, JJ94, Jus95, Kai98, Kar99, KNI96, KDM95, KW91, KS90, Lab99, LS94, Law90, Li99, Lob99, MN95, MS95, MR94, MNS94, NP94, Nel92, NF95, NPL95, NER92, NVvS91, Oht90, OG93, Osi91, Pal95, Per90, Per91, Per96, Phi95a, Poi97a, Poi97b, PW91, Pöt95, RKJ99, Ric95, RS97, SKA94, SK97, SK93, Smi99, Spa90, SS97a, SS97b]. model [Ste95, Tur98, VN93, Vij93, VW93, Weg96, WY97, Wei93, ZWG91]. model-free [HK95]. Modeling [BM96, DG96a, HS98c, Kum97, Spa95, BL95b, BCK92, DH92, DW96, EW98, Fra94, Sim93, TM98]. Modelling [TT93, Lüt97, Urb95]. Models [CCHS90, Mar93, Zel96, Abe94, AP93, AR94, AS95, AL96, AS97a, AM97, AR98, Ana98, AL97, AS97b, AK92, Ang91, Ano92s, AB95, BB92, BGHT95, BOYF94, BMR99, BS94a, BB98, Böc98, BS94b, BS90b, BSH93, Bos94, Bos95a, Bos95b, BDR94, BM98, BR91, BW95, Bro90, Buc95a, BEP96, Cam92, CW97, Cau92, CS98, Cha93, CFJL98, CR97, CF99, CL92, CCN98, CG94, CG95, Chi96, Chi98, CGW98, CK97, Col97, CR96, CKT97, CD97b, Dag94, DD97, DQ93, Des98, Don95, DR97, DK97, Duf93, DK96, ET96, EM92, EH93, Eri95, FOS97, FMB92, Fin93, FLW98, GHT97, GKS91, GH90a, Ghy93, God98a, GMP99, Gön98, GRD99, GM92, GM93a, GM97, GHR94, GD96, GH96, GT92, GRS99, Hah97, Hal93, Ham96, HS93, HRS92, HNIP91]. models [HM94, Hon93, HP94, HKU97, HAD93, Ich93, IASW99, Iwa92a, Iwa92b, JS99b, KCW96, Kim94, KP94, Kiv95, KPS95, KD97, KvD94b, Kle93, KR94, KP93a, KOS95, KPP96, KLOS97, KP97b, KP97c, KM90, Kra93, KR97a, KR97b, Kua98, Kyr98, Lai93, Lau99, LI91, LWG93, Lee94, fL94, fL95, Lee97c, Lee97b, Lee97d, Lee97e, Lee97a, fL99, LY96, LS96b, LH98, Li99, LT99, LH96, MRB98, MM94, MP99, MW99, Mat93, MD98, McA95, McC94, MB90, McM92, Mic99, kMZ93, Mof93, MS97, Mro99, Naw90b, Naw90a, Nel90, Nel92, NF95, Nel96, NRW93, NP93, NV90, Now93, OS93a, OS93b, Paa92, PS90a, PH91, PS90b, PS92, PP93, Pin93, PS98a, Poi94, Poi96, Poi97a, Poi97b, PK94, PN96, PS98b, Qui98b, RS93, Ray91, Rug94, Rui94, Rui97, SS92, SK98, Sch94]. models [Sen90, Seo98, SC93, She94, SK97, SW96, SV99, Sma94b, Smi94, Sow92, SM95, Ste91, Ste92a, Stu95, Sue92, Szr94, Ter98, Tho93, Tsi99, TW91, Var90, VV99, Wan98, Wei91, WC95, Wes96, WA93, Woo91a, Woo95, Woo99, YT98, Zab93, Zel93, vdB97, vdBKOS94]. models-some [Lee97c]. Modified [LB97a, LB97b, MS93, CS90, KP97a]. Moment [KM90, BB98, BQSW99, DD97, Hah97, Kar99, Kem92, KM99, SV99, dL97a]. Moment-based [KM90]. Moments [US94, ACS99, Fri90, GHT90, GT99, HT99b, Kin91, McC98, MG91, OK96, QS99]. monetary [BS90a]. money [FK93, KO90, KO93, Lew90, SR91, Tho94]. money-income [FK93, Tho94]. money-output [KO90, KO93]. monotonic [BL92, CS98]. monotonicity [RD99]. Monte [ACS99, BCL92a, BCL92b, BCL93, BKS96, Buc95a, CCN98, DM92, GLN94, Hau96, Lee97c, Lee97b, Lee97a, SK98, SV99, TM98, Tsi99, Vij97]. moral [DGGV97]. Morrison [HY92]. mortality [BSTY91, fLRP97]. movements [Lew90]. moving [CS90, Fit98, GZW99, KT96, PH91]. moving-average [CS90]. MSE [OK96, VN93]. much [CD98]. Multi [Wei91, EW98, KO90, KO93, NER92, SS90]. multi-category [EW98]. multi-country [KO90, KO93]. multi-dynamic-factor [NER92]. multi-product [SS90]. Multi-step [Wei91]. multidimensional [BHHS94, DKF98]. multilateral [SR92]. multinomial [Ano92b, CR91, DSW90, GKR97, GJP97, MR94]. multiperiod [GKR97]. Multiple [BFH97, Gre90a, MNS94, AV98, Bai99, Chi98, CA99, CD93, Hil91, Hos96b, Iwa93, Rei90, Rob91]. multiplier [TLL^{+}90]. multivariate [BSH93, BDR94, DH92, Fra94, HMR96, JJ92, JHB90, Jus95, KNI96, KPP96, Kyr98, Lob99, MM94, Nel96, NS96, Now93, SS92, SR91, Vij97]. must [Maa90]. mutual [KNI96].

natural [CR97, KDM95, KP93a]. near [HM94, Per96]. near-integrated [HM94, Per96]. nearly [NP97, RS97]. necessarily [SS97a, SS97b]. negative [Kin91]. neglected [LWG93]. Nelson [Luc95]. Nelson-Plosser [Luc95]. nested [BS90b, God98a, KR94, McA95, Mic99, Poi96]. Netherlands [NvPV94]. network [LWG93]. networks [FD97, WW96]. neural [FD97, LWG93]. neutrality [EC96]. No [Ano92b, BCL93, Gil93, KO93]. Noise [Szp97a, Szp97b, CD94]. noisy [Tyb92]. nominal [Cro98]. Non [Vel99, Ana98, CK97, God98a, Han99, KP97a, McA95, Mic99, kMZ93, RM92, SM95, Szp97a, Szp97b, TM98, US94, Zel98]. non-Bayesian [Ana98, kMZ93, Zel98]. non-dynamic [Han99]. non-Gaussian [RM92, TM98]. non-linear [CK97, Szp97a, Szp97b]. non-nested [God98a, McA95, Mic99]. non-normal [SM95, US94]. Non-stationary [Vel99, KP97a]. non-union [LW98a]. noncausality [BDR94]. Nonclassical [HK95]. Nonconvexities [Chi95]. nonexperimental [Spa95]. nonfundamental [LR94]. noninvertible [RM92]. Nonlinear [GIM97, HNP95, PST96, TM98, AM97, Ame90, AV98, Bie97b, Blo96, CW97, DSK99, KCH95, KP96, KPP96, RSU96, RU05, Woo99]. nonlinearities [BEP96]. nonlinearity [BGH^{+}97, KP99, LWG93, dL97a]. nonmonotonic [Vog99]. nonnegative [BP92]. nonnegativity [CL92, VK90]. Nonnested [SK93, Cha93, GH90b, PP93, TW91, Woo90, Zab93]. nonnormal [RS97]. nonnormality [AS93, AS96, Eva92]. Nonparametric [Ahn95, BGHT95, Bie97a, FG95, Goz97, HM93, Hus92, LW90, Lüt97, Mat93, RCCA93, SK96, AM93, AP93, BHL96, BHF90, CR97, CF99, Eas91, Ghy97, HT97, Lew95, MS92, Pin98, WL99, Zhe96]. nonregular [LB97a, LB97b]. nonresponse [HM98]. Nonsense [Ent97]. nonspherical [FMB92]. nonstationary [AR94, CP99, DK91, GAR97, KNI96, Mar98, Phi98]. nonsynchronous [LM90]. Normal [LR96a, ALP96, Cau92, HMR96, JHB90, Phi94, SM95, SS97a, SS97b, US94, Vij97]. normalising [Yan98]. normality [Eas91, GS91, New97]. Normative [Sle91]. North [Ano90f]. note [CS90, GKS91, Had96, Iwa93, Kin91, NPL95, Osi91, XTZ10]. November [Ano90r, Ano98p, Ano96q, Ano97u, Ano98n, Ano99k]. nuisance [CKT97, RK97]. null [CA99, KPSS92, LMN98, Run97a, Run97b]. number [KM99]. numbers [Sle91, dJ98]. numerical [Dor95, PHH91, Ste91, Vin98]. nutrition [fLRP97, Sch97]. Nutritional [Bha97b, Das97a]. observations [BCL92a, CK97, GMP99, NV92, Shi93]. observed [Ai94]. Occupational [BHHS94, DFS94]. occupations [HS94b]. October [Ano91o, Ano94q, Ano90r, Ano92o, Ano93o, Ano95r, Ano96v, Ano97-27, Ano98q, Ano99m]. odds [Con91]. OLS [BM99a, KPS95, PK96]. one [BC94, DS98]. one-sided [DS98]. one-way [BC94]. operators [BK98]. opinion [Ano92h, Bar97, Gra92, Maa92]. optima [MNS94]. Optimal [ALP96, BG98a, DK91, LI95, PS96b, Bro93, NV92, Paa97a, Paa97b]. optimality [And91]. optimization [GFR94]. optimizing [Var90]. option [AN98]. options [DL92, EM92, MT90, Sch96]. order [Bie93, CS90, DK96, LS97b, MS93, RSU96, RU05, SL97, XP98]. ordered [BM98, GJP97, jL92, SCS99, Szr94]. Ordering [EMS99, Wil96]. Ordinal [vP91]. ordinary [FMB92]. Orthogonality [Hon93, CKT97, PS90b]. OTC [Gön98]. other [AKS92, GAR97, TW91]. outcome [Mro99]. outcomes [HM98, MS97]. Outlier [FKL98, GMP99, Luc95]. Outliers [AKS97, Rei90]. output [Dur93, KO90, KO93, PPL93, ZHkM91, ZkM99]. Overdispersion [GT92, CT90]. overidentification [AK92, AK94]. overidentifying [MS96]. overlapping [Yit94].

p [CG94, Dua97]. Pages [Ano90x, Ano90q, Ano90v, Ano90u, Ano90t, Ano90s, Ano90w, Ano90r, Ano91p, Ano91s, Ano91r, Ano91q, Ano91n, Ano91t, Ano91o, Ano91u, Ano92n, Ano92r, Ano92m, Ano92p, Ano92q, Ano92o, Ano93p, Ano93q, Ano93l, Ano93m, Ano93k, Ano93n, Ano93o, Ano93j, Ano94s, Ano94t, Ano94o, Ano94u, Ano94p, Ano94n, Ano94r, Ano94q, Ano95p, Ano95q, Ano95m, Ano95o, Ano95l, Ano95k, Ano95n, Ano95r, Ano96y, Ano96w, Ano96x, Ano96r, Ano96s, Ano96u, Ano96t, Ano96q, Ano96v, Ano96p, Ano97-29, Ano97z, Ano97-28, Ano97x, Ano97y, Ano97r, Ano97v, Ano97u, Ano97-27, Ano97s, Ano97w, Ano97t, Ano98s, Ano98u, Ano98r, Ano98t, Ano98l, Ano98m, Ano98o, Ano98n, Ano98p, Ano98q, Ano98k, Ano99n, Ano99o, Ano99p, Ano99j, Ano99l, Ano99i, Ano99g, Ano99k, Ano99m, Ano99h]. paired [BG98a]. Pairwise [HP94, HKU97]. Pakistan [BFR97]. Panel [Bal95, PK94, AS95, AS97a, Are93, AC93, BCL92a, BHL96, BL98, BBDS92, BB98, CGW98, CK97, CSW92, DL95, DGV98, FOS97, Fre95, GS92, Gön98, GM93a, Hah97, Hah99, HAD93, IASW99, Kao99, Kiv95, Kum90, KP93b, LS96b, LH98, MN95, MN97, Met96, NVvS91, PSS98, VV99, Woo95, Woo99]. panels [BC94, CPW93, EKS98, Han99, HT99a, LI95, NV90, PS95a]. paper [Hin97]. paradigms [Paa92]. Parameter [Dag94, Koo96, And98, CD94, CCN98, CG95, FLW98, Kuo98, LT99, SS97a, SS97b, Vij93, Weg96, ZHkM91]. Parameters [CKT97, Bos95b, Cor99, DGV98, EKS98, LT94, LT99, NV90, OP97, RK97]. Parametric [AH98, BHF90, BH97, LW98a, LW98b, MS92, McM92, SW96, WA93]. parity [Jus95]. Park [CS90]. part [LY96, OK96]. Partial [Urb95, Joh92, Kuo98, MNS94, RT99]. Partially [Phi94, AR94, AM97, BM99a, CP99, CP92, CK97, LS96b]. participation [Ste96]. particularities [KDM95]. partly [Iwa94]. pass [Yik98]. patterns [BT98, Can93, DC98, FD97, Gew93]. pay [WW96]. payoffs [GHT90]. penetration [BT98]. perceptions [HK98]. performance [BM95, FG96, GS92, Orm90]. period [Cre90]. periodic [BFH97]. periodically [Osb91]. periodogram [LR96b, Vel99]. Persistence [PPL93, DG96a, FH90, KLOS97]. Personal [FGM90, BGR95]. perspective [HKT90, vdBKOS94]. Perspectives [Ano90f]. phase [CEH90]. phase-averaging [CEH90]. Phillips [XTZ10]. philosophy [Sha96]. Pierce [Run97a, Run97b]. Pitfalls [GL98]. pivot [Vin98]. plan [HK98]. Plosser [Luc95]. point [AF92, Bro93, Chi98, KP96]. point-optimal [Bro93]. points [ZHkM91, ZkM99]. Poisson [Böc98, CT90, GT92]. policy [KP93b, RW91]. political [FH90]. political-business [FH90]. polychotomous [fL95, Mat93, Sch94]. Polyhedral [CCHS90]. Polynomial [GL94, AF92, HT97, HNIP91, Kua98]. poolability [BHL96]. Pooled [BG97a, AL96, MG91]. Pooling [Ano92b, CR91, ZHkM91]. Population [PW95]. populations [fLRP97, MWS91]. positioning [BG98a]. positive [MB90, OK96, Pel96]. positive-part [OK96]. possible [ÖJ91]. possibly [KP97a, LN99, SW96, TY95, YT98]. Posterior [CP98, CGW98, Lan97, Chi96, Con91, Iwa96]. potential [Kem92]. Pötscher [Ste95]. poverty [Das97a, FH90, VB90, FH90]. Power [SG91, Bar92, Cor99, DNSW92, HS96, Jus95, KZ90, LS96a, NT90, PS95c, PS96c, Vog99]. powers [Sma94a]. pp [Ano92b, BCL93, Gil93, Hal94b, KO93]. PPP [JJ92]. Pre [Gil91, Gil93, NV92, ÖJ91, ST95]. pre-experimental [NV92]. pre-test [ÖJ91, ST95]. Pre-testing [Gil91, Gil93]. predeterminedness [AK92, AK94]. Prediction [BB92, KOS95, Mil92, Pal95, Phi95a, Phi95b, Ric95]. Predictive [GGH98, Iwa92a, WC95]. preference [HK98, HLS98, OP97, Sen91, SW94]. preferences [HK98]. preliminary [Gil92]. preliminary-test [Gil92]. premium [MS92]. presence [CEH96, DT94, God98b, Han92a, LMN98, LN99, RK97, Wei90]. present [Koo91]. pretest [Hal92, Hal94b]. Price [Dea90, AH93, BGR95, Bie97b, DHSR98, EKS98, Fis95, Gön98, HW96, HK95, MD98, Mou91, Paa97a, Paa97b, Sle90]. price-setting [MD98]. prices [CN96, Dan94, EM92, Koo91, MN95, VB90]. Pricing [MT90, AN98, BM96, ENR90, GH90a, MR91, Sha90, Smi99, Stu95]. principal [PM99]. principles [Mag90]. prior [CP98, Poi94, Zel96]. Priors [KCW96, FOS97, KP93a, RS93, YP96]. Private [GS97a, Paa92]. Probabilities [Car95, Ahn95, HMR96, Vij97]. probability [BSH93, DSW90, Sha96]. probit [BL98, BS94b, GKR97, GM93b, MR94, SV99]. problem [BL91a, PP93, Ril93]. problems [DNSW92]. procedure [DKF98]. procedures [CEH90, DC96, Hil91]. process [CS90, Chu96, Dua97]. processes [BK96, Bai96, BP92, CP99, Dor95, EG93, GZW99, GS97b, HT99b, Kuo98, LC97, LS97b, Mar98, NS96, Osb91, RM92, SL97, She94, TSB96, TY95, dV91]. procyclical [Chi95]. Product [KVC98, BG98a, BT98, KS97, NPL95, SS90]. Production [CSS90, Kum90, BGW91, BFR97, BHF90, BFR95, BW95, Fis95, FK94, Hal93, Kra93, PP98, Tyb92]. production-smoothing [Hal93, Kra93]. production-theoretic [Fis95]. productive [Sch97]. productivity [AB91, Chi95, DSK99, GR95, HM95, LO91, PN96, Rei90]. profiles [FBA91]. programming [FL90, ST90]. promotion [FLW98]. promotions [DHSR98]. Properties [FMB92, HT99b, Dur93, Hod92, Koo91, LS94, Lau99, PS98b, Smi99, SM95, WY97, Zel98]. property [BL92]. proportional [Rug94, Sue92]. prospects [Ano90f]. proxy [OG93]. Pseudo [BG98b, CPW93, CD97b, Lai93, Maa92]. pseudo- [Lai93]. Pseudo-maximum [BG98b, CD97b]. pseudo-science [Maa92]. public [GS97a]. purchase [MR92]. purchasing [Jus95]. purpose [BT95].

q [CG94, Dua97]. QML [AS97b, Rui97]. Quadratic [jL93, JHB90, Kem91, US94]. quadrature [GH90b]. quadrature-based [GH90b]. Qualitative [FG96, GM92, BS94a, Lai93]. quality [HMS91, fLRP97, Mou95]. Quantile [Buc95b, Buc95a, Fit98, KP96]. quantiles [Kle93]. quantity [Fis95, HK95]. Quasi [Cra92, NF98, Rui94, Hos96b, McC98, Woo91b]. Quasi-Aitken [Cra92]. quasi-FIML [McC98]. quasi-likelihood [Hos96b]. Quasi-maximum [Rui94]. quasi-maximum-likelihood [Woo91b]. Quasi-rational [NF98]. quasilikelihood [SC93].

R&D [HM95, Rei90]. radio [Mou91]. Random [Ent97, BL92, BW95, Che99, EGW90, FBA91, GKS91, GM93b, HSPHS90, Hol95, IT98, MB90, PP94, Run97a, Run97b]. random-coefficient [GKS91]. random-effects [EGW90]. range [CR94a, Hos96b]. Rank [BG97b, CS98, CP99, CD97a, Gew96, Ino99, Kar17, Seo98]. rate [AL97, Bie97b, Cro98, HKT90, Jus95, KDM95, PS92, PST96, WC95, vdB97]. rate-making [HKT90]. rates [GHT90, kMZ93, New97, SvD91, ZHkM91, ZkM99]. Ratio [CCHS90, Att95, Bai99, CD93, PL91, SG91, US94]. Rational [Cro98, Sim93, HS93, JS99b, Lab99, NF98, PS92, Ray91, SS92]. rational-expectations [Ray91]. rationality [EC96]. rationing [Squ94]. real [Gri96, SvD91]. Reasonable [GU90]. Recognizing [FD97]. rectangle [HMR96]. recursion [Zha99]. recursive [Ste91]. reduced [CP99, Gew96, Kar17, McC98]. reduction [KP94]. Redundancy [BQSW99, BK98]. reference [GM93a, HS98b]. reformulation [AL96]. regime [Ai94, GH96, Ham90, HS94a, KP97b, KP97c]. Regression [CT90, DM92, MB90, AL96, AS96, And91, ALP96, BC94, Böc98, Bro93, Bro90, Buc95a, Buc95b, Cam92, CW97, Cau92, Chi92, Chi93, CG94, CP93, Dag94, DD97, DC98, Eas91, EH93, Ent97, FMB92, FH91, GKS91, Gew96, Gil91, Gil92, Gil93, God98a, GH99, GT92, GRS99, Hal96, HP94, HS94c, Iwa92a, Kao99, Kar17, KCH95, KW91, KPS95, KP96, KP97b, KP97c, KZ90, Kyr98, jL92, jL93, Lee97c, Lee97b, LW98b, LT94, MRB98, Mar98, Mic99, Mou91, NT90, Naw90b, Naw90a, NPL95, Oht90, OG93, Osi91, OS93a, OS93b, Pöt95, RK97, Rob91, RCCA93, Ryu93a, Shi93, SKA94, SK97, SK93, SK96, Soo90, Ste92a, Ste95, Vel99, VN93, Wei93, mWK96, Woo91a, XP98, ZWG91]. Regression-based [CT90, DM92, Woo91a]. regressions [Abe94, BCL92a, Bol96, Cho94, DK91, Fit98, Hal94a, Hod98, Kuo98, LM97, MMF93, Phi91, XTZ10]. regressor [Bol96]. regressors [Cho94, HM98, KP97a, Mic99, Ril93]. regularity [FG95, vSKK93]. regulation [AB91]. reinterpretation [MS96]. Rejection [KB95]. rejections [LMN98]. related [KW91, NP94]. relation [dV91]. relations [Dav98, HK95]. relationship [BFR97, Dag90, HM95, Kum97]. relationships [Gon94, GL98, GNW96, Ind93, KNI96, Koo91, PS95a]. Relative [MMF99, GT99, Sle90, Szr94]. rents [LO91]. repeated [Mof93]. Reply [Ste95, Bos95a]. Representation [BYN98]. Representations [HS98a, LR94, ST98]. representative [PW95]. Research [CD97b, Ano92c]. reserve [Paa97a, Paa97b]. residential [FBA91]. Residual [GH96, Leh90, Cho94, Kao99]. Residual-based [GH96, Cho94, Kao99]. residuals [PK96]. resistant [PK96]. resource [CR97]. Response [Rui97, CR97, DL97b, GJP97, HM93, HASM98, HS98c, KPP96, Koo96, jL92, LS97b, Phi95b, Phi98, ZkM99]. responses [BM93b]. restricted [And98]. Restricting [FFM96]. restrictions [BB98, Gil91, Gil93, Hah97, Joh95, Kem91, Kem92, MS96, OG93, RR90]. results [BCL92b, BCL93, Dav98, Dea90, God96, God98b, God98a, HMR96, Lee97c, Lee97b, Lee97a, McC98, ÖJ91]. retirement [MW97]. returns [CL96, CEK92, DG96a, NER92, Run97a, Run97b]. revealed [DKF98, HK98, SW94]. Review [BYN98, BCK92]. revisited [Leh90, ORS96, Spa90]. Ridder [Ano92b]. right [Nel92, NF95]. risk [CEK92, Leh90, MS92, ÖJ91]. risk-premium [MS92]. risks [Sue92]. Robust [KR97a, KR97b, Naw90b, Naw90a, OS93b, CA91, Fit98, FKL98, HLvD95, Luc95, Vin98, Woo91a]. Robustness [AS93, AS96, Eva92, AK94, SG91, dL97a]. role [MD98, PS95b, TLL^{+}90]. root [Bie93, Bie97b, Bos94, CP93, DNSW92, Die93, GP93, GAR97, GS97b, Hal92, Hal94b, HLvD95, HS96, KCW96, KS93, KPSS92, Luc95, NT90, PS95c, SvD91, Seo99, ST98]. roots [BM93a, BFH97, BG97b, Dic93, FHP97a, FHP97b, GLN94, Gre90a, HT99a, Hyl95, Iwa93, Lub95, PH91, Saï91, TP93, Vin98]. rotating [NVvS91]. rotating-panel [NVvS91]. Roy [Vij93]. Ruiz [AS97b]. rule [Eas91, OK96, Oht99]. Rules [Maa92, fL94]. run [DHSR98, FKL98, Gon94, GL98, Ind93, JJ94, Jus95, KOS95, PS95a]. Rwandese [Bha97b].

S [Ano92b, Hin97]. Saddlepath [SS92]. safety [DGGV97]. sales [Paa97a, Paa97b]. Sample [Ter98, Bie93, BK98, CA91, CP92, Don95, God98a, GS92, Hod92, Hos96a, KZ90, LS94, Lee94, fL94, LY96, McC98, MG91, Orm90, PS98b, Sch94, Smi99, Ste92b, Szr94, YT98, Zel98]. samples [Lan97]. sampling [BW93, Chi93, GS95, GW93, IL96, MRB98, SV99, WWC97, Wes96]. sanitation [fLRP97]. Sargan [Had96]. Sato [RD99]. Sato-Vartia [RD99]. scalar [HST98]. scale [Mai90, She94]. scales [BL91b, Pen99]. scaling [DKF98]. scanner [DHSR98, EKS98, Gön98]. scanning [FKL98]. scarcity [LO91]. science [Maa92]. scope [RW91]. score [JT94, MD98, RK97]. score-based [RK97]. search [Blo97, CK97, DC98, RS93, TT93]. Seasonal [BM93a, Dic93, HEGY90, Osb93, RS93, Zel93, Abe94, AR94, Can93, Die93, EC96, Fra94, FHP97a, FHP97b, FD97, Gew93, GP93, GLN94, GS95, Hyl95, JS99a, Lee92, Osb91, ST98]. Seasonality [Ghy93, HS93, Duf93, Ghy97, Hal93, Kra93, NRW93]. seasonally [Sim93]. second [Kar99, RSU96, RU05]. second-order [RSU96, RU05]. section [AL96, Ano92h]. sectional [Con99, CSS90, Fre95]. sections [Col97, Mof93, NV90, PW95, VN93]. securities [BM99b]. seek [Era97]. Seemingly [MMF93, KPS95]. segregation [BHHS94, DFS94, NS94]. selecting [SW96, TW91]. Selection [Woo95, AP93, CP99, Don95, Era97, EKS98, GKW95, Hal92, Hal94b, KS90, Lee94, fL94, LY96, Mou91, NP93, Pal95, Phi95a, PW91, Ric95, Sch94, SK97, SK96, Ter98, VV99, VW93]. Selectivity [Sch94, Vij93]. Self [Era97, EKS98]. Self-selection [Era97, EKS98]. Semi [CR97, Eas91, LW98a]. Semi-nonparametric [CR97, Eas91]. semi-parametric [LW98a]. Seminonparametric [BGW91]. Semiparametric [AP93, Che97, CR94a, GH99, GRS99, Hor93, Ich93, Lee94, fL94, fL95, LS96b, Pen99, SC93, Ste96, Sue92, Ai94, CF99, Das91, DK97, HM93, LH98, Lob99, NP93, PO95, PSS98, PS91, Pin93, Ril93, Rug94, SKA94, Tho93, WA93, mWK96]. sensitive [HT93]. sensitivity [BM99a]. separability [DW95]. separable [CM97, SW94]. separate [JT94]. September [Ano92p, Ano91u, Ano93j, Ano94q, Ano96p, Ano97s, Ano98k, Ano99h]. sequential [Hah97, fL95]. serial [Böc98, CG95, Kyr98, LH98, Pin98, Rob91, Wei90]. series [AL96, AKS97, BL95b, BW93, BP92, Can93, CF99, CR94a, CA99, Col97, CSS90, DQ93, DeJ92, DNSW92, DH92, Dur93, Fin93, Fra94, FK94, Gew93, GLN94, GP96, GAR97, GW93, Hal92, Hal94b, Ham90, Ham96, HRS92, Hos96a, Hos96b, Hus92, Jus95, KNI96, KOS95, KP99, KPSS92, Lau99, LI91, LWG93, Li99, LH96, MM94, Mof93, NT90, New97, Osb91, PRW97, Pöt95, Saï91, Shi93, SK97, Sow92, Ste92a, Ste95, Szp97a, Szp97b, VN93, Vog99, WL99, XP98]. service [BHF90, HAD93]. services [Mou91]. set [CCN98]. sets [FH91, HU94]. setting [HASM98, MD98]. settings [Man93]. several [BCL92b, BCL93, Thu92, WC95]. sex [MS97]. sexes [MZ90]. share [FFM96, Kum97, MMF93]. shares [FFM96]. shelter [Mou95]. shift [Thu92, Vog99]. shifts [FHP97a, FHP97b, GH96]. shocks [AKS97, Chi95]. short [AL97, JJ94]. short-run [JJ94]. short-term [AL97]. Shrinkage [KCH95, KB99, Vin95]. sided [DS98]. significance [BHHS94, KM95a, KM95b, McA95, Sha96]. Simple [And94, Ang91, CD93, Cre90, Dav98, LW98b, Miz95]. simplex [FFM96]. simplification [Men94]. Simplified [BDR94, AS97a]. Simulated [BM98, Lee97a, CD97b, Dan94, GFR94]. Simulation [GM93a, HMR96, LI91, Lee97c, Lee97b, CGW98, PP93, Van92]. Simulation-based [GM93a]. simulations [LR96a, TM98]. simulator [Lee97d, Lee97e]. simulators [BSH93]. Simultaneous [CSW92, Lil93, RR90, AK92, BS94a, CP98, Dhr94, Fri90, Joh95, Kai98, KR97a, KR97b, Lee94, Mro99, ST95, Smi94, Spa90, Ste91, Tur98, Zel98]. simultaneous-equations [Spa90]. sing [Tyb92]. single [BGH^{+}97, Hal94a, Hod92, Ich93, Joh92, KD97, PS90b]. single-blind [BGH^{+}97]. single-equation [Hal94a, Hod92, Joh92, PS90b]. single-index [Ich93]. situation [Pol91]. size [BK98, DSW90, DC96, Eva92, HSPHS90, MS95, Mai90, Wol90]. size-efficient [Mai90]. skill [CL96]. Skipping [GMP99]. slope [PK94, SS97a, SS97b]. slopes [Zab93]. slowly [HS98b]. SLS [Ich93, Ich93]. Slutsky [Lew95]. Small [KW91, God98a, Orm90]. Small-disturbance [KW91]. small-sample [Orm90]. Smith [HS98a]. Smooth [BSH93, ET96, Lee97d, Lee97e, PO95]. Smoothing [Sto96, FK94, Hal93, Kra93]. smoothness [FK94]. SNP [FG96]. social [BR97, Dag90, Man93]. Society [Ano90f]. Socioeconomic [KWvD97]. solutions [Ana98, SS92]. Some [EG93, GLN94, God96, ÖJ91, Tho93, Wei93, AM97, BP92, CW97, Dav98, God98b, God98a, HNP95, Lee97c, Lee97b, Lee97a, MS95, Men94, NP93, RS93, Sma94a, Woo99, YT98]. Sørensen [Rui97]. Sources [PP98, Vog99, HLS98]. space [And98, DH92, PS98a, She94, TM98]. spatial [BK98, DKF98, PS98a]. special [GM93a, HS98b]. specific [Hyl95, PK94]. Specification [Dhr94, Ham96, Kle93, LH96, McC94, Met96, Woo91b, BW95, DT94, FFM96, JM91, Li99, MP99, NN98, PS91, WA93]. specifications [GKS91]. specified [AM97]. Spectral [Dur91, HST98]. speculative [DG96a]. speed [PS96a]. spherically [Gil91, Gil92, Gil93, OG93]. spreads [MN95]. Spurious [Cho94, Kao99, LMN98, Mar98, Abe94, Ent97, TP93]. square [HS94c, VW93]. squared [CW97, Kar99, RSU96, RU05, TW91]. squares [FMB92, Fit98, Ich93, JM91, LH96, MS93, Sch90, SM95, Szr94, Yik98]. Stability [Qui98b, Bro93, DC96, Mil92, ORS96]. stable [KT96, Tsi99, dV91]. stage [Ahn95, Ame90, BS94b, fL94, MS93, ÖJ91, Sch90]. standard [AS97b, JS90]. State [She94, DH92, JT94, Slo90, TM98]. state-space [TM98]. statement [Aig96]. states [Ano92b, CR91]. static [AH98]. Stationarity [BP92, Bie97b, CA99, DeJ92, KPSS92, LS96a, RKJ99]. stationary [DK91, GHR94, KP97a, LT99, Seo98, Sow92, Vel99, ZWG91]. statistic [HM94, PP93, Run97a, Run97b]. Statistical [Ano92s, GKR97, MWS91, Seo98, Spa90, TY95, GFR94, Hos96b, HKT90, KWvD97, Lee97a, Vin98]. statistics [AKS92, CA91, DFS94, GZW99, MS95, Phi97, PS98a]. status [Bha97b, Das97a, MWS91]. Steigerwald [Pöt95]. Stein [HU94, OK96, Oht99]. Stein-rule [OK96, Oht99]. step [Don95, Lob99, VV99, Wei91]. Stochastic [BW95, Dan94, Mar93, PSS98, vdBKOS94, AL97, AS97b, ACS99, BCdL98, FOS97, FFM96, GHT97, GS92, Gön98, GS97b, GIM97, Gre90b, HS98b, KM90, KO90, KO93, LT99, MT90, Rui94, Rui97, SK98, Sen90, SK97, WL99, Wic96, dJ98]. Stock [DL92, BM96, BM99b, CEK92, Koo91, LI95, MN95, NER92, PS90a, Run97a, Run97b]. stock/flow [LI95]. Stratified [RT99, IL96]. strictly [IASW99]. strong [Rob91]. Structural [Dav98, Ryu93b, Bai99, BGHT95, BS94b, Bos94, Bos95a, Bos95b, CEH96, Chi95, CP92, CD98, Das91, DC96, DG96b, DK96, Eri95, GH90a, GP96, GGH98, GNW96, HR96, Hod92, JJ92, LT94, Mil92, PK96, PS96c, SK97, Smi94, Thu92, TM93, Zha99]. Structure [KvD94a, BFR95, Buc95b, CP99, DC98, ENR90, JJ94, PST96, PP94]. studies [Kle98, LI96]. study [ACS99, BC94, Buc95a, Cha90, Kum96, MN95, MM94, TP93]. Subject [AC94, AC97, Ano92t, Ano94v, Ham90, fL94]. suboptimally [Szr94]. Subsample [Tho94]. Subsampling [PRW97]. subsidies [DR93]. substitution [BMN93, BT98, GHT90]. successive [Wes96]. Summary [DFS94]. superexogeneity [EH93, PS96c]. supply [Ang91, Blo97, Blo96, BMN93, DL97b, Law90, NF98, Ste96]. sure [KPSS92, Per91, SM95]. surprise [HU94]. surprises [MN95]. survey [DL95, Dea90, GM93a, HM98, HS98c, KS97, MRB98, NRW93, PW95]. surveys [Rom97a, Rom97b]. survival [JM91]. switching [BMR99, Ham96, Kim94, KP97b, KP97c, Lee97c, Lee97b, Phi91, PS98b, Ter98, XTZ10]. symmetric [Gil91, Gil92, Gil93, HKU97, OG93, Tsi99]. symmetry [Lew95]. System [Yan98, Att91, Att95, Urb95, VK90]. Systematic [GS95]. systems [AV98, Cha90, Cha98, Dhr94, DR97, HS98a, HS90, Joh92, MR92, NP97, Par96, PR99, Urb95, WW96, Woo96, vSKK93].

take [Blo96]. target [PS92]. tax [Era97]. taxes [Blo96]. tech [BM95]. technical [AC93, Kum90]. techniques [Mou95, Van92, ZHkM91, Zhe96]. technological [DGV98]. technologies [BT95, Tyb92]. technology [Chi95, MN97, Phi97]. teflon [KS91]. telecommunication [HW96]. telephone [HAD93]. television [WW96]. Temporal [PS95c, GS95]. term [AL97, BM99b, PST96]. Test [Kuo98, AL96, AS93, AKS92, Att95, BHL96, BS90a, Bar92, CN96, CD93, DNSW92, Dor95, Eas91, Gil92, GL94, Hor94, LS96a, LW98b, Luc95, MS95, Orm90, ÖJ91, Pin98, PK96, Saï91, ST95, SW94, Zhe96, dJ96]. Testing [AV98, BL95a, Bie97b, Bos94, CP93, CA99, Das97b, DS98, DT94, DC96, ET96, EH93, EC96, GH90b, GLN94, GAR97, GNW96, GJ92, Gri96, Hal92, Hal94b, HS98b, HR96, JJ92, JS99b, KPSS92, Kyr98, LWG93, LH98, LT94, LT99, Lub95, LN99, OG93, PH91, Rob91, SL97, Shi93, SKA94, Sne99, And98, Ang91, Are93, Att91, Bou95, Cha93, CF99, Dur91, Gil91, Gil93, GL98, GKW95, GM93b, Ham96, Han92a, Han99, HK91, Iwa94, KM95a, KB95, KvD94b, KM90, Mag90, McA95, MR91, Met96, Mir95, Oht90, PP93, SK93, Spa95, Vog99, Woo90, Woo91b]. Tests [AK92, Can98, God98a, Hau96, Hyl95, Ino99, Kua98, WA93, AS96, AK94, And94, ALP96, AH98, Bai99, BCL92b, BCL93, BGH^{+}97, BD96, BHHS94, BG97b, Bro93, CT90, CEH96, CP92, Cho94, Dhr94, Die93, DW95, DK91, DK96, Eva92, FG95, GP93, GGH98, God96, God98b, GH96, GT92, HS96, JT94, JM91, Kao99, Kem91, KM95b, KS93, KW91, KD97, Kle93, Koo91, KZ90, LWG93, LT96, Lew95, LMN98, Li99, LB97a, LB97b, MS96, Mic99, Mil92, NT90, Pen99, PS90b, PS91, PS95c, PS96c, Qui98b, RK97, Seo99, SG91, SV99, Sma94a, ST98, Smi99, Thu92, TM93, VW93, Woo90, Zab93, dL97a]. Theil [SR92]. their [BG97a, BM93b, Dav98, HMR96, LR96a]. theorem [PO95]. theoretic [Fis95]. theoretical [GLN94, HMR96]. theories [Bou95, GKW95]. Theory [MMF93, PN96, AM93, Blo97, BCK92, CP92, CD98, Gra92, GS95, KB95, LS94, Mar98, MR91, Nel96, PS95b, Seo99, Spa95]. think [Mir95]. third [MS93, Poi97a, Poi97b]. third-order [MS93]. threaten [BM97]. Three [Mir95, Sch90, BYN98, DL95, MS93]. three-dimensional [DL95]. three-facet [BYN98]. Three-stage [Sch90, MS93]. Threshold [Han99, GM92]. thumb [Maa92]. Tiao [BOYF94]. timber [Paa97a, Paa97b]. Time [Dur93, Lau99, AL96, AL97, AKS97, BB92, BL95b, BW93, Bha97b, BH97, BP92, Can93, Cha93, CF99, CR94a, CG95, CA99, Col97, CR96, Com99, CSS90, DQ93, DeJ92, DNSW92, DR97, DH92, DW96, Fin93, Fra94, FK94, Gew93, GLN94, GP96, GAR97, GW93, Gri96, HW96, Hal92, Hal94b, Ham90, Ham96, HT99a, HRS92, Hos96a, Hos96b, Hus92, Jus95, KNI96, KOS95, KP99, Kum90, KP93b, KPSS92, LI91, LWG93, LW90, LS97a, Li99, LH96, MM94, Mof93, NT90, NV90, Osb91, PRW97, Pöt95, RS93, Saï91, Shi93, SK97, Sow92, Ste92a, Ste95, Sue92, Szp97a, Szp97b, Vog99, WL99, XP98, ZHkM91]. time-dependent [BB92, NV90]. time-series [CSS90, Ham96, Hos96b, Jus95, LI91, NT90, Osb91, Saï91]. time-varying [CG95, DR97, HW96, Kum90, LW90, Sue92, ZHkM91]. timing [KP93b, Lil93]. tobacco [Lab99]. Tobin [BBDS92]. Tobit [Che97, Chi92, Gre90a, Hon93, HKU97, Iwa93, fL94, fL99, Per90, SV99]. Tobit-type [fL94]. today [Sha96]. Tornqvist [SR92]. trading [LM90, MN95]. training [Gri93]. transformation [Abr99, BL91a, Buc95b, GH99, KCH95]. Transformations [Sen90]. Transforming [GZW95]. transition [ET96, GM97]. transitions [EGW90, MW97]. translog [Kum97, VK90]. traps [Das97a]. treasury [ENR90]. treatment [Ter98]. Trend [RKJ99, Bie97b, DeJ92, Hal96, Ino99, Kua98, Per97, PK96, Saï91]. trend-break [Ino99]. trend-resistant [PK96]. trend-stationarity [DeJ92]. trending [KM97]. trends [GIM97, Han92a, KO90, KO93, Mar93, Wic96]. Tribute [Zel91]. trimming [HKU97]. trip [Hor93]. truncated [GT92, HP94, TM93]. truncation [Eas91]. turning [ZHkM91, ZkM99]. Two [Ame90, Don95, RS96, VV99, Ahn95, BL92, BS94b, Cre90, FH91, fL94, LY96, Lob99, ÖJ91, Poi97a, Poi97b, vP91]. two-part [LY96]. two-period [Cre90]. Two-stage [Ame90, Ahn95, BS94b, fL94, ÖJ91]. Two-step [Don95, VV99, Lob99]. two-way [BL92]. type [fL94, Che97, HKU97]. Type-3 [Che97].

U.S. [AH93, BM93a, BM95, Buc95b, Dic93, PPL93, PN96]. UIP [JJ92]. UK [JJ92, MW97, SR91]. ultrastructural [SS97a, SS97b]. unbalanced [BC94, NP97]. unbiased [BSH93]. unbiasedness [KS90]. uncertainty [AM93, Ahn95, Koo96]. uncovered [Jus95]. unemployment [Blo97, FGM90, NvPV94]. unidentified [Vij93]. unified [PS90b]. Uniform [dJ98, PP94]. union [LW98a]. union/non [LW98a]. union/non-union [LW98a]. Unit [KS93, Saï91, BM93a, Bie93, Bie97b, BFH97, BG97b, CP93, DNSW92, Dic93, Die93, FHP97a, FHP97b, FFM96, GP93, GLN94, GAR97, GS97b, Hal92, Hal94b, HT99a, HLvD95, HS96, Hyl95, KCW96, KPSS92, Lub95, Luc95, NT90, PH91, PS95c, SvD91, Seo99, ST98, TP93]. unit-root [GS97b, NT90]. Unit-roots [Saï91]. units [Mai90]. univariate [EMS99, Fra94, MM94, Sow92]. unknown [Abr99, BM99a, CD94, Cra92, GGH98, GH99, GJ92, HU94, IT98]. unknown-variance [HU94]. Unobserved [HRS92, GV97, JT94, MN95, MP99]. unrelated [KPS95, MMF93]. unspecified [Szp97a, Szp97b]. unstable [Bos94]. upwards [Eas91]. urban [TS97]. use [FOS97, FBA91, GKW95, Kem92, Kum96, MRB98]. Using [CFJL98, GHT90, MS97, BHL96, BKS96, BR97, CP98, DM92, DL95, Eas91, FD97, GS92, HS98a, Hal92, Hal94b, HKU97, HM98, Hus92, JM91, KP93a, KLOS97, Mag90, Mou95, NV90, Qui98a, Ruu91, SK96, Van92, Zab93, ZHkM91]. Utah [Ano90f]. utility [BS90b, KR94, SW94, vP91]. utilization [HY92, PN96, SS90].

validation [And91]. valuation [CR97, Sch96]. value [CM97, Koo91, Paa92, Sma94b]. values [AKS92, Hor94, KM97, ST98]. VAR [LR94, PR99]. VARFIMA [MW99]. variable [Abr99, AB95, BSH93, BM98, Che99, DL97b, GH99, Hal92, Hal94b, HASM98, Lai93, Lee94, Mro99, PH91, Pin93, SK96]. variables [Ame90, Att91, BYN98, BS94a, Dag94, DD97, DT94, Fri90, God98b, Hal94a, HNIP91, HNP95, Hon93, IASW99, Iwa92b, Iwa94, Kai98, KNI96, LS97a, Mag90, MN97, MB90, Met96, Now93, OG93, Per97, QS99, Run97a, Run97b, US94, VV99, Wan98]. variance [BCL92a, BM99a, BM93b, Can98, CN96, CD93, EMS99, Gil92, HU94, KT96, Nel92, NVvS91, PL91, Phi98, Run97a, Run97b, SC93, Smi94]. variances [BB92, Woo91a]. variates [DM92]. variation [CSS90, Sle90]. Varieties [GD96]. various [HS94b, Kiv95]. VARMA [LC97]. VARs [CD98, Phi98]. Vartia [RD99]. Varying [FLW98, CG95, DR97, HW96, Kum90, LW90, LH96, Osb91, Sue92, ZHkM91]. vector [AR94, BM93b, CP99, JS99b, LS97b, SL97, TP93, TY95, Tur98, YT98, Zha99]. vectors [BOYF94, BDR94, GJ92, Han92a, Qui98a, Sne99, Wic96, Yik98]. versus [GMP99, MS92, Sne99, Urb95]. vertical [WW96]. via [LH96, Phi94, SK98, Zhe96]. Vol [Ano92b, BCL93, Gil93, KO93, Hal94b]. volatility [AS98, AL97, AS97b, ACS99, Ano92s, BM96, BM99b, BCdL98, Dan94, DL92, DG96a, GHT90, GHT97, HT97, HS98b, MT90, PS90a, Rui94, Rui97, SK98, TZL96, WC95]. Volume [TZL96, Ano98b]. Volumes [AC97, AC94, Ano92t, MN95]. vs [BG97a].

Wage [CL96, Buc95b, DR93, LW98a, OR94]. wages [Kum96, TS97]. Wald [CA91, Kem91, LB97a, LB97b]. walk [HSPHS90]. walks [Ent97]. war [FH90]. water [fLRP97]. Watson [AS93, AKS92, Bar92, HM94, KW91, PL91]. way [BL92, BC94]. ways [Mir95]. Weak [PR99]. weakly [Mic99, SW94]. wealth [Cre90, Wol90]. web [Bar97]. weekend [Con91]. weekly [FKL98]. weighted [Ich93, PS96b, Szr94, ZHkM91]. weights [HM98, MRB98]. Welfare [Pol91, Sen91, Dag90, EGW90, HMS91, MWS91, Tin91, vP91]. WESML [Lan97]. where [HT99a]. which [LS97a, MS93]. white [BSTY91, CL96]. wide [Bar97]. will [BL91a]. winner [HS94c]. without [Ai94, CD98, KB95]. women [TS97]. work [Das97a, EGW90, Hor93, Slo90]. work-trip [Hor93]. work-welfare [EGW90]. world [Bar97]. wrong [Nel92, NF95].

years [HK91].

ZEF [KPS95]. Zin [Smi99]. zone [PS92].

References

Appelbaum:1991:DCR

[AB91] Elie Appelbaum and Joseph Berechman. Demand conditions, regulation, and the measurement of productivity. Journal of Econometrics, 47(2–3):379–400, February 3, 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

           URL http://www.sciencedirect.com/science/article/pii/
           030440769190108P.

Arellano:1995:ALI

[AB95] Manuel Arellano and Olympia Bover. Another look at the instrumental variable estimation of error-components models. Journal of Econometrics, 68(1):29–51, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401642D.

Asea:1998:LC

[AB98] Patrick K. Asea and Brock Blomberg. Lending cycles. Journal of Econometrics, 83(1–2):89–128, March/April 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000663.

Abeysinghe:1994:DSM

[Abe94] Tilak Abeysinghe. Deterministic seasonal models and spurious regressions. Journal of Econometrics, 61(2):259–272, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900868.

Abrevaya:1999:LEF

[Abr99] Jason Abrevaya. Leapfrog estimation of a fixed-effects model with unknown transformation of the dependent variable. Journal of Econometrics, 93(2):203–228, December 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000093.

Atkinson:1993:MTE

[AC93] Scott E. Atkinson and Christopher Cornwell. Measuring technical efficiency with panel data: A dual approach. Journal of Econometrics, 59(3):257–261, October 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390025Z.

Anderson-Courtney:1994:JES

[AC94] Linda Anderson-Courtney. Journal of Econometrics: Subject and author index: volumes 51–60, 1992–1994. Journal of Econometrics, 62(1):1–2, May 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900035.

Anderson-Courtney:1997:JES

[AC97] Linda Anderson-Courtney. Journal of Econometrics: Subject and author index: Volumes 61–75, 1994–1996. Journal of Econometrics, 78(1):1–130, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800036.

Andersen:1999:EMM

[ACS99] Torben G. Andersen, Hyung-Jin Chung, and Bent E. Sørensen. Efficient method of moments estimation of a stochastic volatility model: A Monte Carlo study. Journal of Econometrics, 91(1): 61–87, July 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000499.

Andrews:1992:EPD

[AF92] Donald W. K. Andrews and Ray C. Fair. Estimation of polynomial distributed lags and leads with end point constraints. Journal of Econometrics, 53(1–3):123–139, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900823.

Arguea:1993:EIE

[AH93] Nestor M. Arguea and Cheng Hsiao. Econometric issues of estimating hedonic price functions: With an application to the U.S. market for automobiles. Journal of Econometrics, 56(1–2):243– 267, March 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390108H.

Atkinson:1998:PTS

[AH98] Scott E. Atkinson and Robert Halvorsen. Parametric tests for static and dynamic equilibrium. Journal of Econometrics, 85(1): 33–50, July 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000924.

Ahn:1995:NTS

[Ahn95] Hyungtaik Ahn. Nonparametric two-stage estimation of conditional choice probabilities in a binary choice model under uncertainty. Journal of Econometrics, 67(2):337–378, June 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401636E.

Ai:1994:SEB

[Ai94] Chunrong Ai. A semiparametric efficiency bound of a disequilibrium model without observed regime. Journal of Econometrics, 62(2):143–163, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900205.

Aigner:1990:EI

[Aig90] Dennis J. Aigner. Editor’s introduction. Journal of Econometrics, 44(1–2):1–4, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407690900696.

Aigner:1991:E

[Aig91] Dennis J. Aigner. Editorial. Journal of Econometrics, 50(3): 229, December 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407691900178.

Aigner:1996:ES

[Aig96] Dennis J. Aigner. Editorial statement. Journal of Econometrics, 75(2):397–398, December 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407696900706.

Anderson:1992:TOP

[AK92] T. W. Anderson and Naoto Kunitomo. Tests of overidentification and predeterminedness in simultaneous equation models. Journal of Econometrics, 54(1–3):49–78, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290099D.

Anderson:1994:ART

[AK94] T. W. Anderson and Naoto Kunitomo. Asymptotic robustness of tests of overidentification and predeterminedness. Journal of Econometrics, 62(2):383–414, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900299.

Ansley:1992:CVG

[AKS92] Craig F. Ansley, Robert Kohn, and Thomas S. Shively. Computing p-values for the generalized Durbin–Watson and other invariant test statistics. Journal of Econometrics, 54(1–3):277–300, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692901095.

Atkinson:1997:DSO

[AKS97] A. C. Atkinson, S. J. Koopman, and N. Shephard. Detecting shocks: Outliers and breaks in time series. Journal of Econometrics, 80(2):387–422, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769700050X.

Ahn:1996:RHT

[AL96] Seung C. Ahn and Stuart Low. A reformulation of the Hausman test for regression models with pooled cross-section-time-series data. Journal of Econometrics, 71(1–2):309–319, March/ April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694017077.

Andersen:1997:ECT

[AL97] Torben G. Andersen and Jesper Lund. Estimating continuoustime stochastic volatility models of the short-term interest rate. Journal of Econometrics, 77(2):343–377, April 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696018192.

Andrews:1996:OCT

[ALP96] Donald W. K. Andrews, Inpyo Lee, and Werner Ploberger. Optimal changepoint tests for normal linear regression. Journal of Econometrics, 70(1):9–38, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016828.

Ahn:1993:DTA

[AM93] Hyungtaik Ahn and Charles F. Manski. Distribution theory for the analysis of binary choice under uncertainty with nonparametric estimation of expectations. Journal of Econometrics, 56(3): 291–321, April 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390123M.

Ai:1997:ESP

[AM97] Chunrong Ai and Daniel McFadden. Estimation of some partially specified nonlinear models. Journal of Econometrics, 76 (1–2):1–37, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769501778X.

Amemiya:1990:TSI

[Ame90] Yasuo Amemiya. Two-stage instrumental variables estimators for the nonlinear errors-in-variables model. Journal of Econometrics, 44(3):311–332, June 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090061W.

Asea:1998:HIA

[AN98] Patrick K. Asea and Mthuli Ncube. Heterogeneous information arrival and option pricing. Journal of Econometrics, 83 (1–2):291–323, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000730.

Ananda:1998:BNB

[Ana98] Malwane M. A. Ananda. Bayesian and non-Bayesian solutions to analysis of covariance models under heteroscedasticity. Journal of Econometrics, 86(1):177–192, September 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697001139.

Andrews:1991:AOG

[And91] Donald W. K. Andrews. Asymptotic optimality of generalized C_{L}, cross-validation, and generalized cross-validation in regression with heteroskedastic errors. Journal of Econometrics, 47 (2–3):359–377, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769190107O.

Anderson:1994:STD

[And94] Gordon Anderson. Simple tests of distributional form. Journal of Econometrics, 62(2):265–276, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900248.

Andrews:1998:HTR

[And98] Donald W. K. Andrews. Hypothesis testing with a restricted parameter space. Journal of Econometrics, 84(1):155–199, May 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000833.

Angrist:1991:GDE

[Ang91] Joshua D. Angrist. Grouped-data estimation and testing in simple labor-supply models. Journal of Econometrics, 47(2–3):243–266, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190101I.

Anonymous:1990:Ab

[Ano90a] Anonymous. Acknowledgement. Journal of Econometrics, 43(3): 397–398, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090130L.

Anonymous:1990:Ae

[Ano90b] Anonymous. Acknowledgement. Journal of Econometrics, 46 (1–2):1, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090043S.

Anonymous:1990:Aa

[Ano90c] Anonymous. Announcement. Journal of Econometrics, 43(3): 253, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090119E.

Anonymous:1990:Ac

[Ano90d] Anonymous. Announcement. Journal of Econometrics, 44(3): 239, June 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900582.

Anonymous:1990:Ad

[Ano90e] Anonymous. Announcement. Journal of Econometrics, 45(3): 399, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090008H.

Anonymous:1990:CSN

[Ano90f] Anonymous. The Classification Society of North America’s 1990 meeting: Logan, Utah, 21–23 June 1990: classification and clustering: Perspectives and prospects. Journal of Econometrics, 43 (3):396, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090129H.

Anonymous:1990:EBa

[Ano90g] Anonymous. Editorial board. Journal of Econometrics, 43 (1–2):ifc, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090102Y.

Anonymous:1990:EBb

[Ano90h] Anonymous. Editorial board. Journal of Econometrics, 44(1–2): ifc, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900685.

Anonymous:1990:EBc

[Ano90i] Anonymous. Editorial board. Journal of Econometrics, 45(1–2): ifc, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090090G.

Anonymous:1990:EBd

[Ano90j] Anonymous. Editorial board. Journal of Econometrics, 46 (1–2):ii, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090042R.

Anonymous:1990:Ea

[Ano90k] Anonymous. Erratum. Journal of Econometrics, 43(3):395, March 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090128G.

Anonymous:1990:Ec

[Ano90l] Anonymous. Erratum. Journal of Econometrics, 45(3):397, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090007G.

Anonymous:1990:Ia

[Ano90m] Anonymous. Index. Journal of Econometrics, 43(3):399–400, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090131C.

Anonymous:1990:Ib

[Ano90n] Anonymous. Index. Journal of Econometrics, 44(3):401, June 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690900674.

Anonymous:1990:Ic

[Ano90o] Anonymous. Index. Journal of Econometrics, 45(3):401, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090009I.

Anonymous:1990:Id

[Ano90p] Anonymous. Index. Journal of Econometrics, 46(3):407–408, December 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090018O.

Anonymous:1990:PAM

[Ano90q] Anonymous. Pages 1–238 (April–May 1990). Journal of Econometrics, 44(1–2):??, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1990:PON

[Ano90r] Anonymous. Pages 1–245 (October–November 1990). Journal of Econometrics, 46(1–2):??, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1990:PJF

[Ano90s] Anonymous. Pages 1–251 (January–February 1990). Journal of Econometrics, 43(1–2):??, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1990:PJA

[Ano90t] Anonymous. Pages 1–290 (July–August 1990). Journal of Econometrics, 45(1–2):??, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1990:PJ

[Ano90u] Anonymous. Pages 239–401 (June 1990). Journal of Econometrics, 44(3):??, June 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1990:PD

[Ano90v] Anonymous. Pages 247–408 (December 1990). Journal of Econometrics, 46(3):??, December 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1990:PM

[Ano90w] Anonymous. Pages 253–400 (March 1990). Journal of Econometrics, 43(3):??, March 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1990:P

[Ano90x] Anonymous. Pages 291–401 (1990). Journal of Econometrics, 45 (3):??, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:Ab

[Ano91a] Anonymous. Acknowledgement. Journal of Econometrics, 47 (2–3):401–402, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769190109Q.

Anonymous:1991:Aa

[Ano91b] Anonymous. Announcement. Journal of Econometrics, 47(2–3): 195, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190097W.

Anonymous:1991:Ac

[Ano91c] Anonymous. Announcement. Journal of Econometrics, 49(3):403, September 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190005X.

Anonymous:1991:Ad

[Ano91d] Anonymous. Announcement. Journal of Econometrics, 50(3): 233, December 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190019A.

Anonymous:1991:EBa

[Ano91e] Anonymous. Editorial board. Journal of Econometrics, 47(1): ifc, January 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190074N.

Anonymous:1991:EBb

[Ano91f] Anonymous. Editorial board. Journal of Econometrics, 48(1–2): ii, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190028C.

Anonymous:1991:EBc

[Ano91g] Anonymous. Editorial board. Journal of Econometrics, 49(1–2): ifc, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190007Z.

Anonymous:1991:EBd

[Ano91h] Anonymous. Editorial board. Journal of Econometrics, 50(1–2): ifc, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190084Q.

Anonymous:1991:E

[Ano91i] Anonymous. Erratum. Journal of Econometrics, 49(3):401, September 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190004W.

Anonymous:1991:Ia

[Ano91j] Anonymous. Index. Journal of Econometrics, 47(2–3):403, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190110Y.

Anonymous:1991:Ib

[Ano91k] Anonymous. Index. Journal of Econometrics, 48(3):411, June 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190073M.

Anonymous:1991:Ic

[Ano91l] Anonymous. Index. Journal of Econometrics, 49(3):405, September 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190006Y.

Anonymous:1991:Id

[Ano91m] Anonymous. Index. Journal of Econometrics, 50(3):399, December 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190027B.

Anonymous:1991:PJa

[Ano91n] Anonymous. Pages 1–194 (January 1991). Journal of Econometrics, 47(1):??, January 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:PO

[Ano91o] Anonymous. Pages 1–228 (11 October 1991). Journal of Econometrics, 50(1–2):??, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:PAM

[Ano91p] Anonymous. Pages 1–285 (April–May 1991). Journal of Econometrics, 48(1–2):??, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:PJAa

[Ano91q] Anonymous. Pages 1–304 (July–August 1991). Journal of Econometrics, 49(1–2):??, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:PF

[Ano91r] Anonymous. Pages 195–403 (3 February 1991). Journal of Econometrics, 47(2–3):??, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:PD

[Ano91s] Anonymous. Pages 229–399 (December 1991). Journal of Econometrics, 50(3):??, December 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:PJb

[Ano91t] Anonymous. Pages 287–411 (June 1991). Journal of Econometrics, 48(3):??, June 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1991:PS

[Ano91u] Anonymous. Pages 305–405 (September 1991). Journal of Econometrics, 49(3):??, September 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1992:A

[Ano92a] Anonymous. Acknowledgement. Journal of Econometrics, 53 (1–3):401–403, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407692900948.

Anonymous:1992:AEJ

[Ano92b] Anonymous. Acknowledgement/erratum: J. S. Cramer and G. Ridder, ‘Pooling states in the multinomial logit model’, Journal of Econometrics, Vol. 47, No. 2/3 (1991) pp. 267-272. Journal of Econometrics, 51(1–2):285–286, January/February 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290039T. See [CR91].

Anonymous:1992:CCE

[Ano92c] Anonymous. CERGE-the center for economic research and graduate education. Journal of Econometrics, 51(1–2):287, January/ February 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290041O.

Anonymous:1992:EBa

[Ano92d] Anonymous. Editorial board. Journal of Econometrics, 51 (1–2):ifc, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769290024L.

Anonymous:1992:EBb

[Ano92e] Anonymous. Editorial board. Journal of Econometrics, 52(1–2): ifc, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290062V.

Anonymous:1992:EBc

[Ano92f] Anonymous. Editorial board. Journal of Econometrics, 53 (1–3):ifc, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407692900753.

Anonymous:1992:EBd

[Ano92g] Anonymous. Editorial board. Journal of Econometrics, 54 (1–3):ii, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769290097B.

Anonymous:1992:FOS

[Ano92h] Anonymous. Fellow’s opinion section. Journal of Econometrics, 51(1–2):1, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769290025M.

Anonymous:1992:Ia

[Ano92i] Anonymous. Index. Journal of Econometrics, 51(1–2):289, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290042P.

Anonymous:1992:Ib

[Ano92j] Anonymous. Index. Journal of Econometrics, 52(3):423–424, June 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290023K.

Anonymous:1992:Ic

[Ano92k] Anonymous. Index. Journal of Econometrics, 53(1–3):405, July/ September 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692900959.

Anonymous:1992:Id

[Ano92l] Anonymous. Index. Journal of Econometrics, 54(1–3):401, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692901158.

Anonymous:1992:PJF

[Ano92m] Anonymous. Pages 1–289 (January–February 1992). Journal of Econometrics, 51(1–2):??, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1992:PAM

[Ano92n] Anonymous. Pages 1–311 (April–May 1992). Journal of Econometrics, 52(1–2):??, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1992:POD

[Ano92o] Anonymous. Pages 1–401 (October–December 1992). Journal of Econometrics, 54(1–3):??, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1992:PJS

[Ano92p] Anonymous. Pages 1–405 (July–September 1992). Journal of Econometrics, 53(1–3):??, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1992:PM

[Ano92q] Anonymous. Pages 291–382 (March 1992). Journal of Econometrics, 51(3):??, March 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1992:PJ

[Ano92r] Anonymous. Pages 313–424 (June 1992). Journal of Econometrics, 52(3):??, June 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1992:SMF

[Ano92s] Anonymous. Statistical models for financial volatility. Journal of Econometrics, 52(1–2):1–4, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290063W.

Anonymous:1992:SAI

[Ano92t] Anonymous. Subject and author index: volumes 41–50, 1989– 1991. Journal of Econometrics, 51(3):291–382, March 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290096A.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial board. Journal of Econometrics, 56(1– 2):ii, March 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390097O.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial board. Journal of Econometrics, 57(1–3): ifc, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390055A.

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial board. Journal of Econometrics, 58(1– 2):ii, July 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390109I.

Anonymous:1993:EBd

[Ano93d] Anonymous. Editorial board. Journal of Econometrics, 59(1–2): iv, September 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900354.

Anonymous:1993:Ia

[Ano93e] Anonymous. Index. Journal of Econometrics, 55(1–2):357, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390022W.

Anonymous:1993:Ib

[Ano93f] Anonymous. Index. Journal of Econometrics, 56(3):447, April 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390130W.

Anonymous:1993:Ic

[Ano93g] Anonymous. Index. Journal of Econometrics, 57(1–3):407, May/ June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390074F.

Anonymous:1993:Id

[Ano93h] Anonymous. Index. Journal of Econometrics, 58(3):407, August 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693900549.

Anonymous:1993:Ie

[Ano93i] Anonymous. Index. Journal of Econometrics, 59(3):407–408, October 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693900343.

Anonymous:1993:PS

[Ano93j] Anonymous. Pages 1–211 (September 1993). Journal of Econometrics, 59(1–2):??, September 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1993:PM

[Ano93k] Anonymous. Pages 1–267 (March 1993). Journal of Econometrics, 56(1–2):??, March 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1993:PJ

[Ano93l] Anonymous. Pages 1–274 (July 1993). Journal of Econometrics, 58(1–2):??, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1993:PJF

[Ano93m] Anonymous. Pages 1–357 (January–February 1993). Journal of Econometrics, 55(1–2):??, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1993:PMJ

[Ano93n] Anonymous. Pages 1–407 (May–June 1993). Journal of Econometrics, 57(1–3):??, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1993:PO

[Ano93o] Anonymous. Pages 213–408 (October 1993). Journal of Econometrics, 59(3):??, October 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1993:PAa

[Ano93p] Anonymous. Pages 269–447 (April 1993). Journal of Econometrics, 56(3):??, April 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1993:PAb

[Ano93q] Anonymous. Pages 275–407 (August 1993). Journal of Econometrics, 58(3):??, August 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:Aa

[Ano94a] Anonymous. Acknowledgement. Journal of Econometrics, 60(1– 2):323–324, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900515.

Anonymous:1994:Ab

[Ano94b] Anonymous. Announcement. Journal of Econometrics, 64 (1–2):1, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769490054X.

Anonymous:1994:AI

[Ano94c] Anonymous. Author index. Journal of Econometrics, 62(1): 51–66, May 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900051.

Anonymous:1994:EBa

[Ano94d] Anonymous. Editorial board. Journal of Econometrics, 60 (1–2):ifc, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900353.

Anonymous:1994:EBb

[Ano94e] Anonymous. Editorial board. Journal of Econometrics, 61(1): iv, March 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900728.

Anonymous:1994:EBc

[Ano94f] Anonymous. Editorial board. Journal of Econometrics, 62(1): ifc, May 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900027.

Anonymous:1994:EBd

[Ano94g] Anonymous. Editorial board. Journal of Econometrics, 63(1):ii, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900345.

Anonymous:1994:EBe

[Ano94h] Anonymous. Editorial board. Journal of Econometrics, 64 (1–2):ifc, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900531.

Anonymous:1994:Ia

[Ano94i] Anonymous. Index. Journal of Econometrics, 60(1–2):325, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900523.

Anonymous:1994:Ib

[Ano94j] Anonymous. Index. Journal of Econometrics, 61(2):429–430, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900949.

Anonymous:1994:Ic

[Ano94k] Anonymous. Index. Journal of Econometrics, 62(2):443, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900310.

Anonymous:1994:Id

[Ano94l] Anonymous. Index. Journal of Econometrics, 63(2):407, August 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900337.

Anonymous:1994:Ie

[Ano94m] Anonymous. Index. Journal of Econometrics, 64(1–2): 401, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769490071X.

Anonymous:1994:PMa

[Ano94n] Anonymous. Pages 1–196 (March 1994). Journal of Econometrics, 61(1):??, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:PJF

[Ano94o] Anonymous. Pages 1–325 (January–February 1994). Journal of Econometrics, 60(1–2):??, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:PJb

[Ano94p] Anonymous. Pages 1–325 (July 1994). Journal of Econometrics, 63(1):??, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:PSO

[Ano94q] Anonymous. Pages 1–401 (September–October 1994). Journal of Econometrics, 64(1–2):??, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:PMb

[Ano94r] Anonymous. Pages 1–66 (May 1994). Journal of Econometrics, 62(1):??, May 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:PAa

[Ano94s] Anonymous. Pages 197–430 (April 1994). Journal of Econometrics, 61(2):??, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:PAb

[Ano94t] Anonymous. Pages 327–408 (August 1994). Journal of Econometrics, 63(2):??, August 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:PJa

[Ano94u] Anonymous. Pages 67–444 (June 1994). Journal of Econometrics, 62(2):??, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1994:SI

[Ano94v] Anonymous. Subject index. Journal of Econometrics, 62(1): 3–50, May 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900043.

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial board. Journal of Econometrics, 65(1): ii, January 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769590011X.

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial board. Journal of Econometrics, 66 (1–2):iv–v, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695900039.

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial board. Journal of Econometrics, 67(1):ii, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695900330.

Anonymous:1995:EBd

[Ano95d] Anonymous. Editorial board. Journal of Econometrics, 68(1): ii–iii, July 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695900071.

Anonymous:1995:EBe

[Ano95e] Anonymous. Editorial board. Journal of Econometrics, 69(1): ii, September 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695900012.

Anonymous:1995:Ia

[Ano95f] Anonymous. Index. Journal of Econometrics, 65(2):429, February 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695900195.

Anonymous:1995:Ib

[Ano95g] Anonymous. Index. Journal of Econometrics, 66(1–2):371, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695900055.

Anonymous:1995:Ic

[Ano95h] Anonymous. Index. Journal of Econometrics, 67(2):403, June 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695900179.

Anonymous:1995:Id

[Ano95i] Anonymous. Index. Journal of Econometrics, 68(2):399, August 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695900314.

Anonymous:1995:Ie

[Ano95j] Anonymous. Index. Journal of Econometrics, 69(2):429, October 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695900004.

Anonymous:1995:PM

[Ano95k] Anonymous. Pages 1–257 (May 1995). Journal of Econometrics, 67(1):??, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1995:PJc

[Ano95l] Anonymous. Pages 1–268 (July 1995). Journal of Econometrics, 68(1):??, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1995:PJa

[Ano95m] Anonymous. Pages 1–332 (January 1995). Journal of Econometrics, 65(1):??, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1995:PMA

[Ano95n] Anonymous. Pages 1–371 (March–April 1995). Journal of Econometrics, 66(1–2):??, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1995:PJb

[Ano95o] Anonymous. Pages 259–403 (June 1995). Journal of Econometrics, 67(2):??, June 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1995:PA

[Ano95p] Anonymous. Pages 269–399 (August 1995). Journal of Econometrics, 68(2):??, August 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1995:PF

[Ano95q] Anonymous. Pages 333–429 (February 1995). Journal of Econometrics, 65(2):??, February 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1995:PO

[Ano95r] Anonymous. Pages 367–429 (October 1995). Journal of Econometrics, 69(2):??, October 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:Aa

[Ano96a] Anonymous. Acknowledgement. Journal of Econometrics, 71 (1–2):399–402, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407632900013.

Anonymous:1996:Ab

[Ano96b] Anonymous. Announcement. Journal of Econometrics, 72(1–2): 397, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696900056.

Anonymous:1996:JEF

[Ano96c] Anonymous. Journal of Econometrics Fellows —- 1996. Journal of Econometrics, 75(2):399–400, December 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696900718.

Anonymous:1996:EBa

[Ano96d] Anonymous. Editorial board. Journal of Econometrics, 70(1): ii, January 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696900019.

Anonymous:1996:EBb

[Ano96e] Anonymous. Editorial board. Journal of Econometrics, 71(1–2):ii– iii, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619900017.

Anonymous:1996:EBc

[Ano96f] Anonymous. Editorial board. Journal of Econometrics, 72(1–2):ii– iii, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696900032.

Anonymous:1996:EBd

[Ano96g] Anonymous. Editorial board. Journal of Econometrics, 73(1): iv–v, July 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696900676.

Anonymous:1996:EBe

[Ano96h] Anonymous. Editorial board. Journal of Econometrics, 74(1): ii, September 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696900639.

Anonymous:1996:EBf

[Ano96i] Anonymous. Editorial board. Journal of Econometrics, 75(1): ii, November 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769690069X.

Anonymous:1996:Ia

[Ano96j] Anonymous. Index. Journal of Econometrics, 70(2):407, February 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696900731.

Anonymous:1996:Ib

[Ano96k] Anonymous. Index. Journal of Econometrics, 71(1–2):403, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440765690001X.

Anonymous:1996:Ic

[Ano96l] Anonymous. Index. Journal of Econometrics, 72(1–2):399, May/ June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440769690007X.

Anonymous:1996:Id

[Ano96m] Anonymous. Index. Journal of Econometrics, 73(2):411, August 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698800031.

Anonymous:1996:Ie

[Ano96n] Anonymous. Index. Journal of Econometrics, 74(2):407, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696900652.

Anonymous:1996:If

[Ano96o] Anonymous. Index. Journal of Econometrics, 75(2):401, December 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440769690072X.

Anonymous:1996:PS

[Ano96p] Anonymous. Pages 1–208 (September 1996). Journal of Econometrics, 74(1):??, September 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PN

[Ano96q] Anonymous. Pages 1–238 (November 1996). Journal of Econometrics, 75(1):??, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PJa

[Ano96r] Anonymous. Pages 1–316 (January 1996). Journal of Econometrics, 70(1):??, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PJb

[Ano96s] Anonymous. Pages 1–324 (July 1996). Journal of Econometrics, 73(1):??, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PMJ

[Ano96t] Anonymous. Pages 1–399 (May–June 1996). Journal of Econometrics, 72(1–2):??, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PMA

[Ano96u] Anonymous. Pages 1–403 (March–April 1996). Journal of Econometrics, 71(1–2):??, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PO

[Ano96v] Anonymous. Pages 209–407 (October 1996). Journal of Econometrics, 74(2):??, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PD

[Ano96w] Anonymous. Pages 239–401 (December 1996). Journal of Econometrics, 75(2):??, December 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PF

[Ano96x] Anonymous. Pages 317–407 (February 1996). Journal of Econometrics, 70(2):??, February 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1996:PA

[Ano96y] Anonymous. Pages 325–411 (August 1996). Journal of Econometrics, 73(2):??, August 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:Aa

[Ano97a] Anonymous. Acknowledgement. Journal of Econometrics, 76(1– 2):407–408, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697900178.

Anonymous:1997:AFJ

[Ano97b] Anonymous. Announcement: Fellows of the journal of econometrics. Journal of Econometrics, 78(1):131–133, ???? 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800048.

Anonymous:1997:Ab

[Ano97c] Anonymous. Announcements. Journal of Econometrics, 82(1): 193–195, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698800018.

Anonymous:1997:JEA

[Ano97d] Anonymous. Journal of Econometrics Annals 1979–1996. Journal of Econometrics, 78(1):135–138, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769780005X.

Anonymous:1997:EBa

[Ano97e] Anonymous. Editorial board. Journal of Econometrics, 76 (1–2):ii, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697900154.

Anonymous:1997:EBb

[Ano97f] Anonymous. Editorial board. Journal of Econometrics, 77(1): ii, March 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697900099.

Anonymous:1997:EBc

[Ano97g] Anonymous. Editorial board. Journal of Econometrics, 78(1): i, iv, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697800024.

Anonymous:1997:EBd

[Ano97h] Anonymous. Editorial board. Journal of Econometrics, 79(1): ii–iii, July 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697900191.

Anonymous:1997:EBe

[Ano97i] Anonymous. Editorial board. Journal of Econometrics, 80(1):ii– iii, September 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697900117.

Anonymous:1997:EBf

[Ano97j] Anonymous. Editorial board. Journal of Econometrics, 81(1): ii, November 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697900130.

Anonymous:1997:EBg

[Ano97k] Anonymous. Editorial board. Journal of Econometrics, 82(1): iii, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698900008.

Anonymous:1997:EI

[Ano97l] Anonymous. Editors introduction. Journal of Econometrics, 80(2):195–197, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000304.

Anonymous:1997:Ia

[Ano97m] Anonymous. Index. Journal of Econometrics, 77(2):405–406, April 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697873782.

Anonymous:1997:Ib

[Ano97n] Anonymous. Index. Journal of Econometrics, 78(1):381, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697800188.

Anonymous:1997:Ic

[Ano97o] Anonymous. Index. Journal of Econometrics, 78(2):381, June 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697880499.

Anonymous:1997:Id

[Ano97p] Anonymous. Index. Journal of Econometrics, 80(2):423, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697825882.

Anonymous:1997:Ie

[Ano97q] Anonymous. Index. Journal of Econometrics, 81(2):395, December 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697865724.

Anonymous:1997:PJb

[Ano97r] Anonymous. Pages 1–193 (July 1997). Journal of Econometrics, 79(1):??, July 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PS

[Ano97s] Anonymous. Pages 1–193 (September 1997). Journal of Econometrics, 80(1):??, September 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:Pc

[Ano97t] Anonymous. Pages 1–195 (1997). Journal of Econometrics, 82 (1):??, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PN

[Ano97u] Anonymous. Pages 1–280 (November 1997). Journal of Econometrics, 81(1):??, November 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PM

[Ano97v] Anonymous. Pages 1–295 (March 1997). Journal of Econometrics, 77(1):??, March 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:Pb

[Ano97w] Anonymous. Pages 1–381 (1997). Journal of Econometrics, 78 (1):??, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PJF

[Ano97x] Anonymous. Pages 1–408 (January–February 1997). Journal of Econometrics, 76(1–2):??, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PJa

[Ano97y] Anonymous. Pages 139–381 (June 1997). Journal of Econometrics, 78(2):??, June 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PAb

[Ano97z] Anonymous. Pages 195–402 (August 1997). Journal of Econometrics, 79(2):??, August 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PO

[Ano97-27] Anonymous. Pages 195–423 (October 1997). Journal of Econometrics, 80(2):??, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PD

[Ano97-28] Anonymous. Pages 281–395 (December 1997). Journal of Econometrics, 81(2):??, December 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1997:PAa

[Ano97-29] Anonymous. Pages 297–406 (April 1997). Journal of Econometrics, 77(2):??, April 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:AFJ

[Ano98a] Anonymous. Announcement: Fellows of the journal of econometrics. Journal of Econometrics, 84(1):201–203, May 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698900112.

Anonymous:1998:AIV

[Ano98b] Anonymous. Author index to volume 89. Journal of Econometrics, 89(1–2):457–458, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000839.

Anonymous:1998:EBa

[Ano98c] Anonymous. Editorial board. Journal of Econometrics, 83(1–2): ii, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697802278.

Anonymous:1998:EBb

[Ano98d] Anonymous. Editorial board. Journal of Econometrics, 84(1):ii, May 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698900094.

Anonymous:1998:EBc

[Ano98e] Anonymous. Editorial board. Journal of Econometrics, 85(1):ii, July 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440769890001X.

Anonymous:1998:Ia

[Ano98f] Anonymous. Index. Journal of Econometrics, 82(2):393, February 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697815758.

Anonymous:1998:Ib

[Ano98g] Anonymous. Index. Journal of Econometrics, 84(2):401, June 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698800055.

Anonymous:1998:Ic

[Ano98h] Anonymous. Index. Journal of Econometrics, 85(2):401, August 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000189.

Anonymous:1998:Id

[Ano98i] Anonymous. Index. Journal of Econometrics, 86(2):397, October 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000190.

Anonymous:1998:Ie

[Ano98j] Anonymous. Index. Journal of Econometrics, 87(2):373, December 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000724.

Anonymous:1998:PS

[Ano98k] Anonymous. Pages 1–192 (September 1998). Journal of Econometrics, 86(1):??, September 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PJb

[Ano98l] Anonymous. Pages 1–203 (July 1998). Journal of Econometrics, 85(1):??, July 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PM

[Ano98m] Anonymous. Pages 1–203 (May 1998). Journal of Econometrics, 84(1):??, May 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PNa

[Ano98n] Anonymous. Pages 1–206 (November 1998). Journal of Econometrics, 87(1):??, November 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PMA

[Ano98o] Anonymous. Pages 1–363 (March–April 1998). Journal of Econometrics, 83(1–2):??, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PNb

[Ano98p] Anonymous. Pages 1–458 (26 November 1998). Journal of Econometrics, 89(1–2):??, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PO

[Ano98q] Anonymous. Pages 193–398 (October 1998). Journal of Econometrics, 86(2):??, October 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PF

[Ano98r] Anonymous. Pages 197–393 (February 1998). Journal of Econometrics, 82(2):??, February 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PA

[Ano98s] Anonymous. Pages 205–401 (August 1998). Journal of Econometrics, 85(2):??, August 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PJa

[Ano98t] Anonymous. Pages 205–401 (June 1998). Journal of Econometrics, 84(2):??, June 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1998:PD

[Ano98u] Anonymous. Pages 207–374 (December 1998). Journal of Econometrics, 87(2):??, December 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:E

[Ano99a] Anonymous. Erratum. Journal of Econometrics, 90(2):337–343, June 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000736.

Anonymous:1999:Ia

[Ano99b] Anonymous. Index. Journal of Econometrics, 88(2):403–404, February 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000827.

Anonymous:1999:Ib

[Ano99c] Anonymous. Index. Journal of Econometrics, 90(2):345, June 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000020.

Anonymous:1999:Ic

[Ano99d] Anonymous. Index. Journal of Econometrics, 91(2):403, August 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440769900024X.

Anonymous:1999:Id

[Ano99e] Anonymous. Index. Journal of Econometrics, 92(2):391, October 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000548.

Anonymous:1999:Ie

[Ano99f] Anonymous. Index. Journal of Econometrics, 93(2):403, December 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440769900055X.

Anonymous:1999:PM

[Ano99g] Anonymous. Pages 1–154 (May 1999). Journal of Econometrics, 90(1):??, May 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PS

[Ano99h] Anonymous. Pages 1–192 (September 1999). Journal of Econometrics, 92(1):??, September 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PJc

[Ano99i] Anonymous. Pages 1–200 (July 1999). Journal of Econometrics, 91(1):??, July 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PJa

[Ano99j] Anonymous. Pages 1–202 (January 1999). Journal of Econometrics, 88(1):??, January 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PN

[Ano99k] Anonymous. Pages 1–202 (November 1999). Journal of Econometrics, 93(1):??, November 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PJb

[Ano99l] Anonymous. Pages 155–346 (June 1999). Journal of Econometrics, 90(2):??, June 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PO

[Ano99m] Anonymous. Pages 193–392 (October 1999). Journal of Econometrics, 92(2):??, October 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PA

[Ano99n] Anonymous. Pages 201–404 (August 1999). Journal of Econometrics, 91(2):??, August 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PD

[Ano99o] Anonymous. Pages 203–404 (December 1999). Journal of Econometrics, 93(2):??, December 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1999:PF

[Ano99p] Anonymous. Pages 203–404 (February 1999). Journal of Econometrics, 88(2):??, February 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Ahn:1993:SEC

[AP93] Hyungtaik Ahn and James L. Powell. Semiparametric estimation of censored selection models with a nonparametric selection mechanism. Journal of Econometrics, 58(1–2):3–29, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390111H.

Ahn:1994:EPN

[AR94] Sung K. Ahn and Gregory C. Reinsel. Estimation of partially nonstationary vector autoregressive models with seasonal behavior. Journal of Econometrics, 62(2):317–350, June 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900272.

Allenby:1998:MMC

[AR98] Greg M. Allenby and Peter E. Rossi. Marketing models of consumer heterogeneity. Journal of Econometrics, 89(1–2):57–78, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000554.

Arellano:1993:TCE

[Are93] Manuel Arellano. On the testing of correlated effects with panel data. Journal of Econometrics, 59(1–2):87–97, September 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390040C.

Ali:1993:RND

[AS93] Mukhtar M. Ali and Subhash C. Sharma. Robustness to nonnormality of the Durbin–Watson test for autocorrelation. Journal of Econometrics, 57(1–3):117–136, May/June 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900619.

Ahn:1995:EEM

[AS95] Seung C. Ahn and Peter Schmidt. Efficient estimation of models for dynamic panel data. Journal of Econometrics, 68(1): 5–27, July 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401641C.

Ali:1996:RNR

[AS96] Mukhtar M. Ali and Subhash C. Sharma. Robustness to nonnormality of regression F -tests. Journal of Econometrics, 71 (1–2):175–205, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401700X.

Ahn:1997:EED

[AS97a] Seung C. Ahn and Peter Schmidt. Efficient estimation of dynamic panel data models: Alternative assumptions and simplified estimation. Journal of Econometrics, 76(1–2):309–321, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017933.

Andersen:1997:GQA

[AS97b] Torben G. Andersen and Bent E. Sørensen. GMM and QML asymptotic standard deviations in stochastic volatility models: Comments on Ruiz (1994). Journal of Econometrics, 76(1–2):397–403, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407695017992. See [Rui94] and reply [Rui97].

Ait-Sahalia:1998:DEV

[AS98] Yacine Aït-Sahalia. Dynamic equilibrium and volatility in financial asset markets. Journal of Econometrics, 84(1):93–127, May 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697800012.

Attfield:1991:ETW

[Att91] C. L. F. Attfield. Estimation and testing when explanatory variables are endogenous: An application to a demand system. Journal of Econometrics, 48(3):395–408, June 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190071K.

Attfield:1995:BAL

[Att95] C. L. F. Attfield. A Bartlett adjustment to the likelihood ratio test for a system of equations. Journal of Econometrics, 66(1–2):207–223, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016157.

Anderson:1998:TME

[AV98] Heather M. Anderson and Farshid Vahid. Testing multiple equation systems for common nonlinear components. Journal of Econometrics, 84(1):1–36, May 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000766.

Baillie:1996:LMP

[Bai96] Richard T. Baillie. Long memory processes and fractional integration in econometrics. Journal of Econometrics, 73(1):5–59, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017321.

Bai:1999:LRT

[Bai99] Jushan Bai. Likelihood ratio tests for multiple structural changes. Journal of Econometrics, 91(2):299–323, August 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000797.

Baltagi:1995:EIP

[Bal95] Badi H. Baltagi. Editor’s introduction panel data. Journal of Econometrics, 68(1):1–4, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695900098.

Bartels:1992:PFD

[Bar92] Robert Bartels. On the power function of the Durbin–Watson test. Journal of Econometrics, 51(1–2):101–112, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290031L.

Barnett:1997:FOE

[Bar97] William A. Barnett. Fellow’s opinion: Econometrics, data, and the world wide web. Journal of Econometrics, 77(2):297–302, April 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696018179.

Bauer:1990:RDE

[Bau90] Paul W. Bauer. Recent developments in the econometric estimation of frontiers. Journal of Econometrics, 46(1–2):39–56, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090046V.

Baillie:1992:PDM

[BB92] Richard T. Baillie and Tim Bollerslev. Prediction in dynamic models with time-dependent conditional variances. Journal of Econometrics, 52(1–2):91–113, April/May 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290066Z.

Blundell:1998:ICM

[BB98] Richard Blundell and Stephen Bond. Initial conditions and moment restrictions in dynamic panel data models. Journal of Econometrics, 87(1):115–143, November 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000098.

Blundell:1992:ITE

[BBDS92] Richard Blundell, Stephen Bond, Michael Devereux, and Fabio Schiantarelli. Investment and Tobin’s Q: Evidence from company panel data. Journal of Econometrics, 51(1–2):233–257, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290037R.

Baillie:1996:FIG

[BBM96] Richard T. Baillie, Tim Bollerslev, and Hans Ole Mikkelsen. Fractionally integrated generalized autoregressive conditional heteroskedasticity. Journal of Econometrics, 74(1):3–30, September 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407695017496.

Baltagi:1994:IPC

[BC94] Badi H. Baltagi and Young-Jae Chang. Incomplete panels: A comparative study of alternative estimators for the unbalanced one-way error component regression model. Journal of Econometrics, 62(2):67–89, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900175.

Breidt:1998:DEL

[BCdL98] F. Jay Breidt, Nuno Crato, and Pedro de Lima. The detection and estimation of long memory in stochastic volatility. Journal of Econometrics, 83(1–2):325–348, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000729.

Bollerslev:1992:AMF

[BCK92] Tim Bollerslev, Ray Y. Chou, and Kenneth F. Kroner. ARCH modeling in finance: A review of the theory and empirical evidence. Journal of Econometrics, 52(1–2):5–59, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290064X.

Baltagi:1992:MCE

[BCL92a] Badi H. Baltagi, Young-Jae Chang, and Qi Li. Monte Carlo evidence on panel data regressions with AR(1) disturbances and an arbitrary variance on the initial observations. Journal of Econometrics, 52(3):371–380, June 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290017L.

Baltagi:1992:MCR

[BCL92b] Badi H. Baltagi, Young-Jae Chang, and Qi Li. Monte Carlo results on several new and existing tests for the error component model. Journal of Econometrics, 54(1–3):95–120, October/ December 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290101V.

Baltagi:1993:MCR

[BCL93] Badi H. Baltagi, Young-Jae Chang, and Qi Li. Monte Carlo results on several new and existing tests for the error component model (vol. 54, no. 1-3 (1992) pp. 95-120). Journal of Econometrics, 59 (3):405, October 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900332.

Bernard:1996:ITC

[BD96] Andrew B. Bernard and Steven N. Durlauf. Interpreting tests of the convergence hypothesis. Journal of Econometrics, 71 (1–2):161–173, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016992.

Boudjellaba:1994:SCN

[BDR94] Hafida Boudjellaba, Jean-Marie Dufour, and Roch Roy. Simplified conditions for noncausality between vectors in multivariate ARMA models. Journal of Econometrics, 63(1):271–287, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693015687.

Burgess:1996:EIA

[BEP96] Simon Burgess, Alvaro Escribano, and Gerard Pfann. Editor’s introduction: Asymmetries and nonlinearities in dynamic economic models. Journal of Econometrics, 74(1):1–2, September 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017488.

Boswijk:1997:MUR

[BFH97] H. Peter Boswijk, Philip Hans Franses, and Niels Haldrup. Multiple unit roots in periodic autoregression. Journal of Econometrics, 80(1):167–193, September 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440769781127X.

Bregman:1995:PCS

[BFR95] Arie Bregman, Melvyn Fuss, and Haim Regev. The production and cost structure of Israeli industry evidence from individual firm data. Journal of Econometrics, 65(1):45–81, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401597S.

Behrman:1997:DAP

[BFR97] Jere R. Behrman, Andrew D. Foster, and Mark R. Rosenzweig. The dynamics of agricultural production and the calorie-income relationship: Evidence from Pakistan. Journal of Econometrics, 77(1):187–207, March 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769601812X.

Baltagi:1997:PEV

[BG97a] Badi H. Baltagi and James M. Griffin. Pooled estimators vs. their heterogeneous counterparts in the context of dynamic demand for gasoline. Journal of Econometrics, 77(2):303–327, April 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696018027.

Breitung:1997:RTU

[BG97b] Jörg Breitung and Christian Gouriéroux. Rank tests for unit roots. Journal of Econometrics, 81(1):7–27, November 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000316.

Baier:1998:OPP

[BG98a] Daniel Baier and Wolfgang Gaul. Optimal product positioning based on paired comparison data. Journal of Econometrics, 89(1–2):365–392, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000670.

Broze:1998:PML

[BG98b] Laurence Broze and Christian Gouriéroux. Pseudo-maximum likelihood method, adjusted pseudo-maximum likelihood method and covariance estimators. Journal of Econometrics, 85(1):75–98, July 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440769700095X.

Barnett:1997:SBC

[BGH^{+}97] William A. Barnett, A. Ronald Gallant, Melvin J. Hinich, Jochen A. Jungeilges, Daniel T. Kaplan, and Mark J. Jensen. A single-blind controlled competition among tests for nonlinearity and chaos. Journal of Econometrics, 82(1):157–192, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440769700081X.

Bansal:1995:NES

[BGHT95] Ravi Bansal, A. Ronald Gallant, Robert Hussey, and George Tauchen. Nonparametric estimation of structural models for high-frequency currency market data. Journal of Econometrics, 66(1–2):251–287, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401618A.

Balestra:1992:A

[BGK92] Pietro Balestra, Stephen Goldfeld, and Teun Kloek. Announcement. Journal of Econometrics, 52(3):313, June 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290014I.

Berndt:1995:EEP

[BGR95] Ernst R. Berndt, Zvi Griliches, and Neal J. Rappaport. Econometric estimates of price indexes for personal computers in the 1990’s. Journal of Econometrics, 68(1):243–268, July 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401651F.

Barnett:1991:SBE

[BGW91] William A. Barnett, John Geweke, and Michael Wolfe. Seminonparametric Bayesian estimation of the asymptotically ideal production model. Journal of Econometrics, 49(1–2):5–50, July/ August 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407691900093.

Bossaerts:1997:LPA

[BH97] Peter Bossaerts and Pierre Hillion. Local parametric analysis of hedging in discrete time. Journal of Econometrics, 81(1):243– 272, November 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000468.

Bhargava:1997:EIA

[Bha97a] Alok Bhargava. Editor’s introduction: Analysis of data on health. Journal of Econometrics, 77(1):1–4, March 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696018039.

Bhargava:1997:NSA

[Bha97b] Alok Bhargava. Nutritional status and the allocation of time in rwandese households. Journal of Econometrics, 77(1):277– 295, March 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696018167.

Bjurek:1990:DPN

[BHF90] Hans Bjurek, Lennart Hjalmarsson, and Finn R. Forsund. Deterministic parametric and nonparametric estimation of efficiency in service production: A comparison. Journal of Econometrics, 46(1–2):213–227, October/November 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090056Y.

Boisso:1994:OSM

[BHHS94] Dale Boisso, Kathy Hayes, Joseph Hirschberg, and Jacques Silber. Occupational segregation in the multidimensional case: Decomposition and tests of significance. Journal of Econometrics, 61(1): 161–171, March 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900825.

Baltagi:1996:NTP

[BHL96] Badi H. Baltagi, Javier Hidalgo, and Qi Li. A nonparametric test for poolability using panel data. Journal of Econometrics, 75(2):345–367, December 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017798.

Basmann:1990:GFF

[BHSJ90] R. L. Basmann, K. J. Hayes, D. J. Slottje, and J. D. Johnson. A general functional form for approximating the Lorenz curve. Journal of Econometrics, 43(1–2):77–90, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690901086.

Bierens:1993:HOS

[Bie93] Herman J. Bierens. Higher-order sample autocorrelations and the unit root hypothesis. Journal of Econometrics, 57(1–3):137–160, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390062A.

Bierens:1997:NCA

[Bie97a] Herman J. Bierens. Nonparametric cointegration analysis. Journal of Econometrics, 77(2):379–404, April 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696018209.

Bierens:1997:TUR

[Bie97b] Herman J. Bierens. Testing the unit root with drift hypothesis against nonlinear trend stationarity, with an application to the US price level and interest rate. Journal of Econometrics, 81(1):29– 64, November 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769700033X.

Baillie:1996:EIF

[BK96] Richard T. Baillie and Maxwell L. King. Editors’ introduction: Fractional differencing and long memory processes. Journal of Econometrics, 73(1):1–3, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017313.

Blommestein:1998:ISS

[BK98] Hans J. Blommestein and Nick A. M. Koper. The influence of sample size on the degree of redundancy in spatial lag operators. Journal of Econometrics, 82(2):317–333, February 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000602.

Barnett:1996:BEA

[BKS96] Glen Barnett, Robert Kohn, and Simon Sheather. Bayesian estimation of an autoregressive model using Markov chain Monte Carlo. Journal of Econometrics, 74(2):237–254, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017445.

Baltagi:1991:TWC

[BL91a] Badi H. Baltagi and Qi Li. A transformation that will circumvent the problem of autocorrelation in an error-component model. Journal of Econometrics, 48(3):385–393, June 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190070T.

Blundell:1991:ICE

[BL91b] Richard Blundell and Arthur Lewbel. The information content of equivalence scales. Journal of Econometrics, 50(1–2):49–68, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190089V.

Baltagi:1992:MPI

[BL92] Badi H. Baltagi and Qi Li. A monotonic property for iterative GLS in the two-way random effects model. Journal of Econometrics, 53(1–3):45–51, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900797.

Baltagi:1995:TAA

[BL95a] Badi H. Baltagi and Qi Li. Testing AR(1) against MA(1) disturbances in an error component model. Journal of Econometrics, 68(1):133–151, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401646H.

Bauwens:1995:EIB

[BL95b] Luc Bauwens and Michel Lubrano. Editors’ introduction Bayesian and classical econometric modeling of time series. Journal of Econometrics, 69(1):1–4, September 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401659N.

Bertschek:1998:CEP

[BL98] Irene Bertschek and Michael Lechner. Convenient estimators for the panel probit model. Journal of Econometrics, 87(2):329–371, December 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000086.

Blomquist:1996:EMM

[Blo96] Sören Blomquist. Estimation methods for male labor supply functions how to take account of nonlinear taxes. Journal of Econometrics, 70(2):383–405, February 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401678X.

Bloemen:1997:JST

[Blo97] Hans G. Bloemen. Job search theory, labour supply and unemployment duration. Journal of Econometrics, 79(2):305–325, August 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000250.

Beaulieu:1993:SUR

[BM93a] J. Joseph Beaulieu and Jeffrey A. Miron. Seasonal unit roots in aggregate U.S. data. Journal of Econometrics, 55(1–2):305–328, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390018Z.

Braun:1993:MVA

[BM93b] Phillip A. Braun and Stefan Mittnik. Misspecifications in vector autoregressions and their effects on impulse responses and variance decompositions. Journal of Econometrics, 59(3):319– 341, October 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900295.

Berndt:1995:HTC

[BM95] Ernst R. Berndt and Catherine J. Morrison. High-tech capital formation and economic performance in U.S. manufacturing industries an exploratory analysis. Journal of Econometrics, 65(1): 9–43, January 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401596R.

Bollerslev:1996:MPL

[BM96] Tim Bollerslev and Hans Ole Mikkelsen. Modeling and pricing long memory in stock market volatility. Journal of Econometrics, 73(1): 151–184, July 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017364.

Bloom:1997:DAE

[BM97] David E. Bloom and Ajay S. Mahal. Does the AIDS epidemic threaten economic growth? Journal of Econometrics, 77(1):105– 124, March 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696018088.

Breslaw:1998:SLV

[BM98] Jon A. Breslaw and James McIntosh. Simulated latent variable estimation of models with ordered categorical data. Journal of Econometrics, 87(1):25–47, November 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000049.

Banerjee:1999:SOW

[BM99a] Anurag N. Banerjee and Jan R. Magnus. The sensitivity of OLS when the variance matrix is (partially) unknown. Journal of Econometrics, 92(2):295–323, October 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000931.

Bollerslev:1999:LTE

[BM99b] Tim Bollerslev and Hans Ole Mikkelsen. Long-term equity anticipation securities and stock market volatility dynamics. Journal of Econometrics, 92(1):75–99, September 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000864.

Blundell:1993:LSI

[BMN93] Richard Blundell, Costas Meghir, and Pedro Neves. Labour supply and intertemporal substitution. Journal of Econometrics, 59(1–2):137–160, September 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407693900435.

Billio:1999:BES

[BMR99] M. Billio, A. Monfort, and C. P. Robert. Bayesian estimation of switching ARMA models. Journal of Econometrics, 93(2):229– 255, December 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769900010X.

Bockenholt:1998:MIP

[Böc98] Ulf Böckenholt. Mixed INAR(1) Poisson regression models: Analyzing heterogeneity and serial dependencies in longitudinal count data. Journal of Econometrics, 89(1–2):317–338, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000694.

Bollinger:1996:BMR

[Bol96] Christopher R. Bollinger. Bounding mean regressions when a binary regressor is mismeasured. Journal of Econometrics, 73(2):387–399, August 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407695017305.

Boswijk:1994:TUR

[Bos94] H. Peter Boswijk. Testing for an unstable root in conditional and structural error correction models. Journal of Econometrics, 63 (1):37–60, July 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693015609.

Boswijk:1995:CSE

[Bos95a] H. Peter Boswijk. Conditional and structural error correction models reply. Journal of Econometrics, 69(1):173–175, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401667O.

Boswijk:1995:EIC

[Bos95b] H. Peter Boswijk. Efficient inference on cointegration parameters in structural error correction models. Journal of Econometrics, 69(1):133–158, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401665M.

Boumans:1995:FTB

[Bou95] Marcel Boumans. Frisch on testing of business cycle theories. Journal of Econometrics, 67(1):129–147, May 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401630I.

Bewley:1994:CBT

[BOYF94] Ronald Bewley, David Orden, Minxian Yang, and Lance A. Fisher. Comparison of Box–Tiao and Johansen canonical estimators of cointegrating vectors in VEC(1) models. Journal of Econometrics, 64(1–2):3–27, September/October 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900558.

Bougerol:1992:SGP

[BP92] Philippe Bougerol and Nico Picard. Stationarity of GARCH processes and of some nonnegative time series. Journal of Econometrics, 52(1–2):115–127, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900672.

Bauwens:1996:EI

[BPvD96] Luc Bauwens, Wolfgang Polasek, and Herman K. van Dijk. Editor’s introduction. Journal of Econometrics, 75(1):1–5, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769501764X.

Breusch:1999:RMC

[BQSW99] Trevor Breusch, Hailong Qian, Peter Schmidt, and Donald Wyhowski. Redundancy of moment conditions. Journal of Econometrics, 91(1):89–111, July 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000505.

Bresnahan:1991:EMD

[BR91] Timothy F. Bresnahan and Peter C. Reiss. Empirical models of discrete games. Journal of Econometrics, 48(1–2):57–81, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407691900329.

Bidani:1997:DSI

[BR97] Benu Bidani and Martin Ravallion. Decomposing social indicators using distributional data. Journal of Econometrics, 77(1):125– 139, March 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407695018093.

Brownstone:1990:BIE

[Bro90] David Brownstone. Bootstrapping improved estimators for linear regression models. Journal of Econometrics, 44(1–2):171–187, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900788.

Brooks:1993:APO

[Bro93] Robert D. Brooks. Alternative point-optimal tests for regression coefficient stability. Journal of Econometrics, 57(1–3):365–376, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390071C.

Barnett:1990:DDD

[BS90a] William A. Barnett and Apostolos Serletis. A dispersiondependency diagnostic test for aggregation error: With applications to monetary economics and income distribution. Journal of Econometrics, 43(1–2):5–34, January/February 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690901042.

Borsch-Supan:1990:CNL

[BS90b] Axel Börsch-Supan. On the compatibility of nested logit models with utility maximization. Journal of Econometrics, 43(3):373– 388, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090126E.

Blundell:1994:CES

[BS94a] Richard Blundell and Richard J. Smith. Coherency and estimation in simultaneous models with censored or qualitative dependent variables. Journal of Econometrics, 64(1–2):355–373, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900698.

Borjas:1994:TSE

[BS94b] George J. Borjas and Glenn T. Sueyoshi. A two-stage estimator for probit models with structural group effects. Journal of Econometrics, 64(1–2):165–182, September/October 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900620.

Borsch-Supan:1993:SUM

[BSH93] Axel Börsch-Supan and Vassilis A. Hajivassiliou. Smooth unbiased multivariate probability simulators for maximum likelihood estimation of limited dependent variable models. Journal of Econometrics, 58(3):347–368, August 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390049B.

Behrman:1991:BWM

[BSTY91] Jere R. Behrman, Robin Sickles, Paul Taubman, and Abdo Yazbeck. Black-white mortality inequalities. Journal of Econometrics, 50(1–2):183–203, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190095U.

Bresnahan:1995:GPT

[BT95] Timothy F. Bresnahan and M. Trajtenberg. General purpose technologies ‘engines of growth’? Journal of Econometrics, 65(1):83– 108, January 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401598T.

Brownstone:1998:FNP

[BT98] David Brownstone and Kenneth Train. Forecasting new product penetration with flexible substitution patterns. Journal of Econometrics, 89(1–2):109–129, November 26, 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000578.

Buchinsky:1995:EAC

[Buc95a] Moshe Buchinsky. Estimating the asymptotic covariance matrix for quantile regression models a Monte Carlo study. Journal of Econometrics, 68(2):303–338, August 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401652G.

Buchinsky:1995:QRB

[Buc95b] Moshe Buchinsky. Quantile regression, Box–Cox transformation model, and the U.S. wage structure, 1963–1987. Journal of Econometrics, 65(1):109–154, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401599U.

Bell:1993:ESE

[BW93] William R. Bell and David W. Wilcox. The effect of sampling error on the time series behavior of consumption data. Journal of Econometrics, 55(1–2):235–265, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390014V.

Brown:1995:SSR

[BW95] Bryan W. Brown and Mary Beth Walker. Stochastic specification in random production models of cost-minimizing firms. Journal of Econometrics, 66(1–2):175–205, March/April 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016146.

Bagozzi:1998:RME

[BYN98] Richard P. Bagozzi, Youjae Yi, and Kent D. Nassen. Representation of measurement error in marketing variables: Review of approaches and extension to three-facet designs. Journal of Econometrics, 89(1–2):393–421, November 26, 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000682.

Chesher:1991:FSD

[CA91] Andrew Chesher and Gerard Austin. The finite-sample distributions of heteroskedasticity robust Wald statistics. Journal of Econometrics, 47(1):153–173, January 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190082O.

Choi:1999:TNS

[CA99] In Choi and Byung Chul Ahn. Testing the null of stationarity for multiple time series. Journal of Econometrics, 88(1):41– 77, January 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000219.

Cameron:1992:IGC

[Cam92] Trudy Ann Cameron. The impact of grouping coarseness in alternative grouped-data regression models. Journal of Econometrics, 52(3):419–421, June 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769290022J.

Canova:1993:FTS

[Can93] Fabio Canova. Forecasting time series with common seasonal patterns. Journal of Econometrics, 55(1–2):173–200, January/ February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900103.

Caner:1998:TCI

[Can98] Mehmet Caner. Tests for cointegration with infinite variance errors. Journal of Econometrics, 86(1):155–175, September 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697001127.

Cartwright:1995:PE

[Car95] Nancy Cartwright. Probabilities and experiments. Journal of Econometrics, 67(1):47–59, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401626B.

Caudill:1992:MGC

[Cau92] Steven B. Caudill. More on grouping coarseness in linear normal regression models. Journal of Econometrics, 52(3):407–417, June 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290021I.

Charnes:1990:PCR

[CCHS90] A. Charnes, W. W. Cooper, Z. M. Huang, and D. B. Sun. Polyhedral cone-ratio DEA models with an illustrative application to large commercial banks. Journal of Econometrics, 46(1–2):73–91, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090048X.

Chiang:1998:MCM

[CCN98] Jeongwen Chiang, Siddhartha Chib, and Chakravarthi Narasimhan. Markov chain Monte Carlo and models of consideration set and parameter heterogeneity. Journal of Econometrics, 89(1–2):223–248, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000621.

Chow:1993:SMV

[CD93] K. Victor Chow and Karen C. Denning. A simple multiple variance ratio test. Journal of Econometrics, 58(3):385–401, August 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693900516.

Cheung:1994:MLE

[CD94] Yin-Wong Cheung and Francis X. Diebold. On maximum likelihood estimation of the differencing parameter of fractionallyintegrated noise with unknown mean. Journal of Econometrics, 62(2):301–316, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900264.

Cragg:1997:IRM

[CD97a] John G. Cragg and Stephen G. Donald. Inferring the rank of a matrix. Journal of Econometrics, 76(1–2):223–250, January/ February 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017909.

Crepon:1997:RDC

[CD97b] Bruno Crepon and Emmanuel Duguet. Research and development, competition and innovation pseudo-maximum likelihood and simulated maximum likelihood methods applied to count data models with heterogeneity. Journal of Econometrics, 79(2):355– 378, August 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000274.

Cooley:1998:BCA

[CD98] Thomas F. Cooley and Mark Dwyer. Business cycle analysis without much theory a look at structural VARs. Journal of Econometrics, 83(1–2):57–88, March/April 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000651.

Campos:1990:AMP

[CEH90] Julia Campos, Neil R. Ericsson, and David F. Hendry. An analogue model of phase-averaging procedures. Journal of Econometrics, 43(3):275–292, March 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407690901219.

Campos:1996:CTP

[CEH96] Julia Campos, Neil R. Ericsson, and David F. Hendry. Cointegration tests in the presence of structural breaks. Journal of Econometrics, 70(1):187–220, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016895.

Chou:1992:MRA

[CEK92] Ray Chou, Robert F. Engle, and Alex Kane. Measuring risk aversion from excess returns on a stock index. Journal of Econometrics, 52(1–2):201–224, April/May 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900708.

Chen:1999:CHT

[CF99] Xiaohong Chen and Yanqin Fan. Consistent hypothesis testing in semiparametric and nonparametric models for econometric time series. Journal of Econometrics, 91(2):373–401, August 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000815.

Chambers:1998:UDF

[CFJL98] Robert G. Chambers, Rolf Färe, Edward Jaenicke, and Erik Lichtenberg. Using dominance in forming bounds on DEA models: The case of experimental agricultural data. Journal of Econometrics, 85(1):189–203, July 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697001000.

Chib:1994:BIR

[CG94] Siddhartha Chib and Edward Greenberg. Bayes inference in regression models with ARMA ( p, q ) errors. Journal of Econometrics, 64(1–2):183–206, September/October 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900639.

Chib:1995:HAM

[CG95] Siddhartha Chib and Edward Greenberg. Hierarchical analysis of SUR models with extensions to correlated serial errors and timevarying parameter models. Journal of Econometrics, 68(2):339– 360, August 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401653H.

Chib:1998:PSB

[CGW98] Siddhartha Chib, Edward Greenberg, and Rainer Winkelmann. Posterior simulation and Bayes factors in panel count data models. Journal of Econometrics, 86(1):33–54, September 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697001085.

Chambers:1990:FDS

[Cha90] Marcus J. Chambers. Forecasting with demand systems: A comparative study. Journal of Econometrics, 44(3):363–376, June 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090064Z.

Chambers:1993:NAT

[Cha93] Marcus J. Chambers. A nonnested approach to testing continuous time models against discrete alternatives. Journal of Econometrics, 57(1–3):319–343, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390069H.

Chambers:1998:ESJ

[Cha98] Marcus J. Chambers. The estimation of systems of joint differential-difference equations. Journal of Econometrics, 85(1): 1–31, July 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000912.

Chen:1997:SET

[Che97] Songnian Chen. Semiparametric estimation of the Type-3 Tobit model. Journal of Econometrics, 80(1):1–34, September 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696000115.

Chen:1999:DFE

[Che99] Songnian Chen. Distribution-free estimation of the random coefficient dummy endogenous variable model. Journal of Econometrics, 91(1):171–199, July 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769800075X.

Chib:1992:BIT

[Chi92] Siddhartha Chib. Bayes inference in the Tobit censored regression model. Journal of Econometrics, 51(1–2):79–99, January/ February 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290030U.

Chib:1993:BRA

[Chi93] Siddhartha Chib. Bayes regression with autoregressive errors: A Gibbs sampling approach. Journal of Econometrics, 58(3):275– 294, August 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900468.

Chirinko:1995:NLH

[Chi95] Robert S. Chirinko. Nonconvexities, labor hoarding, technology shocks, and procyclical productivity a structural econometric analysis. Journal of Econometrics, 66(1–2):61–98, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694016083.

Chib:1996:CPD

[Chi96] Siddhartha Chib. Calculating posterior distributions and modal estimates in Markov mixture models. Journal of Econometrics, 75(1):79–97, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017704.

Chib:1998:ECM

[Chi98] Siddhartha Chib. Estimation and comparison of multiple changepoint models. Journal of Econometrics, 86(2):221–241, October 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697001152.

Choi:1994:SRR

[Cho94] In Choi. Spurious regressions and residual-based tests for cointegration when regressors are cointegrated. Journal of Econometrics, 60(1–2):313–320, January/February 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900493.

Chung:1996:EGL

[Chu96] Ching-Fan Chung. Estimating a generalized long memory process. Journal of Econometrics, 73(1):237–259, July 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017399.

Christensen:1997:INL

[CK97] Bent Jesper Christensen and Nicholas M. Kiefer. Inference in nonlinear panel models with partially missing observations the case of the equilibrium search model. Journal of Econometrics, 79(2):201– 219, August 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696000206.

Crepon:1997:PIN

[CKT97] Bruno Crepon, Francis Kramarz, and Alain Trognon. Parameters of interest, nuisance parameters and orthogonality conditions an application to autoregressive error component models. Journal of Econometrics, 82(1):135–156, ???? 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000547.

Chiang:1992:DCM

[CL92] Jeongwen Chiang and Lung-Fei Lee. Discrete/continuous models of consumer demand with binding nonnegativity constraints. Journal of Econometrics, 54(1–3):79–93, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407692901006.

Card:1996:WDR

[CL96] David Card and Thomas Lemieux. Wage dispersion, returns to skill, and black-white wage differentials. Journal of Econometrics, 74(2):319–361, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017577.

Campbell:1990:EI

[CM90] John Y. Campbell and Angelo Melino. Editors’ introduction. Journal of Econometrics, 45(1–2):1–5, July/August 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690900917.

Choi:1997:GEV

[CM97] Ki-Hong Choi and Choon-Geol Moon. Generalized extreme value model and additively separable generator function. Journal of Econometrics, 76(1–2):129–140, January/February 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017852.

Cheung:1996:CVT

[CN96] Yin-Wong Cheung and Lilian K. Ng. A causality-in-variance test and its application to financial market prices. Journal of Econometrics, 72(1–2):33–48, May/June 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401714X.

Collado:1997:EDM

[Col97] M. Dolores Collado. Estimating dynamic models from time series of independent cross-sections. Journal of Econometrics, 82(1): 37–62, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000158.

Comte:1999:DCT

[Com99] F. Comte. Discrete and continuous time cointegration. Journal of Econometrics, 88(2):207–226, February 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000256.

Connolly:1991:POA

[Con91] Robert A. Connolly. A posterior odds analysis of the weekend effect. Journal of Econometrics, 49(1–2):51–104, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190010B.

Conley:1999:GEC

[Con99] T. G. Conley. GMM estimation with cross sectional dependence. Journal of Econometrics, 92(1):1–45, September 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000840.

Corts:1999:CPM

[Cor99] Kenneth S. Corts. Conduct parameters and the measurement of market power. Journal of Econometrics, 88(2):227–250, February 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000281.

Cowell:1991:GBI

[Cow91] Frank A. Cowell. Grouping bounds for inequality measures under alternative informational assumptions. Journal of Econometrics, 48(1–2):1–14, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190029D.

Choi:1992:AFS

[CP92] In Choi and Peter C. B. Phillips. Asymptotic and finite sample distribution theory for IV estimators and tests in partially identified structural equations. Journal of Econometrics, 51(1–2):113–150, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290032M.

Choi:1993:TUR

[CP93] In Choi and Peter C. B. Phillips. Testing for a unit root by frequency domain regression. Journal of Econometrics, 59(3):263– 286, October 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900262.

Chao:1998:PDL

[CP98] J. C. Chao and P. C. B. Phillips. Posterior distributions in limited information analysis of the simultaneous equations model using the Jeffreys prior. Journal of Econometrics, 87(1):49–86, November 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000062.

Chao:1999:MSP

[CP99] John C. Chao and Peter C. B. Phillips. Model selection in partially nonstationary vector autoregressive processes with reduced rank structure. Journal of Econometrics, 91(2):227–271, August 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000773.

Carraro:1993:EIE

[CPW93] Carlo Carraro, Franco Peracchi, and Guglielmo Weber. Editors’ introduction: The econometrics of panels and pseudo panels. Journal of Econometrics, 59(1–2):1–4, September 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900365.

Cramer:1991:PSM

[CR91] J. S. Cramer and G. Ridder. Pooling states in the multinomial logit model. Journal of Econometrics, 47(2–3):267–272, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190102J. See erratum [Ano92b].

Cheng:1994:SET

[CR94a] Bing Cheng and P. M. Robinson. Semiparametric estimation from time series with long-range dependence. Journal of Econometrics, 64(1–2):335–353, September/October 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769490068X.

Conway:1994:AED

[CR94b] Delores A. Conway and Harry V. Roberts. Analysis of employment discrimination through homogeneous job groups. Journal of Econometrics, 61(1):103–131, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900795.

Comte:1996:LMC

[CR96] F. Comte and E. Renault. Long memory continuous time models. Journal of Econometrics, 73(1):101–149, July 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017356.

Chen:1997:SNE

[CR97] Heng Z. Chen and Alan Randall. Semi-nonparametric estimation of binary response models with an application to natural resource valuation. Journal of Econometrics, 76(1–2):323–340, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017941.

Cragg:1992:QAE

[Cra92] John G. Cragg. Quasi-Aitken estimation for heteroskedasticity of unknown form. Journal of Econometrics, 54(1–3):179–201, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290105Z.

Creedy:1990:MWS

[Cre90] John Creedy. Measuring wealth in a simple two-period model. Journal of Econometrics, 43(1–2):167–177, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690901149.

Crockett:1998:REI

[Cro98] Jean A. Crockett. Rational expectations, inflation and the nominal interest rate. Journal of Econometrics, 83(1–2):349–363, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000754.

Choudhury:1990:NPH

[CS90] Askar H. Choudhury and Robert D. St. Louis. A note on Park and Heikes’ (1983) modified approximate estimator for the first-order moving-average process. Journal of Econometrics, 46(3):399–406, December 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090017N.

Cavanagh:1998:REM

[CS98] Christopher Cavanagh and Robert P. Sherman. Rank estimators for monotonic index models. Journal of Econometrics, 84(2): 351–381, June 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000900.

Cornwell:1990:PFC

[CSS90] Christopher Cornwell, Peter Schmidt, and Robin C. Sickles. Production frontiers with cross-sectional and time-series variation in efficiency levels. Journal of Econometrics, 46(1–2):185–200, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090054W.

Cornwell:1992:SEP

[CSW92] Christopher Cornwell, Peter Schmidt, and Donald Wyhowski. Simultaneous equations and panel data. Journal of Econometrics, 51(1–2):151–181, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290033N.

Cameron:1990:RBT

[CT90] A. Colin Cameron and Pravin K. Trivedi. Regression-based tests for overdispersion in the Poisson model. Journal of Econometrics, 46(3):347–364, December 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090014K.

Cameron:1997:SMG

[CW97] A. Colin Cameron and Frank A. G. Windmeijer. An R-squared measure of goodness of fit for some common nonlinear regression models. Journal of Econometrics, 77(2):329–342, April 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696018180.

Dagum:1990:RBI

[Dag90] Camilo Dagum. On the relationship between income inequality measures and social welfare functions. Journal of Econometrics, 43(1–2):91–102, January/February 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690901097.

Dagenais:1994:PER

[Dag94] Marcel G. Dagenais. Parameter estimation in regression models with errors in the variables and autocorrelated disturbances. Journal of Econometrics, 64(1–2):145–163, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900612.

Danielsson:1994:SVA

[Dan94] Jon Danielsson. Stochastic volatility in asset prices estimation with simulated maximum likelihood. Journal of Econometrics, 64(1–2):375–400, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900701.

Das:1991:SSA

[Das91] Sanghamitra Das. A semiparametric structural analysis of the idling of cement kilns. Journal of Econometrics, 50(3):235–256, December 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190020E.

Dasgupta:1997:NSC

[Das97a] Partha Dasgupta. Nutritional status, the capacity for work, and poverty traps. Journal of Econometrics, 77(1):5–37, March 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696018040.

Dastoor:1997:TCH

[Das97b] Naorayex K. Dastoor. Testing for conditional heteroskedasticity with misspecified alternative hypotheses. Journal of Econometrics, 82(1):63–80, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697889567.

Davidson:1998:SRC

[Dav98] James Davidson. Structural relations, cointegration and identification: some simple results and their application. Journal of Econometrics, 87(1):87–113, November 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000074.

Diebold:1996:TSS

[DC96] Francis X. Diebold and Celia Chen. Testing structural stability with endogenous breakpoint a size comparison of analytic and bootstrap procedures. Journal of Econometrics, 70(1):221– 241, January 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694016909.

DeSarbo:1998:LSD

[DC98] Wayne S. DeSarbo and Jungwhan Choi. A latent structure double hurdle regression model for exploring heterogeneity in consumer search patterns. Journal of Econometrics, 89(1–2):423–455, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000700.

Dagenais:1997:HME

[DD97] Marcel G. Dagenais and Denyse L. Dagenais. Higher moment estimators for linear regression models with errors in the variables. Journal of Econometrics, 76(1–2):193–221, January/ February 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017895.

Deaton:1990:PES

[Dea90] Angus Deaton. Price elasticities from survey data: Extensions and Indonesian results. Journal of Econometrics, 44(3):281–309, June 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690900607.

DeJong:1992:CIT

[DeJ92] David N. DeJong. Co-integration and trend-stationarity in macroeconomic time series: Evidence from the likelihood function. Journal of Econometrics, 52(3):347–370, June 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290016K.

Deschamps:1998:FML

[Des98] Philippe J. Deschamps. Full maximum likelihood estimation of dynamic demand models. Journal of Econometrics, 82(2):335– 359, February 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697815746.

Deutsch:1994:MOS

[DFS94] Joseph Deutsch, Yves Flückiger, and Jacques Silber. Measuring occupational segregation: Summary statistics and the impact of classification errors and aggregation. Journal of Econometrics, 61(1):133–146, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900809.

Ding:1996:MVP

[DG96a] Zhuanxin Ding and Clive W. J. Granger. Modeling volatility persistence of speculative returns: A new approach. Journal of Econometrics, 73(1):185–215, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017372.

Dufour:1996:EIR

[DG96b] Jean-Marie Dufour and Eric Ghysels. Editors’ introduction recent developments in the econometrics of structural change. Journal of Econometrics, 70(1):1–8, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401681X.

Dionne:1997:DMH

[DGGV97] Georges Dionne, Robert Gagné, François Gagnon, and Charles Vanasse. Debt, moral hazard and airline safety an empirical evidence. Journal of Econometrics, 79(2):379–402, August 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697829892.

Dionne:1998:ITP

[DGV98] Georges Dionne, Robert Gagné, and Charles Vanasse. Inferring technological parameters from incomplete panel data. Journal of Econometrics, 87(2):303–327, December 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000025.

Dorfman:1992:BAS

[DH92] Jeffrey H. Dorfman and Arthur M. Havenner. A Bayesian approach to state space multivariate time series modeling. Journal of Econometrics, 52(3):315–346, June 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290015J.

Dhrymes:1994:STS

[Dhr94] Phoebus J. Dhrymes. Specification tests in simultaneous equations systems. Journal of Econometrics, 64(1–2):45–76, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900574.

Dhrymes:1998:IKI

[Dhr98] Phoebus J. Dhrymes. Identification and Kullback information in the GLSEM. Journal of Econometrics, 83(1–2):163–184, March/ April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000687.

Dekimpe:1998:LRE

[DHSR98] Marnik G. Dekimpe, Dominique M. Hanssens, and Jorge M. Silva-Risso. Long-run effects of price promotions in scanner markets. Journal of Econometrics, 89(1–2):269–291, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000645.

Dickey:1993:SUR

[Dic93] David A. Dickey. Seasonal unit roots in aggregate U.S. data. Journal of Econometrics, 55(1–2):329–331, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693900192.

Diebold:1993:DES

[Die93] Francis X. Diebold. Discussion: The effect of seasonal adjustment filters on tests for a unit root. Journal of Econometrics, 55(1–2):99–103, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390005P.

deJong:1996:BTU

[dJ96] Robert M. de Jong. The bierens test under data dependence. Journal of Econometrics, 72(1–2):1–32, May/June 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017123.

deJong:1998:ULL

[dJ98] Robert M. de Jong. Uniform laws of large numbers and stochastic Lipschitz-continuity. Journal of Econometrics, 86(2):243–268, October 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697001164.

Dufour:1991:OIT

[DK91] Jean-Marie Dufour and Maxwell L. King. Optimal invariant tests for the autocorrelation coefficient in linear regressions with stationary or nonstationary AR(1) errors. Journal of Econometrics, 47(1):115–143, January 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190080W.

Dufour:1996:ETS

[DK96] Jean-Marie Dufour and Jan F. Kiviet. Exact tests for structural change in first-order dynamic models. Journal of Econometrics, 70(1):39–68, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016836.

Drost:1997:EES

[DK97] Feike C. Drost and Chris A. J. Klaassen. Efficient estimation in semiparametric GARCH models. Journal of Econometrics, 81(1):193–221, November 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000420.

DeSarbo:1998:BMS

[DKF98] Wayne S. DeSarbo, Youngchan Kim, and Duncan Fong. A Bayesian multidimensional scaling procedure for the spatial analysis of revealed choice data. Journal of Econometrics, 89(1– 2):79–108, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407698000566.

Day:1992:SMV

[DL92] Theodore E. Day and Craig M. Lewis. Stock market volatility and the information content of stock index options. Journal of Econometrics, 52(1–2):267–287, April/May 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290073Z.

Davies:1995:NFA

[DL95] Anthony Davies and Kajal Lahiri. A new framework for analyzing survey forecasts using three-dimensional panel data. Journal of Econometrics, 68(1):205–227, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401649K.

deLima:1997:RNT

[dL97a] Pedro J. F. de Lima. On the robustness of nonlinearity tests to moment condition failure. Journal of Econometrics, 76(1–2):251–280, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017917.

Diebold:1997:WEA

[DL97b] Francis X. Diebold and Russell L. Lamb. Why are estimates of agricultural supply response so variable? Journal of Econometrics, 76(1–2):357–373, January/February 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017976.

Davidson:1992:RBM

[DM92] Russell Davidson and James G. MacKinnon. Regression-based methods for using control variates in Monte Carlo experiments. Journal of Econometrics, 54(1–3):203–222, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407692901062.

DeJong:1992:PPU

[DNSW92] David N. DeJong, John C. Nankervis, N. E. Savin, and Charles H. Whiteman. The power problems of unit root test in time series with autoregressive errors. Journal of Econometrics, 53(1–3):323–343, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290090E.

Donald:1995:TSE

[Don95] Stephen G. Donald. Two-step estimation of heteroskedastic sample selection models. Journal of Econometrics, 65(2):347–380, February 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769301590I.

Dorfman:1995:NBT

[Dor95] Jeffrey H. Dorfman. A numerical Bayesian test for cointegration of AR processes. Journal of Econometrics, 66(1–2):289–324, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401619B.

Dagum:1993:DLM

[DQ93] Estela Bee Dagum and Benoît Quenneville. Dynamic linear models for time series components. Journal of Econometrics, 55(1–2):333–351, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900206.

Dubin:1993:EEI

[DR93] Jeffrey A. Dubin and Douglas Rivers. Experimental estimates of the impact of wage subsidies. Journal of Econometrics, 56(1–2): 219–242, March 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390107G.

Doran:1997:ALT

[DR97] Howard E. Doran and Alicia N. Rambaldi. Applying linear timevarying constraints to econometric models: With an application to demand systems. Journal of Econometrics, 79(1):83–95, July 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000080.

Demos:1998:TGE

[DS98] Antonis Demos and Enrique Sentana. Testing for GARCH effects: a one-sided approach. Journal of Econometrics, 86(1):97– 127, September 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697001103.

Duggal:1999:IPN

[DSK99] Vijaya G. Duggal, Cynthia Saltzman, and Lawrence R. Klein. Infrastructure and productivity: a nonlinear approach. Journal of Econometrics, 92(1):47–74, September 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000852.

Diamond:1990:MPM

[DSW90] Charles A. Diamond, Curtis J. Simon, and John T. Warner. A multinomial probability model of size income distribution. Journal of Econometrics, 43(1–2):43–61, January/February 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690901064.

Dezhbakhsh:1994:TAP

[DT94] Hashem Dezhbakhsh and Jerry G. Thursby. Testing for autocorrelation in the presence of lagged dependent variables: A specification error approach. Journal of Econometrics, 60(1–2):251–272, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900469.

Duan:1997:AGP

[Dua97] Jin-Chuan Duan. Augmented GARCH ( p, q ) process and its diffusion limit. Journal of Econometrics, 79(1):97–127, July 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000092.

Dufour:1993:ISI

[Duf93] Jean-Marie Dufour. The importance of seasonality in inventory models. Journal of Econometrics, 55(1–2):129–133, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390007R.

Durlauf:1991:SBT

[Dur91] Steven N. Durlauf. Spectral based testing of the martingale hypothesis. Journal of Econometrics, 50(3):355–376, December 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407691900259.

Durlauf:1993:TSP

[Dur93] Steven N. Durlauf. Time series properties of aggregate output fluctuations. Journal of Econometrics, 56(1–2):39–56, March 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390100J.

deVries:1991:RBG

[dV91] Casper G. de Vries. On the relation between GARCH and stable processes. Journal of Econometrics, 48(3):313–324, June 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190066M.

Diewert:1995:FFF

[DW95] W. E. Diewert and T. J. Wales. Flexible functional forms and tests of homogeneous separability. Journal of Econometrics, 67(2): 259–302, June 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694016179.

Drost:1996:CGG

[DW96] Feike C. Drost and Bas J. M. Werker. Closing the GARCH gap: Continuous time GARCH modeling. Journal of Econometrics, 74(1):31–57, September 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769501750X.

Eastwood:1991:ANC

[Eas91] Brian J. Eastwood. Asymptotic normality and consistency of seminonparametric regression estimators using an upwards f test truncation rule. Journal of Econometrics, 48(1–2):151–181, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190036D.

Ermini:1996:TJH

[EC96] Luigi Ermini and Dongkoo Chang. Testing the joint hypothesis of rationality and neutrality under seasonal cointegration: The case of Korea. Journal of Econometrics, 74(2):363–386, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017585.

Ermini:1993:SGA

[EG93] Luigi Ermini and Clive W. J. Granger. Some generalizations on the algebra of i(1) processes. Journal of Econometrics, 58(3):369– 384, August 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390050F.

Engle:1993:JCF

[EGHL93] R. F. Engle, C. W. J. Granger, S. Hylleberg, and H. S. Lee. The Japanese consumption function. Journal of Econometrics, 55(1–2):275–298, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390016X.

Enberg:1990:REL

[EGW90] John Enberg, Peter Gottschalk, and Douglas Wolf. A randomeffects logit model of work-welfare transitions. Journal of Econometrics, 43(1–2):63–75, January/February 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690901075.

Engle:1993:TSI

[EH93] Robert F. Engle and David F. Hendry. Testing superexogeneity and invariance in regression models. Journal of Econometrics, 56(1–2):119–139, March 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390103C.

Erdem:1998:MPC

[EKS98] Tülin Erdem, Michael P. Keane, and Baohong Sun. Missing price and coupon availability data in scanner panels: Correcting for the self-selection bias in choice model parameters. Journal of Econometrics, 89(1–2):177–196, November 26, 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000608.

Engle:1992:IAM

[EM92] Robert F. Engle and Chowdhury Mustafa. Implied ARCH models from options prices. Journal of Econometrics, 52(1–2):289–311, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692900742.

Ebrahimi:1999:OUD

[EMS99] Nader Ebrahimi, Esfandiar Maasoumi, and Ehsan S. Soofi. Ordering univariate distributions by entropy and variance. Journal of Econometrics, 90(2):317–336, June 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000463.

Engle:1990:APF

[ENR90] Robert F. Engle, Victor K. Ng, and Michael Rothschild. Asset pricing with a factor-arch covariance structure: Empirical estimates for treasury bills. Journal of Econometrics, 45 (1–2):213–237, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090099F.

Entorf:1997:RWD

[Ent97] Horst Entorf. Random walks with drifts: Nonsense regression and spurious fixed-effect estimation. Journal of Econometrics, 80(2):287–296, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000419.

Erard:1997:SSM

[Era97] Brian Erard. Self-selection with measurement errors a microeconometric analysis of the decision to seek tax assistance and its implications for tax compliance. Journal of Econometrics, 81(2):319– 356, December 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697865700.

Ericsson:1995:CSE

[Eri95] Neil R. Ericsson. Conditional and structural error correction models. Journal of Econometrics, 69(1):159–171, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401666N.

Eitrheim:1996:TAS

[ET96] Øyvind Eitrheim and Timo Teräsvirta. Testing the adequacy of smooth transition autoregressive models. Journal of Econometrics, 74(1):59–75, September 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017518.

Evans:1992:RST

[Eva92] Merran Evans. Robustness of size of tests of autocorrelation and heteroscedasticity to nonnormality. Journal of Econometrics, 51(1–2):7–24, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290027O.

Erdem:1998:EMC

[EW98] Tülin Erdem and Russell S. Winer. Econometric modeling of competition: A multi-category choice-based mapping approach. Journal of Econometrics, 89(1–2):159–175, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000591.

Fiebig:1991:RCA

[FBA91] Denzil G. Fiebig, Robert Bartels, and Dennis J. Aigner. A random coefficient approach to the estimation of residential end-use load profiles. Journal of Econometrics, 50(3):297–327, December 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407691900237.

Franses:1997:RCS

[FD97] Philip Hans Franses and Gerrit Draisma. Recognizing changing seasonal patterns using artificial neural networks. Journal of Econometrics, 81(1):273–280, November 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769700047X.

Fry:1996:SSD

[FFM96] Jane M. Fry, Tim R. L. Fry, and Keith R. McLaren. The stochastic specification of demand share equations: Restricting budget shares to the unit simplex. Journal of Econometrics, 73(2):377– 385, August 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407695017275.

Fare:1995:NTR

[FG95] Rolf Färe and Shawna Grosskopf. Nonparametric tests of regularity, Farrell efficiency, and goodness-of-fit. Journal of Econometrics, 69(2):415–425, October 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401677R.

Fenton:1996:QAP

[FG96] Victor M. Fenton and A. Ronald Gallant. Qualitative and asymptotic performance of SNP density estimators. Journal of Econometrics, 74(1):77–118, September 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017526.

Filardo:1998:BCD

[FG98] Andrew J. Filardo and Stephen F. Gordon. Business cycle durations. Journal of Econometrics, 85(1):99–123, July 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000961.

Follmann:1990:PCU

[FGM90] Dean A. Follmann, Matthew S. Goldberg, and Laurie May. Personal characteristics, unemployment insurance, and the duration of unemployment. Journal of Econometrics, 45(3):351–366, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090004D.

Fomby:1990:IAW

[FH90]  Thomas B. Fomby and Kathy J. Hayes.  An interven-

tion analysis of the war on poverty: Poverty’s persistence and political-business cycle implications. Journal of Econometrics, 43(1–2):197–212, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090116B.

Fukushige:1991:ERM

[FH91] Mototsugu Fukushige and Michio Hatanaka. Estimation of a regression model on two or more sets of differently grouped data. Journal of Econometrics, 47(2–3):207–226, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190099Y.

Franses:1997:BASa

[FHP97a] Philip Hans Franses, Henk Hoek, and Richard Paap. Bayesian analysis of seasonal unit roots and seasonal mean shifts. Journal of Econometrics, 78(1):359–380, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800176.

Franses:1997:BASb

[FHP97b] Philip Hans Franses, Henk Hoek, and Richard Paap. Bayesian analysis of seasonal unit roots and seasonal mean shifts. Journal of Econometrics, 78(2):359–380, June 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000183.

Findley:1993:DLM

[Fin93] David F. Findley. Dynamic linear models for time series components. Journal of Econometrics, 55(1–2):353–356, January/ February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390021V.

Fisher:1995:PTM

[Fis95] Franklin M. Fisher. The production-theoretic measurement of input price and quantity indices. Journal of Econometrics, 65(1):155–174, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016005.

Fitzenberger:1998:MBB

[Fit98] Bernd Fitzenberger. The moving blocks bootstrap and robust inference for linear least squares and quantile regressions. Journal of Econometrics, 82(2):235–287, February 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000584.

Friedman:1993:ALE

[FK93] Benjamin M. Friedman and Kenneth N. Kuttner. Another look at the evidence on money-income causality. Journal of Econometrics, 57(1–3):189–203, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390064C.

Froeb:1994:MCS

[FK94] Luke Froeb and Robert Koyak. Measuring and comparing smoothness in time series the production smoothing hypothesis. Journal of Econometrics, 64(1–2):97–122, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900590.

Franses:1998:ORA

[FKL98] Philip Hans Franses, Teun Kloek, and André Lucas. Outlier robust analysis of long-run marketing effects for weekly scanning data. Journal of Econometrics, 89(1–2):293–315, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000657.

Ferrier:1990:MCE

[FL90] Gary D. Ferrier and C. A. Knox Lovell. Measuring cost efficiency in banking: Econometric and linear programming evidence. Journal of Econometrics, 46(1–2):229–245, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090057Z.

Lee:1994:STS

[fL94] Lung fei Lee. Semiparametric two-stage estimation of sample selection models subject to Tobit-type selection rules. Journal of Econometrics, 61(2):305–344, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900884.

Lee:1995:SML

[fL95] Lung fei Lee. Semiparametric maximum likelihood estimation of polychotomous and sequential choice models. Journal of Econometrics, 65(2):381–428, February 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693015919.

Lee:1999:EDA

[fL99] Lung fei Lee. Estimation of dynamic and ARCH Tobit models. Journal of Econometrics, 92(2):355–390, October 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000955.

Lee:1997:EIN

[fLRP97] Lung fei Lee, Mark R. Rosenzweig, and Mark M. Pitt. The effects of improved nutrition, sanitation, and water quality on child health in high-mortality populations. Journal of Econometrics, 77(1): 209–235, March 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696018131.

Foekens:1998:VPM

[FLW98] Eijte W. Foekens, Peter S. H. Leeflang, and Dick R. Wittink. Varying parameter models to accommodate dynamic promotion effects. Journal of Econometrics, 89(1–2):249–268, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000633.

Fiebig:1992:POL

[FMB92] Denzil G. Fiebig, Michael McAleer, and Robert Bartels. Properties of ordinary least squares estimators in regression models with nonspherical disturbances. Journal of Econometrics, 54(1–3):321–334, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692901114.

Fernandez:1997:UPD

[FOS97] Carmen Fernández, Jacek Osiewalski, and Mark F. J. Steel. On the use of panel data in stochastic frontier models with improper priors. Journal of Econometrics, 79(1):169–193, July 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697880505.

Fuss:1995:EI

[FP95] Melvyn Fuss and Ariel Pakes. Editors’ introduction. Journal of Econometrics, 65(1):1–8, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401595Q.

Franses:1994:MAM

[Fra94] Philip Hans Franses. A multivariate approach to modeling univariate seasonal time series. Journal of Econometrics, 63(1):133– 151, July 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693015632.

Frees:1995:ACS

[Fre95] Edward W. Frees. Assessing cross-sectional correlation in panel data. Journal of Econometrics, 69(2):393–414, October 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401658M.

Friedmann:1990:BEM

[Fri90] Ralph Friedmann. Bounds for exact moments of estimators in the errors-in-variables model and simultaneous equations. Journal of Econometrics, 45(3):367–384, ???? 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090005E.

Gil-Alana:1997:TUR

[GAR97] L. A. Gil-Alaña and P. M. Robinson. Testing of unit root and other nonstationary hypotheses in macroeconomic time series. Journal of Econometrics, 80(2):241–268, October 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000389.

Granger:1996:VLM

[GD96] Clive W. J. Granger and Zhuanxin Ding. Varieties of long memory models. Journal of Econometrics, 73(1):61–77, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769501733X.

Geweke:1993:FTS

[Gew93] John Geweke. Forecasting time series with common seasonal patterns. Journal of Econometrics, 55(1–2):201–202, January/ February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390011S.

Geweke:1996:BRR

[Gew96] John Geweke. Bayesian reduced rank regression in econometrics. Journal of Econometrics, 75(1):121–146, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017739. See corrigendum [Kar17].

Goffe:1994:GOS

[GFR94] William L. Goffe, Gary D. Ferrier, and John Rogers. Global optimization of statistical functions with simulated annealing. Journal of Econometrics, 60(1–2):65–99, January/February 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900388.

Ghysels:1998:PTS

[GGH98] Eric Ghysels, Alain Guay, and Alastair Hall. Predictive tests for structural change with unknown breakpoint. Journal of Econometrics, 82(2):209–233, February 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000572.

Ghysels:1990:CBI

[GH90a] Eric Ghysels and Alastair Hall. Are consumption-based intertemporal capital asset pricing models structural? Journal of Econometrics, 45(1–2):121–139, July/August 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090096C.

Ghysels:1990:TNE

[GH90b] Eric Ghysels and Alastair Hall. Testing nonnested Euler conditions with quadrature-based methods of approximation. Journal of Econometrics, 46(3):273–308, December 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090011H.

Gregory:1996:RBT

[GH96] Allan W. Gregory and Bruce E. Hansen. Residual-based tests for cointegration in models with regime shifts. Journal of Econometrics, 70(1):99–126, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407669416857.

Gorgens:1999:SEC

[GH99] Tue Gørgens and Joel L. Horowitz. Semiparametric estimation of a censored regression model with an unknown transformation of the dependent variable. Journal of Econometrics, 90(2):155– 191, June 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000402.

Govaerts:1994:ESL

[GHR94] Bernadette Govaerts, David F. Hendry, and Jean-François Richard. Encompassing in stationary linear dynamic models. Journal of Econometrics, 63(1):245–270, July 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693015676.

Gallant:1990:UCM

[GHT90] A. Ronald Gallant, Lars Peter Hansen, and George Tauchen. Using conditional moments of asset payoffs to infer the volatility of intertemporal marginal rates of substitution. Journal of Econometrics, 45(1–2):141–179, July/August 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090097D.

Gallant:1997:ESV

[GHT97] A. Ronald Gallant, David Hsieh, and George Tauchen. Estimation of stochastic volatility models with diagnostics. Journal of Econometrics, 81(1):159–192, November 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000390.

Ghysels:1993:EIS

[Ghy93] Eric Ghysels. Editor’s introduction: Seasonality and econometric models. Journal of Econometrics, 55(1–2):1–8, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390001L.

Ghysels:1997:SBC

[Ghy97] Eric Ghysels. On seasonality and business cycle durations: A nonparametric investigation. Journal of Econometrics, 79(2):269– 290, August 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000237.

Giles:1991:PTL

[Gil91] Judith A. Giles. Pre-testing for linear restrictions in a regression model with spherically symmetric disturbances. Journal of Econometrics, 50(3):377–398, December 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190026A.

Giles:1992:EEV

[Gil92] Judith A. Giles. Estimation of the error variance after a preliminary-test of homogeneity in a regression model with spherically symmetric disturbances. Journal of Econometrics, 53(1– 3):345–361, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407692900915.

Giles:1993:PTL

[Gil93] Judith A. Giles. Pre-testing for linear restrictions in a regression model with spherically symmetric disturbances (vol. 50, no. 3 (1991) pp. 377-398). Journal of Econometrics, 58(3):403, August 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693900527.

Granger:1997:NST

[GIM97] Clive W. J. Granger, Tomoo Inoue, and Norman Morin. Nonlinear stochastic trends. Journal of Econometrics, 81(1):65–92, November 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000341.

Griffiths:1992:TEL

[GJ92] William Griffiths and George Judge. Testing and estimating location vectors when the error covariance matrix is unknown. Journal of Econometrics, 54(1–3):121–138, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290102W.

Golan:1997:EIC

[GJP97] Amos Golan, George Judge, and Jeffrey Perloff. Estimation and inference with censored and ordered multinomial response data. Journal of Econometrics, 79(1):23–51, July 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000067.

Geweke:1997:SIM

[GKR97] John F. Geweke, Michael P. Keane, and David E. Runkle. Statistical inference in the multinomial multiperiod probit model. Journal of Econometrics, 80(1):125–165, September 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000055.

Gatto:1991:NCS

[GKS91] Joseph P. Gatto, Harry H. Kelejian, and Scott W. Stephan. A note concerning specifications of interactive random-coefficient regression models. Journal of Econometrics, 47(2–3):273–284, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190103K.

Granger:1995:CTE

[GKW95] Clive W. J. Granger, Maxwell L. King, and Halbert White. Comments on testing economic theories and the use of model selection criteria. Journal of Econometrics, 67(1):173–187, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401632A.

Gregoir:1994:PCE

[GL94] Stéphane Gregoir and Guy Laroque. Polynomial cointegration estimation and test. Journal of Econometrics, 63(1):183–214, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693015654.

Gonzalo:1998:PTL

[GL98] Jesús Gonzalo and Tae-Hwy Lee. Pitfalls in testing for long run relationships. Journal of Econometrics, 86(1):129–154, September 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697001115.

Ghysels:1994:TUR

[GLN94] Eric Ghysels, Hahn S. Lee, and Jaesum Noh. Testing for unit roots in seasonal time series: Some theoretical extensions and a Monte Carlo investigation. Journal of Econometrics, 62(2):415– 442, June 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900302.

Gourieroux:1992:QTA

[GM92] Christian Gourieroux and Alain Monfort. Qualitative threshold ARCH models. Journal of Econometrics, 52(1–2):159–199, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692900694.

Gourieroux:1993:SBI

[GM93a] Christian Gourieroux and Alain Monfort. Simulation-based inference: A survey with special reference to panel data models. Journal of Econometrics, 59(1–2):5–33, September 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900376.

Guilkey:1993:ETR

[GM93b] David K. Guilkey and James L. Murphy. Estimation and testing in the random effects probit model. Journal of Econometrics, 59(3):301–317, October 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407693900284.

Gourieroux:1997:DTC

[GM97] Christian Gouriéroux and Thierry Magnac. Duration, transition and count data models introduction. Journal of Econometrics, 79(2):195–199, August 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000195.

Gomez:1999:MOA

[GMP99] Víctor Gómez, Agustín Maravall, and Daniel Peña. Missing observations in ARIMA models: Skipping approach versus additive outlier approach. Journal of Econometrics, 88(2):341–363, February 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000360.

Gregory:1996:TSB

[GNW96] Allan W. Gregory, James M. Nason, and David G. Watt. Testing for structural breaks in cointegrated relationships. Journal of Econometrics, 71(1–2):321–341, March/April 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407696845088.

Godfrey:1996:SRG

[God96] Leslie G. Godfrey. Some results on the glejser and koenker tests for heteroskedasticity. Journal of Econometrics, 72(1–2):275–299, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694017239.

Godfrey:1998:TNN

[God98a] L. G. Godfrey. Tests of non-nested regression models some results on small sample behaviour and the bootstrap. Journal of Econometrics, 84(1):59–74, May 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000791.

Godfrey:1998:HTA

[God98b] Leslie G. Godfrey. Hausman tests for autocorrelation in the presence of lagged dependent variables some further results. Journal of Econometrics, 82(2):197–207, February 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000560.

Gonzalo:1994:FAM

[Gon94] Jesus Gonzalo. Five alternative methods of estimating longrun equilibrium relationships. Journal of Econometrics, 60(1– 2):203–233, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900442.

Gonul:1998:EPE

[Gön98] Füsun F. Gönül. Estimating price expectations in the OTC medicine market: An application of dynamic stochastic discrete choice models to scanner panel data. Journal of Econometrics, 89(1–2):41–56, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000542.

Gozalo:1997:NBA

[Goz97] Pedro L. Gozalo. Nonparametric bootstrap analysis with applications to demographic effects in demand functions. Journal of Econometrics, 81(2):357–393, December 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697865712.

Ghysels:1993:ESA

[GP93] Eric Ghysels and Pierre Perron. The effect of seasonal adjustment filters on tests for a unit root. Journal of Econometrics, 55(1–2):57–98, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390004O.

Ghysels:1996:ELF

[GP96] Eric Ghysels and Pierre Perron. The effect of linear filters on dynamic time series with structural change. Journal of Econometrics, 70(1):69–97, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016844.

Griliches:1995:FPI

[GR95] Zvi Griliches and Haim Regev. Firm productivity in Israeli industry 1979–1988. Journal of Econometrics, 65(1):175–203, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401601U.

Granger:1992:FOE

[Gra92] Clive W. J. Granger. Fellow’s opinion: Evaluating economic theory. Journal of Econometrics, 51(1–2):3–5, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290026N.

Gonzalez-Rivera:1999:ECM

[GRD99] Gloria González-Rivera and Feike C. Drost. Efficiency comparisons of maximum-likelihood-based estimators in GARCH models. Journal of Econometrics, 93(1):93–111, November 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000056.

Greene:1990:MRT

[Gre90a] William Greene. Multiple roots of the Tobit log-likelihood. Journal of Econometrics, 46(3):365–380, December 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090015L.

Greene:1990:GDS

[Gre90b] William H. Greene. A Gamma-distributed stochastic frontier model. Journal of Econometrics, 46(1–2):141–163, October/ November 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090052U.

Gritz:1993:ITF

[Gri93] R. Mark Gritz. The impact of training on the frequency and duration of employment. Journal of Econometrics, 57(1–3):21– 51, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390057C.

Grillenzoni:1996:TCR

[Gri96] Carlo Grillenzoni. Testing for causality in real time. Journal of Econometrics, 73(2):355–376, August 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407695017291.

Gurmu:1999:SEC

[GRS99] Shiferaw Gurmu, Paul Rilstone, and Steven Stern. Semiparametric estimation of count regression models 1. Journal of Econometrics, 88(1):123–150, January 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000268.

Gallant:1991:ANF

[GS91] A. Ronald Gallant and Geraldo Souza. On the asymptotic normality of Fourier flexible form estimates. Journal of Econometrics, 50(3):329–353, December 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407691900248.

Gong:1992:FSE

[GS92] Byeong-Ho Gong and Robin C. Sickles. Finite sample evidence on the performance of stochastic frontiers and data envelopment analysis using panel data. Journal of Econometrics, 51(1–2):259–284, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290038S.

Granger:1995:SST

[GS95] C. W. J. Granger and Pierre L. Siklos. Systematic sampling, temporal aggregation, seasonal adjustment, and cointegration theory and evidence. Journal of Econometrics, 66(1–2):357–369, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694016227.

Gertler:1997:PHI

[GS97a] Paul Gertler and Roland Sturm. Private health insurance and public expenditures in Jamaica. Journal of Econometrics, 77(1):

237–257, March 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696018143.

Granger:1997:ISU

[GS97b] Clive W. J. Granger and Norman R. Swanson. An introduction to stochastic unit-root processes. Journal of Econometrics, 80(1):35– 62, September 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696000164.

Gurmu:1992:OTT

[GT92] Shiferaw Gurmu and Pravin K. Trivedi. Overdispersion tests for truncated Poisson regression models. Journal of Econometrics, 54(1–3):347–370, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692901136.

Gallant:1999:REM

[GT99] A. Ronald Gallant and George Tauchen. The relative efficiency of method of moments estimators 1. Journal of Econometrics, 92(1):149–172, September 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407698000888.

Granger:1990:REB

[GU90] Clive W. J. Granger and Harald F. Uhlig. Reasonable extremebounds analysis. Journal of Econometrics, 44(1–2):159–170, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900777.

Gourieroux:1997:CDM

[GV97] C. Gourieroux and M. Visser. A count data model with unobserved heterogeneity. Journal of Econometrics, 79(2):247–268, August 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000225.

Gregory:1993:ESE

[GW93] Allan W. Gregory and Tony Wirjanto. The effect of sampling error on the time series behavior of consumption data.

Journal of Econometrics, 55(1–2):267–273, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390015W.

Galbraith:1995:TEC

[GZW95] John W. Galbraith and Victoria Zinde-Walsh. Transforming the error-components model for estimation with general ARMA disturbances. Journal of Econometrics, 66(1–2):349–355, March/ April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694016216.

Galbraith:1999:DAD

[GZW99] John W. Galbraith and Victoria Zinde-Walsh. On the distributions of Augmented Dickey–Fuller statistics in processes with moving average components. Journal of Econometrics, 93(1):25– 47, November 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000979.

Hsiao:1993:GFP

[HAD93] Cheng Hsiao, Trent W. Appelbe, and Christopher R. Dineen. A general framework for panel data models with an application to Canadian customer-dialed long distance telephone service. Journal of Econometrics, 59(1–2):63–86, September 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900398.

Hadri:1996:NSD

[Had96] Kaddour Hadri. A note on sargan densities. Journal of Econometrics, 71(1–2):285–290, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017050.

Hahn:1997:EEP

[Hah97] Jinyong Hahn. Efficient estimation of panel data models with sequential moment restrictions. Journal of Econometrics, 79(1): 1–21, July 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000122.

Hahn:1999:HII

[Hah99] Jinyong Hahn. How informative is the initial condition in the dynamic panel model with fixed effects? Journal of Econometrics, 93(2):309–326, December 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000135.

Hall:1992:TUR

[Hal92] Alastair Hall. Testing for a unit root in time series using instrumental variable estimators with pretest data based model selection. Journal of Econometrics, 54(1–3):223–250, October/ December 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692901073.

Hall:1993:ISP

[Hal93] Alastair Hall. Induced seasonality and production-smoothing models of inventory behavior. Journal of Econometrics, 55(1– 2):169–172, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390009T.

Haldrup:1994:ASE

[Hal94a] Niels Haldrup. The asymptotics of single-equation cointegration regressions with i(1) and i(2) variables. Journal of Econometrics, 63(1):153–181, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693015643.

Hall:1994:TUR

[Hal94b] Alastair Hall. Testing for a unit root in time series using instrumental variable estimators with pretest data based model selection (vol. 54 (1992) pp. 223-250). Journal of Econometrics, 60(1–2):321, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900507.

Haldrup:1996:MID

[Hal96] Niels Haldrup. Mirror image distributions and the Dickey–Fuller regression with a maintained trend. Journal of Econometrics, 72(1–2):301–312, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694017247.

Hamilton:1990:ATS

[Ham90] James D. Hamilton. Analysis of time series subject to changes in regime. Journal of Econometrics, 45(1–2):39–70, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690900939.

Hamilton:1996:STM

[Ham96] James D. Hamilton. Specification testing in Markov-switching time-series models. Journal of Econometrics, 70(1):127–157, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407669416869.

Hansen:1992:EET

[Han92a] Bruce E. Hansen. Efficient estimation and testing of cointegrating vectors in the presence of deterministic trends. Journal of Econometrics, 53(1–3):87–121, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900812.

Hansen:1992:HC

[Han92b] Bruce E. Hansen. Heteroskedastic cointegration. Journal of Econometrics, 54(1–3):139–158, October/December 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290103X.

Hansen:1999:TEN

[Han99] Bruce E. Hansen. Threshold effects in non-dynamic panels: Estimation, testing, and inference. Journal of Econometrics, 93(2):345–368, December 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407699000251.

Hausman:1998:MDV

[HASM98] J. A. Hausman, Jason Abrevaya, and F. M. Scott-Morton. Misclassification of the dependent variable in a discrete-response setting. Journal of Econometrics, 87(2):239–269, December 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000153.

Haug:1996:TCM

[Hau96] Alfred A. Haug. Tests for cointegration a Monte Carlo comparison. Journal of Econometrics, 71(1–2):89–115, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694016968.

Hylleberg:1990:SIC

[HEGY90] S. Hylleberg, R. F. Engle, C. W. J. Granger, and B. S. Yoo. Seasonal integration and cointegration. Journal of Econometrics, 44(1–2):215–238, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090080D.

Hillier:1991:MDP

[Hil91] Grant H. Hillier. On multiple diagnostic procedures for the linear model. Journal of Econometrics, 47(1):47–66, January 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190077Q.

Hinkley:1997:DPH

[Hin97] D. V. Hinkley. Discussion of paper by H. Li and G. S. Maddala. Journal of Econometrics, 80(2):319–323, October 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000456. See [LM97].

Hillier:1991:EIY

[HK91] Grant H. Hillier and Maxwell L. King. Editors’ introduction: 40 years of diagnostic testing. Journal of Econometrics, 47(1): 1–4, January 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190075O.

Hardle:1995:NDM

[HK95] Wolfgang Härdle and Alan Kirman. Nonclassical demand: A model-free examination of price–quantity relations in the Marseille fish market. Journal of Econometrics, 67(1):227–257, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401634C.

Harris:1998:MHP

[HK98] Katherine M. Harris and Michael P. Keane. A model of health plan choice:: Inferring preferences and perceptions from a combination of revealed preference and attitudinal data. Journal of Econometrics, 89(1–2):131–157, November 26, 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769800058X.

Hsiao:1990:SPI

[HKT90] Cheng Hsiao, Changseob Kim, and Grant Taylor. A statistical perspective on insurance rate-making. Journal of Econometrics, 44(1–2):5–24, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090070A.

Honore:1997:ETT

[HKU97] Bo E. Honoré, Ekaterini Kyriazidou, and Christopher Udry. Estimation of Type 3 Tobit models using symmetric trimming and pairwise comparisons. Journal of Econometrics, 76(1–2): 107–128, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017844.

Hensher:1998:CSP

[HLS98] David Hensher, Jordan Louviere, and Joffre Swait. Combining sources of preference data. Journal of Econometrics, 89 (1–2):197–221, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440769800061X.

Hoek:1995:CBA

[HLvD95] Henk Hoek, André Lucas, and Herman K. van Dijk. Classical and Bayesian aspects of robust unit root inference. Journal of Econometrics, 69(1):27–59, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401661I.

Hardle:1993:NSA

[HM93] Wolfgang Härdle and Charles F. Manski. Nonparametric and semiparametric approaches to discrete response analysis. Journal of Econometrics, 58(1–2):1–2, July 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390110Q.

Hisamatsu:1994:DDW

[HM94] Hiroyuki Hisamatsu and Koichi Maekawa. The distribution of the Durbin–Watson statistic in integrated and near-integrated models. Journal of Econometrics, 61(2):367–382, April 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900906.

Hall:1995:ERB

[HM95] Bronwyn H. Hall and Jacques Mairesse. Exploring the relationship between R&D and productivity in French manufacturing firms. Journal of Econometrics, 65(1):263–293, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401604X.

Horowitz:1998:COR

[HM98] Joel L. Horowitz and Charles F. Manski. Censoring of outcomes and regressors due to survey nonresponse: Identification and estimation using weights and imputations. Journal of Econometrics, 84(1):37–58, May 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000778.

Hajivassiliou:1996:SMN

[HMR96] Vassilis Hajivassiliou, Daniel McFadden, and Paul Ruud. Simulation of multivariate normal rectangle probabilities and their derivatives theoretical and computational results. Journal of Econometrics, 72(1–2):85–134, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017166.

Hirschberg:1991:CAM

[HMS91] Joseph G. Hirschberg, Esfandiar Maasoumi, and Daniel J. Slottje. Cluster analysis for measuring welfare and quality of life across countries. Journal of Econometrics, 50(1–2):131–150, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190093S.

Hausman:1991:IEP

[HNIP91] Jerry A. Hausman, Whitney K. Newey, Hidehiko Ichimura, and James L. Powell. Identification and estimation of polynomial errors-in-variables models. Journal of Econometrics, 50(3):273– 295, December 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407691900226.

Hausman:1995:NEV

[HNP95] J. A. Hausman, W. K. Newey, and J. L. Powell. Nonlinear errors in variables estimation of some Engel curves. Journal of Econometrics, 65(1):205–233, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401602V.

Hodoshima:1992:FSP

[Hod92] Jiro Hodoshima. Finite-sample properties of single-equation estimators under structural change. Journal of Econometrics, 53 (1–3):189–209, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407692900856.

Hodgson:1998:AEC

[Hod98] Douglas J. Hodgson. Adaptive estimation of cointegrating regressions with ARMA errors. Journal of Econometrics, 85(2):231– 267, August 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697001012.

Holm:1993:MEL

[Hol93] Juhani Holm. Maximum entropy Lorenz curves. Journal of Econometrics, 59(3):377–389, October 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390031Y.

Holly:1995:RLF

[Hol95] Alberto Holly. A random linear functional approach to efficiency bounds. Journal of Econometrics, 65(1):235–261, January 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401603W.

Honore:1993:OCT

[Hon93] Bo E. Honoré. Orthogonality conditions for Tobit models with fixed effects and lagged dependent variables. Journal of Econometrics, 59(1–2):35–61, September 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900387.

Horowitz:1993:SEW

[Hor93] Joel L. Horowitz. Semiparametric estimation of a work-trip mode choice model. Journal of Econometrics, 58(1–2):49–70, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390113J.

Horowitz:1994:BBC

[Hor94] Joel L. Horowitz. Bootstrap-based critical values for the information matrix test. Journal of Econometrics, 61(2):395–411, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900922.

Hosking:1996:ADS

[Hos96a] Jonathan R. M. Hosking. Asymptotic distributions of the sample mean, autocovariances, and autocorrelations of long-memory time series. Journal of Econometrics, 73(1):261–284, July 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017402.

Hosoya:1996:QLA

[Hos96b] Yuzo Hosoya. The quasi-likelihood approach to statistical inference on multiple time-series with long-range dependence.

Journal of Econometrics, 73(1):217–236, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017380.

Honore:1994:PDE

[HP94] Bo E. Honoré and James L. Powell. Pairwise difference estimators of censored and truncated regression models. Journal of Econometrics, 64(1–2):241–278, September/October 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900655.

Hsiao:1993:EI

[HR93] Cheng Hsiao and Paul Ruud. Editors’ introduction. Journal of Econometrics, 56(1–2):1–3, March 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390098P.

Hidalgo:1996:TSC

[HR96] Javier Hidalgo and Peter M. Robinson. Testing for structural change in a long-memory environment. Journal of Econometrics, 70(1):159–174, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016879.

Harvey:1992:UCT

[HRS92] Andrew Harvey, Esther Ruiz, and Enrique Sentana. Unobserved component time series models with arch disturbances. Journal of Econometrics, 52(1–2):129–157, April/May 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900683.

Heineke:1990:AIC

[HS90] J. M. Heineke and H. M. Shefrin. Aggregation and identification in consumer demand systems. Journal of Econometrics, 44(3): 377–390, June 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900652.

Hansen:1993:SAE

[HS93] Lars Peter Hansen and Thomas J. Sargent. Seasonality and approximation errors in rational expectations models. Journal of Econometrics, 55(1–2):21–55, January/February 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390003N.

Hamilton:1994:ACH

[HS94a] James D. Hamilton and Raul Susmel. Autoregressive conditional heteroskedasticity and changes in regime. Journal of Econometrics, 64(1–2):307–333, September/October 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900671.

Hirschberg:1994:EBA

[HS94b] Joseph G. Hirschberg and Daniel J. Slottje. An empirical Bayes approach to analyzing earnings functions for various occupations and industries. Journal of Econometrics, 61(1):65–79, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900779.

Hughes:1994:MCS

[HS94c] Gordon A. Hughes and N. E. Savin. Is the minimum chisquare estimator the winner in logit regression? Journal of Econometrics, 61(2):345–366, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900892.

Hwang:1996:AMD

[HS96] Jaeyoun Hwang and Peter Schmidt. Alternative methods of detrending and the power of unit root tests. Journal of Econometrics, 71(1–2):227–248, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017026.

Haldrup:1998:RCS

[HS98a] Niels Haldrup and Mark Salmon. Representations of I(2) cointegrated systems using the Smith–McMillan form. Journal of Econometrics, 84(2):303–325, June 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000882.

Harvey:1998:TSC

[HS98b] Andrew Harvey and Mariane Streibel. Testing for a slowly changing level with special reference to stochastic volatility. Journal of Econometrics, 87(1):167–189, November 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000128.

Hsiao:1998:MSR

[HS98c] Cheng Hsiao and Bao-Hong Sun. Modeling survey response bias — with an analysis of the demand for an advanced electronic device. Journal of Econometrics, 89(1–2):15–39, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000530.

Hayes:1990:SDI

[HSPHS90] Kathy Hayes, D. J. Slottje, Susan Porter-Hudak, and Gerald Scully. Is the size distribution of income a random walk? Journal of Econometrics, 43(1–2):213–226, January/February 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090117C.

Hansen:1998:SMI

[HST98] Lars Peter Hansen, José Alexandre Scheinkman, and Nizar Touzi. Spectral methods for identifying scalar diffusions. Journal of Econometrics, 86(1):1–32, September 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697001073.

Hardle:1993:HSA

[HT93] Wolfgang Härdle and A. B. Tsybakov. How sensitive are average derivatives? Journal of Econometrics, 58(1–2):31–48, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390112I.

Hardle:1997:LPE

[HT97] W. Härdle and A. Tsybakov. Local polynomial estimators of the volatility function in nonparametric autoregression. Journal of Econometrics, 81(1):223–242, November 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000444.

Harris:1999:IUR

[HT99a] Richard D. F. Harris and Elias Tzavalis. Inference for unit roots in dynamic panels where the time dimension is fixed. Journal of Econometrics, 91(2):201–226, August 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000761.

He:1999:PMF

[HT99b] Changli He and Timo Teräsvirta. Properties of moments of a family of GARCH processes. Journal of Econometrics, 92(1):173– 192, September 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769800089X.

Hwang:1994:CSC

[HU94] J. T. Gene Hwang and Aman Ullah. Confidence sets centered at James–Stein estimators: A surprise concerning the unknownvariance case. Journal of Econometrics, 60(1–2):145–156, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900418.

Hussey:1992:NEA

[Hus92] Robert Hussey. Nonparametric evidence on asymmetry in business cycles using aggregate employment time series. Journal of Econometrics, 51(1–2):217–231, January/February 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290036Q.

Hackl:1996:DIT

[HW96] Peter Hackl and Anders H. Westlund. Demand for international telecommunication time-varying price elasticity. Journal of Econometrics, 70(1):243–260, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016917.

Hauver:1992:MMC

[HY92] James H. Hauver and Jet Yee. Morrison’s measure of capacity utilization: A critique. Journal of Econometrics, 52(3):403– 406, June 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290020R.

Hylleberg:1995:TSU

[Hyl95] Svend Hylleberg. Tests for seasonal unit roots general to specific or specific to general? Journal of Econometrics, 69(1):5– 25, September 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401660R.

Im:1999:EEP

[IASW99] Kyung So Im, Seung C. Ahn, Peter Schmidt, and Jeffrey M. Wooldridge. Efficient estimation of panel data models with strictly exogenous explanatory variables. Journal of Econometrics, 93(1):177–201, November 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000081.

Ichimura:1993:SLS

[Ich93] Hidehiko Ichimura. Semiparametric least squares (SLS) and weighted SLS estimation of single-index models. Journal of Econometrics, 58(1–2):71–120, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390114K.

Imbens:1996:EES

[IL96] Guido W. Imbens and Tony Lancaster. Efficient estimation and stratified sampling. Journal of Econometrics, 74(2):289–318, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017569.

Inder:1993:ELR

[Ind93] Brett Inder. Estimating long-run relationships in economics: A comparison of different approaches. Journal of Econometrics, 57(1–3):53–68, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390058D.

Inoue:1999:TCR

[Ino99] Atsushi Inoue. Tests of cointegrating rank with a trendbreak. Journal of Econometrics, 90(2):215–237, June 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000426.

Ichimura:1998:MLE

[IT98] Hidehiko Ichimura and T. Scott Thompson. Maximum likelihood estimation of a binary choice model with random coefficients of unknown distribution. Journal of Econometrics, 86(2):269– 295, October 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697001176.

Iwata:1992:HPD

[Iwa92a] Shigeru Iwata. Highest predictive density estimator in regression models. Journal of Econometrics, 53(1–3):271–295, July/ September 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692900889.

Iwata:1992:IVE

[Iwa92b] Shigeru Iwata. Instrumental variables estimation in errorsin-variables models when instruments are correlated with errors. Journal of Econometrics, 53(1–3):297–322, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290089A.

Iwata:1993:NMR

[Iwa93] Shigeru Iwata. A note on multiple roots of the Tobit log likelihood. Journal of Econometrics, 56(3):441–445, April 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390129S.

Iwata:1994:ETW

[Iwa94] Shigeru Iwata. On estimation and testing when explanatory variables are partly endogenous. Journal of Econometrics, 61(2):413– 428, April 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900930.

Iwata:1996:BPM

[Iwa96] Shigeru Iwata. Bounding posterior means by model criticism. Journal of Econometrics, 75(2):239–261, December 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017593.

Jenkins:1994:EDM

[Jen94] Stephen P. Jenkins. Earnings discrimination measurement: A distributional approach. Journal of Econometrics, 61(1):81– 102, March 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900787.

Judge:1990:AEB

[JHB90] G. G. Judge, R. Carter Hill, and M. E. Bock. An adaptive empirical Bayes estimator of the multivariate normal mean under quadratic loss. Journal of Econometrics, 44(1–2):189–213, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900799.

Johansen:1992:TSH

[JJ92] Søren Johansen and Katarina Juselius. Testing structural hypotheses in a multivariate cointegration analysis of the PPP and the UIP for UK. Journal of Econometrics, 53(1–3):211–244, July/ September 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692900867.

Johansen:1994:ILR

[JJ94] Søren Johansen and Katarina Juselius. Identification of the longrun and the short-run structure an application to the ISLM model. Journal of Econometrics, 63(1):7–36, July 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693015595.

Lee:1992:MRO

[jL92] Myoung jae Lee. Median regression for ordered discrete response. Journal of Econometrics, 51(1–2):59–77, January/ February 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290029Q.

Lee:1993:QMR

[jL93] Myoung jae Lee. Quadratic mode regression. Journal of Econometrics, 57(1–3):1–19, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390056B.

Jayet:1991:ASD

[JM91] H. Jayet and A. Moreau. Analysis of survival data: Estimation and specification tests using asymptotic least squares. Journal of Econometrics, 48(1–2):263–285, April/May 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190041B.

Johansen:1992:CPS

[Joh92] Søren Johansen. Cointegration in partial systems and the efficiency of single-equation analysis. Journal of Econometrics, 52(3): 389–402, June 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769290019N.

Johansen:1995:IRL

[Joh95] Søren Johansen. Identifying restrictions of linear equations with applications to simultaneous equations and cointegration. Journal of Econometrics, 69(1):111–132, September 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401664L.

Jorgenson:1990:ISL

[JS90] Dale W. Jorgenson and Daniel T. Slesnick. Inequality and the standard of living. Journal of Econometrics, 43(1–2): 103–120, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090110F.

Johansen:1999:LAS

[JS99a] Søren Johansen and Ernst Schaumburg. Likelihood analysis of seasonal cointegration. Journal of Econometrics, 88(2):301–339, February 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000359.

Johansen:1999:TER

[JS99b] Søren Johansen and Anders Rygh Swensen. Testing exact rational expectations in cointegrated vector autoregressive models. Journal of Econometrics, 93(1):73–91, November 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000044.

Jaggia:1994:JSS

[JT94] Sanjiv Jaggia and Pravin K. Trivedi. Joint and separate score tests for state dependence and unobserved heterogeneity. Journal of Econometrics, 60(1–2):273–291, January/February 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900477.

Juselius:1995:DPP

[Jus95] Katarina Juselius. Do purchasing power parity and uncovered interest rate parity hold in the long run? An example of likelihood inference in a multivariate time-series model. Journal of Econometrics, 69(1):211–240, September 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401669Q.

Kai:1998:BIS

[Kai98] Li Kai. Bayesian inference in a simultaneous equation model with limited dependent variables. Journal of Econometrics, 85(2):387– 400, August 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697001061.

Kao:1999:SRR

[Kao99] Chihwa Kao. Spurious regression and residual-based tests for cointegration in panel data. Journal of Econometrics, 90(1): 1–44, May 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000232.

Karanasos:1999:SMA

[Kar99] Menelaos Karanasos. The second moment and the autocovariance function of the squared errors of the GARCH model. Journal of Econometrics, 90(1):63–76, May 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000323.

Karlsson:2017:CBR

[Kar17] Sune Karlsson. Corrigendum to “Bayesian reduced rank regression in econometrics” [J. Econometrics 75 (1996) 121–146]. Journal of Econometrics, 201(1):170–171, November 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407615002754.

Keuzenkamp:1995:RFH

[KB95] Hugo A. Keuzenkamp and Anton P. Barten. Rejection without falsification on the history of testing the homogeneity condition in the theory of consumer demand. Journal of Econometrics, 67(1): 103–127, May 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401629E.

Kazimi:1999:BCB

[KB99] Camilla Kazimi and David Brownstone. Bootstrap confidence bands for shrinkage estimators. Journal of Econometrics, 90(1): 99–127, May 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000372.

Kim:1995:SEN

[KCH95] Minbo Kim and R. Carter-Hill. Shrinkage estimation in nonlinear regression the Box–Cox transformation. Journal of Econometrics, 66(1–2):1–33, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401606Z.

Kadane:1996:PUR

[KCW96] Joseph B. Kadane, Ngai Hang Chan, and Lara J. Wolfson. Priors for unit root models. Journal of Econometrics, 75(1):99–111, November 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017712.

Kiviet:1997:ETS

[KD97] Jan F. Kiviet and Jean-Marie Dufour. Exact tests in single equation autoregressive distributed lag models. Journal of Econometrics, 80(2):325–353, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000481.

Kim:1995:EMP

[KDM95] Jinbang Kim, Neil De Marchi, and Mary S. Morgan. Empirical model particularities and belief in the natural rate hypothesis. Journal of Econometrics, 67(1):81–102, May 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401628D.

Kemp:1991:WTG

[Kem91] Gordon C. R. Kemp. On Wald tests for globally and locally quadratic restrictions. Journal of Econometrics, 50(3):257–272, December 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407691900215.

Kemp:1992:PEG

[Kem92] Gordon C. R. Kemp. The potential for efficiency gains in estimation from the use of additional moment restrictions. Journal of Econometrics, 53(1–3):387–399, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900937.

Kim:1994:DLM

[Kim94] Chang-Jin Kim. Dynamic linear models with Markov-switching. Journal of Econometrics, 60(1–2):1–22, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0304407694900361.

Kinal:1991:NEM

[Kin91] Terrence Kinal. A note on the existence of moments of k-class estimators when k is negative. Journal of Econometrics, 48(3): 409–410, June 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190072L.

Kiviet:1995:BIE

[Kiv95] Jan F. Kiviet. On bias, inconsistency, and efficiency of various estimators in dynamic panel data models. Journal of Econometrics, 68 (1):53–78, July 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401643E.

Klein:1993:STB

[Kle93] Roger W. Klein. Specification tests for binary choice models based on index quantiles. Journal of Econometrics, 59(3):343– 375, October 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900309.

Klein:1998:EIS

[Kle98] Lawrence R. Klein. Editor’s introduction studies in econometrics in honor of Carl F. Christ. Journal of Econometrics, 83(1–2):1–7, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000626.

Koop:1997:BAL

[KLOS97] Gary Koop, Eduardo Ley, Jacek Osiewalski, and Mark F. J. Steel. Bayesian analysis of long memory and persistence using ARFIMA models. Journal of Econometrics, 76(1–2):149–169, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017879.

Kopp:1990:MBE

[KM90] Raymond J. Kopp and John Mullahy. Moment-based estimation and testing of stochastic frontier models. Journal of Econometrics, 46(1–2):165–183, October/November 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090053V.

Keuzenkamp:1995:EIS

[KM95a] Hugo A. Keuzenkamp and Jan R. Magnus. Editors’ introduction: The significance of testing in econometrics. Journal of Econometrics, 67(1):1–3, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016238.

Keuzenkamp:1995:TSE

[KM95b] Hugo A. Keuzenkamp and Jan R. Magnus. On tests and significance in econometrics. Journal of Econometrics, 67(1):5–24, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694016249.

Kramer:1997:AHC

[KM97] Walter Krämer and Sonja Michels. Autocorrelation- and heteroskedasticity-consistent t-values with trending data. Journal of Econometrics, 76(1–2):141–147, January/February 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017860.

Koenker:1999:GIW

[KM99] Roger Koenker and José A. F. Machado. GMM inference when the number of moment conditions is large. Journal of Econometrics, 93(2):327–344, December 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000147.

Min:1993:BNB

[kMZ93] Chung ki Min and Arnold Zellner. Bayesian and non-Bayesian methods for combining models and forecasts with applications to forecasting international growth rates. Journal of Econometrics, 56(1–2):89–118, March 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390102B.

Kato:1996:BMN

[KNI96] Hiroko Kato, Sadao Naniwa, and Makio Ishiguro. A Bayesian multivariate nonstationary time series model for estimating mutual relationships among variables. Journal of Econometrics, 75(1):147– 161, November 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017747.

Krol:1990:ISD

[KO90] Robert Krol and Lee E. Ohanian. The impact of stochastic and deterministic trends on money-output causality: A multicountry investigation. Journal of Econometrics, 45(3):291–308, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090001A.

Krol:1993:ISD

[KO93] Robert Krol and Lee E. Ohanian. The impact of stochastic and deterministic trends on money-output causality: A multi-country investigation (vol. 45, no. 3 (1990) pp. 291-308). Journal of Econometrics, 58(3):405, August 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900538.

Koop:1991:CTP

[Koo91] Gary Koop. Cointegration tests in present value relationships: A Bayesian look at the bivariate properties of stock prices and dividends. Journal of Econometrics, 49(1–2):105–139, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407691900112.

Koop:1996:PUI

[Koo96] Gary Koop. Parameter uncertainty and impulse response analysis. Journal of Econometrics, 72(1–2):135–149, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694017174.

Koop:1995:BLR

[KOS95] Gary Koop, Jacek Osiewalski, and Mark F. J. Steel. Bayesian long-run prediction in time series models. Journal of Econometrics, 69(1):61–80, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401662J.

Koop:1997:BEA

[KOS97] Gary Koop, Jacek Osiewalski, and Mark F. J. Steel. Bayesian efficiency analysis through individual effects: Hospital cost frontiers. Journal of Econometrics, 76(1–2):77–105, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017836.

Koop:1993:BAL

[KP93a] Gary Koop and Dale J. Poirier. Bayesian analysis of logit models using natural conjugate priors. Journal of Econometrics, 56(3): 323–340, April 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390124N.

Kurosawa:1993:MAT

[KP93b] Masako Kurosawa and Stephen Pudney. A method for the analysis of the timing and magnitude of events in a continuous-time panel: The effects of British incomes policy, 1950-1973. Journal of Econometrics, 59(1–2):161–185, September 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900446.

Kiviet:1994:BAR

[KP94] Jan F. Kiviet and Garry D. A. Phillips. Bias assessment and reduction in linear error-correction models. Journal of Econometrics, 63(1):215–243, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693015665.

Koenker:1996:IPA

[KP96] Roger Koenker and Beum J. Park. An interior point algorithm for nonlinear quantile regression. Journal of Econometrics, 71 (1–2):265–283, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407696845076.

Kitamura:1997:FMI

[KP97a] Yuichi Kitamura and Peter C. B. Phillips. Fully modified IV, GIVE and GMM estimation with possibly non-stationary regressors and instruments. Journal of Econometrics, 80(1):85–123, September 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000043.

Koop:1997:LAAa

[KP97b] Gary Koop and Dale J. Poirier. Learning about the acrossregime correlation in switching regression models. Journal of Econometrics, 78(1):217–227, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800115.

Koop:1997:LAAb

[KP97c] Gary Koop and Dale J. Poirier. Learning about the acrossregime correlation in switching regression models. Journal of Econometrics, 78(2):217–227, June 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696000097.

Koop:1999:BFN

[KP99] Gary Koop and Simon M. Potter. Bayes factors and nonlinearity: Evidence from economic time series 1 dagger. Journal of Econometrics, 88(2):251–281, February 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000311.

Koop:1996:IRA

[KPP96] Gary Koop, M. Hashem Pesaran, and Simon M. Potter. Impulse response analysis in nonlinear multivariate models. Journal of Econometrics, 74(1):119–147, September 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017534.

Kiviet:1995:BOG

[KPS95] Jan F. Kiviet, Garry D. A. Phillips, and Bernhard Schipp. The bias of OLS, GLS, and ZEF estimators in dynamic seemingly unrelated regression models. Journal of Econometrics, 69(1):241– 266, September 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401670U.

Kwiatkowski:1992:TNH

[KPSS92] Denis Kwiatkowski, Peter C. B. Phillips, Peter Schmidt, and Yongcheol Shin. Testing the null hypothesis of stationarity against the alternative of a unit root: How sure are we that economic time series have a unit root? Journal of Econometrics, 54(1–3):159–178, October/December 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290104Y.

Koning:1994:CNL

[KR94] Ruud H. Koning and Geert Ridder. On the compatibility of nested logit models with utility maximization: A comment. Journal of Econometrics, 63(2):389–396, August 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693015573.

Krishnakumar:1997:RESa

[KR97a] J. Krishnakumar and E. Ronchetti. Robust estimators for simultaneous equations models. Journal of Econometrics, 78(1):295– 314, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697800140.

Krishnakumar:1997:RESb

[KR97b] J. Krishnakumar and E. Ronchetti. Robust estimators for simultaneous equations models. Journal of Econometrics, 78(2): 295–314, June 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696000140.

Krane:1993:ISP

[Kra93] Spencer D. Krane. Induced seasonality and productionsmoothing models of inventory behavior. Journal of Econometrics, 55(1–2):135–168, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390008S.

Kobayashi:1990:MCU

[KS90] Masahito Kobayashi and Shinichi Sakata. Mallows’ C_{p} criterion and unbiasedness of model selection. Journal of Econometrics, 45(3):385–395, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090006F.

Kennedy:1991:FTF

[KS91] Peter Kennedy and Daniel Simons. Fighting the teflon factor: Comparing classical and Bayesian estimators for autocorrelated errors. Journal of Econometrics, 48(1–2):15–27, April/ May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190030H.

Kim:1993:URT

[KS93] Kiwhan Kim and Peter Schmidt. Unit root tests with conditional heteroskedasticity. Journal of Econometrics, 59(3):287– 300, October 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900273.

Klein:1997:ENP

[KS97] Roger Klein and Robert Sherman. Estimating new product demand from biased survey data. Journal of Econometrics, 76(1–2):53–76, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017828.

Kokoszka:1996:IVS

[KT96] Piotr S. Kokoszka and Murad S. Taqqu. Infinite variance stable moving averages with long memory. Journal of Econometrics, 73 (1):79–99, July 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017348.

Kuan:1998:TCM

[Kua98] Chung-Ming Kuan. Tests for changes in models with a polynomial trend. Journal of Econometrics, 84(1):75–91, May 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000808.

Kumbhakar:1990:PFP

[Kum90] Subal C. Kumbhakar. Production frontiers, panel data, and time-varying technical inefficiency. Journal of Econometrics, 46(1–2):201–211, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090055X.

Kumbhakar:1996:FLS

[Kum96] Subal C. Kumbhakar. A farm-level study of labor use and efficiency wages in Indian agriculture. Journal of Econometrics, 72(1–2):177–195, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694017190.

Kumbhakar:1997:MAI

[Kum97] Subal C. Kumbhakar. Modeling allocative inefficiency in a translog cost function and cost share equations: An exact relationship. Journal of Econometrics, 76(1–2):351–356, January/ February 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017968.

Kuo:1998:TPP

[Kuo98] Biing-Shen Kuo. Test for partial parameter instability in regressions with i(1) processes. Journal of Econometrics, 86(2):337– 368, October 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000050.

Kadiyali:1998:PLE

[KVC98] Vrinda Kadiyali, Naufel Vilcassim, and Pradeep Chintagunta. Product line extensions and competitive market interactions: An empirical analysis. Journal of Econometrics, 89(1–2):339–363, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000669.

Kiviet:1994:SDE

[KvD94a] Jan F. Kiviet and Herman K. van Dijk. Structure and dynamics in econometrics. Journal of Econometrics, 63(1):1–5, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693015584.

Kleibergen:1994:DCT

[KvD94b] Frank Kleibergen and Herman K. van Dijk. Direct cointegration testing in error correction models. Journal of Econometrics, 63(1): 61–103, July 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769301561Y.

King:1991:SDA

[KW91] Maxwell L. King and Ping X. Wu. Small-disturbance asymptotics and the Durbin–Watson and related tests in the dynamic regression model. Journal of Econometrics, 47(1):145–152, January 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190081N.

Kwan:1999:ABA

[Kwa99] Yum K. Kwan. Asymptotic Bayesian analysis based on a limited information estimator. Journal of Econometrics, 88(1):99– 121, January 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000244.

Kakwani:1997:SIH

[KWvD97] Nanak Kakwani, Adam Wagstaff, and Eddy van Doorslaer. Socioeconomic inequalities in health: Measurement, computation, and statistical inference. Journal of Econometrics, 77(1):87– 103, March 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696018076.

Kyriazidou:1998:TSC

[Kyr98] Ekaterini Kyriazidou. Testing for serial correlation in multivariate regression models. Journal of Econometrics, 86(2):193– 220, October 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697001140.

Kramer:1990:FSP

[KZ90] Walter Krämer and Helmut Zeisel. Finite sample power of linear regression autocorrelation tests. Journal of Econometrics, 43(3):

363–372, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090125D.

Labeaga:1999:DHR

[Lab99] José M. Labeaga. A double-hurdle rational addiction model with heterogeneity: Estimating the demand for tobacco. Journal of Econometrics, 93(1):49–72, November 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000032.

Laitila:1993:PRM

[Lai93] Thomas Laitila. A pseudo-R^{2} measure for limited and qualitative dependent variable models. Journal of Econometrics, 56(3): 341–355, April 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390125O.

Lancaster:1997:BWP

[Lan97] Tony Lancaster. Bayes WESML posterior inference from choicebased samples. Journal of Econometrics, 79(2):291–303, August 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696000243.

Lau:1999:ICT

[Lau99] Sau-Him Paul Lau. I (0) in, integration and cointegration out:: Time series properties of endogenous growth models. Journal of Econometrics, 93(1):1–24, November 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000967.

Lawrence:1990:ACM

[Law90] Denis Lawrence. An adjustment-costs model of export supply and import demand. Journal of Econometrics, 46(3):381–398, December 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090016M.

Lutkepohl:1997:MWTa

[LB97a] Helmut Lütkepohl and Maike M. Burda. Modified Wald tests under nonregular conditions. Journal of Econometrics, 78(1):315– 332, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697800152.

Lutkepohl:1997:MWTb

[LB97b] Helmut Lütkepohl and Maike M. Burda. Modified Wald tests under nonregular conditions. Journal of Econometrics, 78(2):315– 332, June 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696000152.

Lutkepohl:1997:ACV

[LC97] Helmut Lütkepohl and Holger Claessen. Analysis of cointegrated VARMA processes. Journal of Econometrics, 80(2):223–239, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000353.

Lee:1992:MLI

[Lee92] Hahn Shik Lee. Maximum likelihood inference on cointegration and seasonal cointegration. Journal of Econometrics, 54(1–3):1–47, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290098C.

Lee:1994:SIV

[Lee94] Lung-Fei Lee. Semiparametric instrumental variable estimation of simultaneous equation sample selection models. Journal of Econometrics, 63(2):341–388, August 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693015713.

Lee:1997:SML

[Lee97a] Lung-Fei Lee. Simulated maximum likelihood estimation of dynamic discrete choice statistical models some Monte Carlo results. Journal of Econometrics, 82(1):1–35, ???? 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000146.

Lee:1997:SEDb

[Lee97b] Lung-Fei Lee. Simulation estimation of dynamic switching regression and dynamic disequilibrium models — some Monte Carlo results. Journal of Econometrics, 78(2):179–184, 186–204, June 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696000073.

Lee:1997:SEDa

[Lee97c] Lung-Fei Lee. Simulation estimation of dynamic switching regression and dynamic disequilibrium models-some Monte Carlo results. Journal of Econometrics, 78(1):179–204, ???? 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800097.

Lee:1997:SLSa

[Lee97d] Lung-Fei Lee. A smooth likelihood simulator for dynamic disequilibrium models. Journal of Econometrics, 78(1):257–294, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697800139.

Lee:1997:SLSb

[Lee97e] Lung-Fei Lee. A smooth likelihood simulator for dynamic disequilibrium models. Journal of Econometrics, 78(2):257–294, June 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696000139.

Lehmann:1990:RRR

[Leh90] Bruce N. Lehmann. Residual risk revisited. Journal of Econometrics, 45(1–2):71–97, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090094A.

Lewbel:1990:IDM

[Lew90] Arthur Lewbel. Income distribution movements and aggregate money illusion. Journal of Econometrics, 43(1–2):35–42, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690901053.

Lewbel:1995:CNH

[Lew95] Arthur Lewbel. Consistent nonparametric hypothesis tests with an application to Slutsky symmetry. Journal of Econometrics, 67(2): 379–401, June 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401637F.

Lutkepohl:1996:SVC

[LH96] Helmut Lütkepohl and Helmut Herwartz. Specification of varying coefficient time series models via generalized flexible least squares. Journal of Econometrics, 70(1):261–290, January 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016925.

Li:1998:TSC

[LH98] Q. Li and C. Hsiao. Testing serial correlation in semiparametric panel data models. Journal of Econometrics, 87(2):207–237, December 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769800013X.

Lee:1991:SET

[LI91] Bong-Soo Lee and Beth Fisher Ingram. Simulation estimation of time-series models. Journal of Econometrics, 47(2–3):197–205, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190098X.

Lancaster:1995:OSF

[LI95] Tony Lancaster and Guido Imbens. Optimal stock/flow panels. Journal of Econometrics, 66(1–2):325–348, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401620F.

Lancaster:1996:CCS

[LI96] Tony Lancaster and Guido Imbens. Case-control studies with contaminated controls. Journal of Econometrics, 71(1–2):145–160, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694016984.

Li:1999:CMS

[Li99] Qi Li. Consistent model specification tests for time series econometric models. Journal of Econometrics, 92(1):101–147, September 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000876.

Lillard:1993:SEH

[Lil93] Lee A. Lillard. Simultaneous equations for hazards: Marriage duration and fertility timing. Journal of Econometrics, 56(1–2): 189–217, March 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390106F.

Lewin:1990:EI

[LL90] Arie Y. Lewin and C. A. Knox Lovell. Editor’s introduction. Journal of Econometrics, 46(1–2):3–5, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090044T.

Lo:1990:EAN

[LM90] Andrew W. Lo and A. Craig MacKinlay. An econometric analysis of nonsynchronous trading. Journal of Econometrics, 45 (1–2):181–211, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090098E.

Li:1997:BCR

[LM97] Hongyi Li and G. S. Maddala. Bootstrapping cointegrating regressions. Journal of Econometrics, 80(2):297–318, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000432. See discussion [CR91].

Leybourne:1998:SRD

[LMN98] Stephen J. Leybourne, Terence C. Mills, and Paul Newbold. Spurious rejections by Dickey–Fuller tests in the presence of a break under the null. Journal of Econometrics, 87(1):191–203, November 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000141.

Lumsdaine:1999:TAP

[LN99] Robin L. Lumsdaine and Serena Ng. Testing for ARCH in the presence of a possibly misspecified conditional mean. Journal of Econometrics, 93(2):257–279, December 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000111.

Lasserre:1991:MPS

[LO91] Pierre Lasserre and Pierre Ouellette. The measurement of productivity and scarcity rents: The case of asbestos in Canada. Journal of Econometrics, 48(3):287–312, June 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190065L.

Lobato:1999:STS

[Lob99] Ignacio N. Lobato. A semiparametric two-step estimator in a multivariate long memory model. Journal of Econometrics, 90(1): 129–153, May 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000384.

Lippi:1994:VAN

[LR94] Marco Lippi and Lucrezia Reichlin. VAR analysis, nonfundamental representations, Blaschke matrices. Journal of Econometrics, 63(1):307–325, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769301570C.

Liu:1996:MNA

[LR96a] Chuanhai Liu and Donald B. Rubin. Markov-normal analysis of iterative simulations before their convergence. Journal of Econometrics, 75(1):69–78, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017690.

Lobato:1996:APE

[LR96b] I. Lobato and P. M. Robinson. Averaged periodogram estimation of long memory. Journal of Econometrics, 73(1):303–324, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017429.

Laroque:1994:ECD

[LS94] Guy Laroque and Bernard Salanié. Estimating the canonical disequilibrium model: Asymptotic theory and finite sample properties. Journal of Econometrics, 62(2):165–210, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900213.

Lee:1996:PKT

[LS96a] Dongin Lee and Peter Schmidt. On the power of the KPSS test of stationarity against fractionally-integrated alternatives. Journal of Econometrics, 73(1):285–302, July 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017410.

Li:1996:SEP

[LS96b] Qi Li and Thanasis Stengos. Semiparametric estimation of partially linear panel data models. Journal of Econometrics, 71 (1–2):389–397, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694017115.

Levy:1997:CTI

[LS97a] Haim Levy and Gideon Schwarz. Correlation and the time interval over which the variables are measured. Journal of Econometrics, 76(1–2):341–350, January/February 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769501795X.

Lutkepohl:1997:IRA

[LS97b] Helmut Lütkepohl and Pentti Saikkonen. Impulse response analysis in infinite order cointegrated vector autoregressive processes. Journal of Econometrics, 81(1):127–157, November 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000377.

Lin:1994:TCR

[LT94] Chien-Fu Jeff Lin and Timo Teräsvirta. Testing the constancy of regression parameters against continuous structural change. Journal of Econometrics, 62(2):211–228, June 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900221.

Lee:1996:CTC

[LT96] Tae-Hwy Lee and Yiuman Tse. Cointegration tests with conditional heteroskedasticity. Journal of Econometrics, 73(2):401– 410, August 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407695017453.

Lin:1999:TPC

[LT99] Chien-Fu Jeff Lin and Timo Teräsvirta. Testing parameter constancy in linear models against stochastic stationary parameters. Journal of Econometrics, 90(2):193–213, June 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000414.

Lubrano:1995:TUR

[Lub95] Michel Lubrano. Testing for unit roots in a Bayesian framework. Journal of Econometrics, 69(1):81–109, September 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401663K.

Lucas:1995:ORU

[Luc95] André Lucas. An outlier robust unit root test with an application to the extended Nelson-plosser data. Journal of Econometrics, 66(1–2):153–173, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016135.

Lutkepohl:1997:NDM

[Lüt97] Helmut Lütkepohl. Nonparametric dynamic modelling. Journal of Econometrics, 81(1):1–5, November 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000298.

Leung:1990:NHE

[LW90] Siu Fai Leung and Wing Hung Wong. Nonparametric hazard estimation with time-varying discrete covariates. Journal of Econometrics, 45(3):309–330, ???? 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090002B.

Lanot:1998:UNU

[LW98a] Gauthier Lanot and Ian Walker. The union/non-union wage differential: An application of semi-parametric methods. Journal of Econometrics, 84(2):327–349, June 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000894.

Li:1998:SCB

[LW98b] Q. Li and Suojin Wang. A simple consistent bootstrap test for a parametric regression function. Journal of Econometrics, 87(1):145–165, November 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407698000116.

Lee:1993:TNN

[LWG93] Tae-Hwy Lee, Halbert White, and Clive W. J. Granger. Testing for neglected nonlinearity in time series models: A comparison of neural network methods and alternative tests. Journal of Econometrics, 56(3):269–290, April 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390122L.

Leung:1996:CBS

[LY96] Siu Fai Leung and Shihti Yu. On the choice between sample selection and two-part models. Journal of Econometrics, 72(1–2):197– 229, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694017204.

Maasoumi:1990:HLM

[Maa90] Esfandiar Maasoumi. How to live with misspecification if you must. Journal of Econometrics, 44(1–2):67–86, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690900733.

Maasoumi:1991:EI

[Maa91] Esfandiar Maasoumi. Editor’s introduction. Journal of Econometrics, 50(1–2):1–5, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769190085R.

Maasoumi:1992:FOR

[Maa92] Esfandiar Maasoumi. Fellow’s opinion: Rules of thumb and pseudo-science. Journal of Econometrics, 53(1–3):1–4, July/ September 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692900764.

Magdalinos:1990:CPT

[Mag90] Michael A. Magdalinos. The classical principles of testing using instrumental variables estimates. Journal of Econometrics, 44(3): 241–279, June 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900593.

Maindiratta:1990:LSE

[Mai90] Ajay Maindiratta. Largest size-efficient scale and size efficiencies of decision-making units in data envelopment analysis. Journal of Econometrics, 46(1–2):57–72, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090047W.

Manski:1993:DCS

[Man93] Charles F. Manski. Dynamic choice in social settings: Learning from the experiences of others. Journal of Econometrics, 58(1–2): 121–136, July 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390115L.

Maravall:1993:SLT

[Mar93] Agustin Maravall. Stochastic linear trends: Models and estimators. Journal of Econometrics, 56(1–2):5–37, March 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390099Q.

Marmol:1998:SRT

[Mar98] Francesc Marmol. Spurious regression theory with nonstationary fractionally integrated processes. Journal of Econometrics, 84(2): 233–250, June 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000857.

Matzkin:1993:NIE

[Mat93] Rosa L. Matzkin. Nonparametric identification and estimation of polychotomous choice models. Journal of Econometrics, 58(1–2): 137–168, July 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390116M.

McDonald:1990:RMP

[MB90] James B. McDonald and Richard J. Butler. Regression models for positive random variables. Journal of Econometrics, 43(1–2):227–251, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090118D.

McAleer:1995:STE

[McA95] Michael McAleer. The significance of testing empirical nonnested models. Journal of Econometrics, 67(1):149–171, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694016319.

McCall:1994:SDD

[McC94] Brian P. McCall. Specification diagnostics for duration models: A martingale approach. Journal of Econometrics, 60(1– 2):293–312, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900485.

McCarthy:1998:FSM

[McC98] Michael D. McCarthy. Finite sample moments results for the quasi-FIML estimator of the reduced form: The linear case. Journal of Econometrics, 83(1–2):239–262, March/April 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000717.

McManus:1992:HCI

[McM92] Douglas A. McManus. How common is identification in parametric models? Journal of Econometrics, 53(1–3):5–23, July/ September 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692900775.

Mayer:1998:MSE

[MD98] Walter J. Mayer and Robert E. Dorsey. Maximum score estimation of disequilibrium models and the role of anticipatory price-setting. Journal of Econometrics, 87(1):1–24, November 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000037.

Mensah:1994:SKD

[Men94] Yaw M. Mensah. A simplification of the Kopp–Diewert method of decomposing cost efficiency and some implications. Journal of Econometrics, 60(1–2):133–144, January/February 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769490040X.

Metcalf:1996:STP

[Met96] Gilbert E. Metcalf. Specification testing in panel data with instrumental variables. Journal of Econometrics, 71(1–2):291–307, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694017069.

Mikhail:1991:PEF

[MG91] William M. Mikhail and G. A. Ghazal. On a pooled estimator and its finite-sample moments. Journal of Econometrics, 48(1–2):195– 214, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190038F.

Michelis:1999:DJC

[Mic99] Leo Michelis. The distributions of the j and Cox non-nested tests in regression models with weakly correlated regressors.

Journal of Econometrics, 93(2):369–401, December 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407699000263.

Mills:1992:BPT

[Mil92] Jeffrey A. Mills. Bayesian prediction tests for structural stability. Journal of Econometrics, 52(3):381–388, June 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290018M.

Mirowski:1995:TWT

[Mir95] Philip Mirowski. Three ways to think about testing in econometrics. Journal of Econometrics, 67(1):25–46, May 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401625A.

Mizon:1995:SMA

[Miz95] Grayham E. Mizon. A simple message for autocorrelation correctors: Don’t. Journal of Econometrics, 69(1):267–288, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401671L.

Maravall:1994:EUM

[MM94] Agustín Maravall and Alexandre Mathis. Encompassing univariate models in multivariate time series: A case study. Journal of Econometrics, 61(2):197–233, April 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900841.

Mandy:1993:SUR

[MMF93] David M. Mandy and Carlos Martins-Filho. Seemingly unrelated regressions under additive heteroscedasticity: Theory and share equation applications. Journal of Econometrics, 58(3):315– 346, August 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390048A.

Mandy:1999:REE

[MMF99] D. M. Mandy and Carlos Martins-Filho. Relative efficiency with equivalence classes of asymptotic covariances. Journal of Econometrics, 88(1):79–98, January 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000220.

Maddala:1995:UCP

[MN95] G. S. Maddala and M. Nimalendran. An unobserved component panel data model to study the effect of earnings surprises on stock prices, trading volumes, and spreads. Journal of Econometrics, 68(1):229–242, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401650O.

McClellan:1997:MCE

[MN97] Mark McClellan and Joseph P. Newhouse. The marginal cost-effectiveness of medical technology: A panel instrumentalvariables approach. Journal of Econometrics, 77(1):39–64, March 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696018052.

McManus:1994:MOA

[MNS94] Douglas A. McManus, John C. Nankervis, and N. E. Savin. Multiple optima and asymptotic approximations in the partial adjustment model. Journal of Econometrics, 62(2):91–128, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900183.

Moffitt:1993:IED

[Mof93] Robert Moffitt. Identification and estimation of dynamic models with a time series of repeated cross-sections. Journal of Econometrics, 59(1–2):99–123, September 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900413.

Moulton:1991:BAR

[Mou91] Brent R. Moulton. A Bayesian approach to regression selection and estimation, with application to a price index for radio services. Journal of Econometrics, 49(1–2):169–193, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407691900134.

Moulton:1995:IIC

[Mou95] Brent R. Moulton. Interarea indexes of the cost of shelter using hedonic quality adjustment techniques. Journal of Econometrics, 68(1):181–204, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401648J.

Maravall:1999:EES

[MP99] Agustín Maravall and Christophe Planas. Estimation error and the specification of unobserved component models. Journal of Econometrics, 92(2):325–353, October 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000943.

McCulloch:1991:BAT

[MR91] Robert McCulloch and Peter E. Rossi. A Bayesian approach to testing the arbitrage pricing theory. Journal of Econometrics, 49(1–2):141–168, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407691900123.

Meghir:1992:FPE

[MR92] Costas Meghir and Jean-Marc Robin. Frequency of purchase and the estimation of demand systems. Journal of Econometrics, 53(1–3):53–85, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290080B.

McCulloch:1994:ELA

[MR94] Robert McCulloch and Peter E. Rossi. An exact likelihood analysis of the multinomial probit model. Journal of Econometrics, 64(1–2):207–240, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900647.

Magee:1998:USW

[MRB98] L. Magee, A. L. Robb, and J. B. Burbidge. On the use of sampling weights when estimating regression models with survey data. Journal of Econometrics, 84(2):251–271, June 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000869.

Mroz:1999:DFA

[Mro99] Thomas A. Mroz. Discrete factor approximations in simultaneous equation models: Estimating the impact of a dummy endogenous variable on a continuous outcome. Journal of Econometrics, 92(2):233–274, October 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000918.

McCurdy:1992:CRP

[MS92] Thomas H. McCurdy and Thanasis Stengos. A comparison of risk-premium forecasts implied by parametric versus nonparametric conditional mean estimators. Journal of Econometrics, 52(1–2):225–244, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290071X.

Morimune:1993:MTS

[MS93] Kimio Morimune and Shinichi Sakata. Modified three-stage least squares estimator which is third-order efficient. Journal of Econometrics, 57(1–3):257–276, May/June 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390067F.

Magdalinos:1995:ASC

[MS95] Michael A. Magdalinos and Spyridon D. Symeonides. Alternative size corrections for some GLS test statistics the case of the AR(1) model. Journal of Econometrics, 66(1–2):35–59, March/ April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694016072.

Magdalinos:1996:RTO

[MS96] Michael A. Magdalinos and Spyridon D. Symeonides. A reinterpretation of the tests of overidentifying restrictions. Journal of Econometrics, 73(2):325–353, August 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017283.

Morduch:1997:UMM

[MS97] Jonathan J. Morduch and Hal S. Stern. Using mixture models to detect sex bias in health outcomes in Bangladesh. Journal of Econometrics, 77(1):259–276, March 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696018155.

MacKinnon:1998:ABC

[MS98] James G. MacKinnon and Anthony A. Smith. Approximate bias correction in econometrics. Journal of Econometrics, 85(2):205– 230, August 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000997.

Melino:1990:PFC

[MT90] Angelo Melino and Stuart M. Turnbull. Pricing foreign currency options with stochastic volatility. Journal of Econometrics, 45 (1–2):239–265, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407690901008.

Meghir:1997:LMT

[MW97] Costas Meghir and Edward Whitehouse. Labour market transitions and retirement of men in the UK. Journal of Econometrics, 79(2):327–354, August 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000262.

Martin:1999:IEA

[MW99] Vance L. Martin and Nigel P. Wilkins. Indirect estimation of ARFIMA and VARFIMA models. Journal of Econometrics, 93(1):149–175, November 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440769900007X.

Wong:1996:BAA

[mWK96] Chi ming Wong and Robert Kohn. A Bayesian approach to additive semiparametric regression. Journal of Econometrics, 74(2):209–235, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017437.

Manton:1991:SMI

[MWS91] Kenneth G. Manton, Max A. Woodbury, and Eric Stallard. Statistical and measurement issues in assessing the welfare status of aged individuals and populations. Journal of Econometrics, 50(1–2):151–181, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190094T.

McDonald:1995:GBDa

[MX95a]  James B. McDonald and Yexiao J. Xu.  A generaliza-

tion of the beta distribution with applications. Journal of Econometrics, 66(1–2):133–152, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016124.

McDonald:1995:GBDb

[MX95b]  James B. McDonald and Yexiao J. Xu.  A generaliza-

tion of the beta distribution with applications. Journal of Econometrics, 69(2):427–428, October 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401680X.

Maasoumi:1990:GEM

[MZ90] Esfandiar Maasoumi and Sourushe Zandvakili. Generalized entropy measures of mobility for different sexes and income levels. Journal of Econometrics, 43(1–2):121–133, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690901116.

Nawata:1990:REBb

[Naw90a] Kazumitsu Nawata. Robust estimation based on grouped-adjusted data in censored regression models. Journal of Econometrics, 43(3):337–362, March 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090124C.

Nawata:1990:REBa

[Naw90b] Kazumitsu Nawata. Robust estimation based on grouped-adjusted data in linear regression models. Journal of Econometrics, 43(3): 317–336, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090123B.

Nelson:1990:AMD

[Nel90] Daniel B. Nelson. ARCH models as diffusion approximations. Journal of Econometrics, 45(1–2):7–38, July/August 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690900928.

Nelson:1992:FFM

[Nel92] Daniel B. Nelson. Filtering and forecasting with misspecified ARCH models i: Getting the right variance with the wrong model. Journal of Econometrics, 52(1–2):61–90, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290065Y.

Nelson:1996:AFT

[Nel96] Daniel B. Nelson. Asymptotic filtering theory for multivariate ARCH models. Journal of Econometrics, 71(1–2):1–47, March/ April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694016798.

Ng:1992:MDF

[NER92] Victor Ng, Robert F. Engle, and Michael Rothschild. A multi-dynamic-factor model for stock returns. Journal of Econometrics, 52(1–2):245–266, April/May 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290072Y.

Newey:1997:CRA

[New97] Whitney K. Newey. Convergence rates and asymptotic normality for series estimators. Journal of Econometrics, 79(1):147– 168, July 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000110.

Nelson:1995:FFM

[NF95] Daniel B. Nelson and Dean P. Foster. Filtering and forecasting with misspecified ARCH models II: Making the right forecast with the wrong model. Journal of Econometrics, 67(2):303– 335, June 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401635D.

Nerlove:1998:QRE

[NF98] Marc Nerlove and Ilaria Fornari. Quasi-rational expectations, an alternative to fully rational expectations: An application to US beef cattle supply. Journal of Econometrics, 83(1–2):129–161, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000675.

Nakamura:1998:MSE

[NN98] Alice Nakamura and Masao Nakamura. Model specification and endogeneity. Journal of Econometrics, 83(1–2):213–237, March/ April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000705.

Nowak:1993:IML

[Now93] Eugen Nowak. The identification of multivariate linear dynamic errors-in-variables models. Journal of Econometrics, 59(3):213– 227, October 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390023X.

Newey:1993:EBS

[NP93] Whitney K. Newey and James L. Powell. Efficiency bounds for some semiparametric selection models. Journal of Econometrics, 58(1–2):169–184, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390117N.

Nabeya:1994:LAD

[NP94] Seiji Nabeya and Pierre Perron. Local asymptotic distribution related to the AR(1) model with dependent errors. Journal of Econometrics, 62(2):229–264, June 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769490023X.

Ng:1997:EIN

[NP97] Serena Ng and Pierre Perron. Estimation and inference in nearly unbalanced nearly cointegrated systems. Journal of Econometrics, 79(1):53–81, July 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697000079.

Neudecker:1995:HLR

[NPL95] Heinz Neudecker, Wolfgang Polasek, and Shuangzhe Liu. The heteroskedastic linear regression model and the Hadamard product a note. Journal of Econometrics, 68(2):361–366, August 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401655J.

Nerlove:1993:ISI

[NRW93] Marc Nerlove, David Ross, and Douglas Willson. The importance of seasonality in inventory models: Evidence from business survey data. Journal of Econometrics, 55(1–2):105–128, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390006Q.

Neuman:1994:ELM

[NS94] Shoshana Neuman and Jacques Silber. The econometrics of labor market segregation and discrimination. Journal of Econometrics, 61(1):1–4, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900736.

Nijman:1996:MCA

[NS96] Theo Nijman and Enrique Sentana. Marginalization and contemporaneous aggregation in multivariate GARCH processes. Journal of Econometrics, 71(1–2):71–87, March/April 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401695X.

Nabeya:1990:LPU

[NT90] Seiji Nabeya and Katsuto Tanaka. Limiting power of unitroot tests in time-series regression. Journal of Econometrics, 46(3):247–271, December 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090010Q.

Nijman:1990:ETD

[NV90] Theo Nijman and Marno Verbeek. Estimation of time-dependent parameters in linear models using cross-sections, panels, or both. Journal of Econometrics, 46(3):333–346, December 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090013J.

Nijman:1992:OCC

[NV92] Theo Nijman and Marno Verbeek. The optimal choice of controls and pre-experimental observations. Journal of Econometrics, 51(1–2):183–189, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290034O.

Niesing:1994:UEM

[NvPV94] Willem Niesing, Bernard M. S. van Praag, and Justus Veenman. The unemployment of ethnic minority groups in the Netherlands. Journal of Econometrics, 61(1):173–196, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900833.

Nijman:1991:ERP

[NVvS91] Theo Nijman, Marno Verbeek, and Arthur van Soest. The efficiency of rotating-panel designs in an analysis-of-variance model. Journal of Econometrics, 49(3):373–399, September 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190003V.

Ohtani:1993:TLR

[OG93] Kazuhiro Ohtani and Judith Giles. Testing linear restrictions on coefficients in a linear regression model with proxy variables and spherically symmetric disturbances. Journal of Econometrics, 57(1–3):393–406, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390073E.

Ohtani:1990:ETL

[Oht90] Kazuhiro Ohtani. On estimating and testing in a linear regression model with autocorrelated errors. Journal of Econometrics, 44(3): 333–346, June 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090062X.

Ohtani:1999:ISR

[Oht99] Kazuhiro Ohtani. Inadmissibility of the Stein-rule estimator under the balanced loss function. Journal of Econometrics, 88(1):193– 201, January 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769800030X.

Ozcam:1991:SRR

[ÖJ91] Ahmet Özcam and George G. Judge. Some risk results for a twostage pre-test estimator in the case of possible heteroskedasticity. Journal of Econometrics, 48(3):355–371, June 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190068O.

Ohtani:1996:EGF

[OK96] Kazuhiro Ohtani and Hideo Kozumi. The exact general formulae for the moments and the MSE dominance of the Stein-rule and positive-part Stein-rule estimators. Journal of Econometrics, 74(2):273–287, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769501747X.

Ogaki:1997:CAE

[OP97] Masao Ogaki and Joon Y. Park. A cointegration approach to estimating preference parameters. Journal of Econometrics, 82(1):

107–134, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000535.

Oaxaca:1994:DDW

[OR94] Ronald L. Oaxaca and Michael R. Ransom. On discrimination and the decomposition of wage differentials. Journal of Econometrics, 61(1):5–21, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900744.

Orme:1990:SSP

[Orm90] Chris Orme. The small-sample performance of the informationmatrix test. Journal of Econometrics, 46(3):309–331, December 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090012I.

Oliner:1996:LCR

[ORS96] Stephen D. Oliner, Glenn D. Rudebusch, and Daniel Sichel. The Lucas critique revisited assessing the stability of empirical Euler equations for investment. Journal of Econometrics, 70(1):291– 316, January 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694016933.

Osiewalski:1993:BME

[OS93a] Jacek Osiewalski and Mark F. J. Steel. Bayesian marginal equivalence of elliptical regression models. Journal of Econometrics, 59(3):391–403, October 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390032Z.

Osiewalski:1993:RBI

[OS93b] Jacek Osiewalski and Mark F. J. Steel. Robust Bayesian inference in elliptical regression models. Journal of Econometrics, 57(1–3):345–363, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390070L.

Osborn:1991:IPV

[Osb91] Denise R. Osborn. The implications of periodically varying coefficients for seasonal time-series processes. Journal of Econometrics, 48(3):373–384, June 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190069P.

Osborn:1993:SC

[Osb93] Denise R. Osborn. Seasonal cointegration. Journal of Econometrics, 55(1–2):299–303, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390017Y.

Osiewalski:1991:NBI

[Osi91] Jacek Osiewalski. A note on Bayesian inference in a regression model with elliptical errors. Journal of Econometrics, 48(1–2):183–193, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769190037E.

Paarsch:1992:DBC

[Paa92] Harry J. Paarsch. Deciding between the common and private value paradigms in empirical models of auctions. Journal of Econometrics, 51(1–2):191–215, January/February 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290035P.

Paarsch:1997:DEOa

[Paa97a] Harry J. Paarsch. Deriving an estimate of the optimal reserve price: An application to British Columbian timber sales. Journal of Econometrics, 78(1):333–357, ???? 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800164.

Paarsch:1997:DEOb

[Paa97b] Harry J. Paarsch. Deriving an estimate of the optimal reserve price: An application to British Columbian timber sales. Journal of Econometrics, 78(2):333–357, June 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000171.

Pai:1997:BAC

[Pai97] Jeffrey S. Pai. Bayesian analysis of compound loss distributions. Journal of Econometrics, 79(1):129–146, July 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000109.

Palm:1995:BMS

[Pal95] Franz C. Palm. Bayesian model selection and prediction with empirical applications comments. Journal of Econometrics, 69(1):333–335, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401673N.

Paruolo:1996:DII

[Par96] Paolo Paruolo. On the determination of integration indices in i(2) systems. Journal of Econometrics, 72(1–2):313–356, May/ June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017259.

Pelloni:1996:FFM

[Pel96] Gianluigi Pelloni. De Finetti, Friedman, and the methodology of positive economics. Journal of Econometrics, 75(1):33–50, November 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017674.

Pendakur:1999:SET

[Pen99] Krishna Pendakur. Semiparametric estimates and tests of baseindependent equivalence scales. Journal of Econometrics, 88(1): 1–40, January 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000207.

Peracchi:1990:BIE

[Per90] Franco Peracchi. Bounded-influence estimators for the Tobit model. Journal of Econometrics, 44(1–2):107–126, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690900755.

Peracchi:1991:BIE

[Per91] Franco Peracchi. Bounded-influence estimators for the SURE model. Journal of Econometrics, 48(1–2):119–134, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190034B.

Perron:1996:AAA

[Per96] Pierre Perron. The adequacy of asymptotic approximations in the near-integrated autoregressive model with dependent errors. Journal of Econometrics, 70(2):317–350, February 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016577.

Perron:1997:FEB

[Per97] Pierre Perron. Further evidence on breaking trend functions in macroeconomic variables. Journal of Econometrics, 80(2):355– 385, October 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000493.

Pantula:1991:TUR

[PH91] Sastry G. Pantula and Alastair Hall. Testing for unit roots in autoregressive moving average models: An instrumental variable approach. Journal of Econometrics, 48(3):325–353, June 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190067N.

Porter-Hudak:1991:NMA

[PHH91] Susan Porter-Hudak and Kathy Hayes. A numerical methods approach to calculating cost-of-living indices. Journal of Econometrics, 50(1–2):91–105, October 11, 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190091Q.

Phillips:1991:CML

[Phi91] Robert F. Phillips. A constrained maximum-likelihood approach to estimating switching regressions. Journal of Econometrics, 48(1–2):241–262, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769190040K.

See note [XTZ10].

Phillips:1994:PAE

[Phi94] Robert F. Phillips. Partially adaptive estimation via a normal mixture. Journal of Econometrics, 64(1–2):123–144, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900604.

Phillips:1995:BMS

[Phi95a] Peter C. B. Phillips. Bayesian model selection and prediction with empirical applications. Journal of Econometrics, 69(1):289–331, September 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401672M.

Phillips:1995:BPR

[Phi95b] Peter C. B. Phillips. Bayesian prediction a response. Journal of Econometrics, 69(1):351–365, September 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401675P.

Philipson:1997:ENH

[Phi97] Tomas Philipson. The evaluation of new health care technology: The labor economics of statistics. Journal of Econometrics, 76(1–2):375–395, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017984.

Phillips:1998:IRF

[Phi98] Peter C. B. Phillips. Impulse response and forecast error variance asymptotics in nonstationary VARs. Journal of Econometrics, 83(1–2):21–56, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769700064X.

Pinkse:1993:CSE

[Pin93] C. A. P. Pinkse. On the computation of semiparametric estimates in limited dependent variable models. Journal of Econometrics, 58(1–2):185–205, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390118O.

Pinkse:1998:CNT

[Pin98] Joris Pinkse. A consistent nonparametric test for serial independence. Journal of Econometrics, 84(2):205–231, June 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000845.

Pope:1996:EIE

[PJ96] Rulon D. Pope and Richard E. Just. Empirical implementation of ex ante cost functions. Journal of Econometrics, 72(1–2):231– 249, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694017212.

Polachek:1994:PEG

[PK94] Solomon W. Polachek and Moon-Kak Kim. Panel estimates of the gender earnings gap: Individual-specific intercept and individual-specific slope models. Journal of Econometrics, 61(1): 23–42, March 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900752.

Ploberger:1996:TRT

[PK96] Werner Ploberger and Walter Krämer. A trend-resistant test for structural change based on OLS residuals. Journal of Econometrics, 70(1):175–185, January 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016887.

Phillips:1991:DWR

[PL91] Peter C. B. Phillips and Mico Loretan. The Durbin–Watson ratio under infinite-variance errors. Journal of Econometrics, 47(1):85– 114, January 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769190079S.

Philipson:1999:MEP

[PM99] Tomas Philipson and Anup Malani. Measurement errors: A principal investigator-agent approach. Journal of Econometrics, 91(2):273–298, August 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407698000785.

Prucha:1996:ECU

[PN96] Ingmar R. Prucha and M. Ishaq Nadiri. Endogenous capital utilization and productivity measurement in dynamic factor demand models theory and an application to the U.S. electrical machinery industry. Journal of Econometrics, 71(1–2):343–379, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694017093.

Pakes:1995:LTS

[PO95] Ariel Pakes and Steven Olley. A limit theorem for a smooth class of semiparametric estimators. Journal of Econometrics, 65(1):295– 332, January 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401605Y.

Poirier:1991:EI

[Poi91] Dale J. Poirier. Editor’s introduction. Journal of Econometrics, 49(1–2):1–4, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407691900082.

Poirier:1994:JPL

[Poi94] Dale Poirier. Jeffreys’ prior for logit models. Journal of Econometrics, 63(2):327–339, August 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407693015562.

Poirier:1996:BAN

[Poi96] Dale J. Poirier. A Bayesian analysis of nested logit models. Journal of Econometrics, 75(1):163–181, November 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017755.

Poirier:1997:CCBa

[Poi97a] Dale J. Poirier. Comparing and choosing between two models with a third model in the background. Journal of Econometrics, 78(1): 139–151, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697800061.

Poirier:1997:CCBb

[Poi97b] Dale J. Poirier. Comparing and choosing between two models with a third model in the background. Journal of Econometrics, 78(2): 139–151, June 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696000024.

Pollak:1991:WCS

[Pol91] Robert A. Pollak. Welfare comparisons and situation comparisons. Journal of Econometrics, 50(1–2):31–48, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190088U.

Potscher:1995:CAE

[Pöt95] Benedikt M. Pötscher. Comment on ‘Adaptive estimation in time series regression models’ by D. G. Steigerwald. Journal of Econometrics, 66(1–2):123–129, March/April 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

030440769401610C. See [Ste92a] reply [Ste95].

Pesaran:1993:SAP

[PP93] M. Hashem Pesaran and Bahram Pesaran. A simulation approach to the problem of computing Cox’s statistic for testing nonnested models. Journal of Econometrics, 57(1–3):377–392, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390072D.

Potscher:1994:GUC

[PP94] Benedikt M. Pötscher and Ingmar R. Prucha. Generic uniform convergence and equicontinuity concepts for random functions: An exploration of the basic structure. Journal of Econometrics, 60(1–2):23–63, January/February 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769490037X.

Palm:1998:SAP

[PP98] Franz C. Palm and Gerard A. Pfann. Sources of asymmetry in production factor dynamics. Journal of Econometrics, 82(2):361– 392, February 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769700078X.

Pesaran:1993:PCA

[PPL93] M. H. Pesaran, R. G. Pierse, and K. C. Lee. Persistence, cointegration, and aggregation: A disaggregated analysis of output fluctuations in the U.S. economy. Journal of Econometrics, 56(1–2): 57–88, March 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390101A.

Paruolo:1999:WEV

[PR99] Paolo Paruolo and Anders Rahbek. Weak exogeneity in i(2) VAR systems. Journal of Econometrics, 93(2):281–308, December 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407699000123.

Politis:1997:SHT

[PRW97] D. N. Politis, Joseph P. Romano, and Michael Wolf. Subsampling for heteroskedastic time series. Journal of Econometrics, 81(2):281–317, December 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697865694.

Pagan:1990:AMC

[PS90a] Adrian R. Pagan and G. William Schwert. Alternative models for conditional stock volatility. Journal of Econometrics, 45(1–2):267–290, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090101X.

Pesaran:1990:UAE

[PS90b] M. Hashem Pesaran and Richard J. Smith. A unified approach to estimation and orthogonality tests in linear single-equation econometric models. Journal of Econometrics, 44(1–2):41–66, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900722.

Peters:1991:DST

[PS91] Simon Peters and Richard J. Smith. Distributional specification tests against semiparametric alternatives. Journal of Econometrics, 47(1):175–194, January 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190083P.

Pesaran:1992:ELD

[PS92] M. Hashem Pesaran and Hossein Samiei. Estimating limiteddependent rational expectations models with an application to exchange rate determination in a target zone. Journal of Econometrics, 53(1–3):141–163, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900834.

Pesaran:1995:ELR

[PS95a]  M. Hashem Pesaran and Ron Smith.  Estimating long-

run relationships from dynamic heterogeneous panels. Journal of Econometrics, 68(1):79–113, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401644F.

Pesaran:1995:RTE

[PS95b] M. Hashem Pesaran and Ron Smith. The role of theory in econometrics. Journal of Econometrics, 67(1):61–79, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401627C.

Pierse:1995:TAP

[PS95c] R. G. Pierse and A. J. Snell. Temporal aggregation and the power of tests for a unit root. Journal of Econometrics, 65(2):333– 345, February 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769301589E.

Pesaran:1996:CSC

[PS96a] M. Hashem Pesaran and Yongcheol Shin. Cointegration and speed of convergence to equilibrium. Journal of Econometrics, 71(1–2):117–143, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016976.

Powell:1996:OBC

[PS96b] James L. Powell and Thomas M. Stoker. Optimal bandwidth choice for density-weighted averages. Journal of Econometrics, 75(2):291–316, December 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017615.

Psaradakis:1996:PTS

[PS96c] Zacharias Psaradakis and Martin Sola. On the power of tests for superexogeneity and structural invariance. Journal of Econometrics, 72(1–2):151–175, May/June 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017182.

Pinkse:1998:CSA

[PS98a] Joris Pinkse and Margaret E. Slade. Contracting in space: An application of spatial statistics to discrete-choice models. Journal of Econometrics, 85(1):125–154, July 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000973.

Psaradakis:1998:FSP

[PS98b] Zacharias Psaradakis and Martin Sola. Finite-sample properties of the maximum likelihood estimator in autoregressive models with Markov switching. Journal of Econometrics, 86(2):369– 386, October 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000104.

Park:1998:SPF

[PSS98] B. U. Park, R. C. Sickles, and L. Simar. Stochastic panel frontiers: A semiparametric approach. Journal of Econometrics, 84(2): 273–301, June 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000870.

Pfann:1996:NIR

[PST96] Gerard A. Pfann, Peter C. Schotman, and Rolf Tschernig. Nonlinear interest rate dynamics and implications for the term structure. Journal of Econometrics, 74(1):149–176, September 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017542.

Pollak:1991:LDC

[PW91] Robert A. Pollak and Terence J. Wales. The likelihood dominance criterion: A new approach to model selection. Journal of Econometrics, 47(2–3):227–242, February 3, 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190100R.

Peracchi:1995:HRM

[PW95] Franco Peracchi and Finis Welch. How representative are matched cross-sections? Evidence from the Current Population Survey. Journal of Econometrics, 68(1):153–179, July 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401647I.

Qian:1999:IIV

[QS99] Hailong Qian and Peter Schmidt. Improved instrumental variables and generalized method of moments estimators. Journal of Econometrics, 91(1):145–169, July 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000748.

Quintos:1998:ACV

[Qui98a] Carmela E. Quintos. Analysis of cointegration vectors using the GMM approach. Journal of Econometrics, 85(1):155–188, July 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000985.

Quintos:1998:STE

[Qui98b] Carmela E. Quintos. Stability tests in error correction models. Journal of Econometrics, 82(2):289–315, February 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000596.

Rayner:1991:ALI

[Ray91] Janne Rayner. Another look at the identification of current rational-expectations models. Journal of Econometrics, 47 (2–3):305–331, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769190105M.

Rodriguez-Campos:1993:NBC

[RCCA93] M. C. Rodríguez-Campos and R. Cao-Abad. Nonparametric bootstrap confidence intervals for discrete regression functions. Journal of Econometrics, 58(1–2):207–222, July 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390119P.

Reinsdorf:1999:SVI

[RD99] Marshall B. Reinsdorf and Alan H. Dorfman. The Sato-vartia index and the monotonicity axiom. Journal of Econometrics, 90 (1):45–61, May 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769800027X.

Reiss:1990:DMO

[Rei90] Peter C. Reiss. Detecting multiple outliers with an application to R&D productivity. Journal of Econometrics, 43(3):293– 315, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090122A.

Richard:1995:BMS

[Ric95] Jean-François Richard. Bayesian model selection and prediction with empirical applications discussion. Journal of Econometrics, 69(1):337–349, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769401674O.

Rilstone:1993:CLS

[Ril93] Paul Rilstone. Calculating the (local) semiparametric efficiency bounds for the generated regressors problem. Journal of Econometrics, 56(3):357–370, April 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390126P.

Rahman:1997:MLS

[RK97] Shahidur Rahman and Maxwell L. King. Marginal-likelihood score-based tests of regression disturbances in the presence of nuisance parameters. Journal of Econometrics, 82(1):81–106, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000523.

Rahbek:1999:TSC

[RKJ99] Anders Rahbek, Hans Christian Kongsted, and Clara Jørgensen. Trend stationarity in the i (2) cointegration model. Journal of Econometrics, 90(2):265–289, June 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769800044X.

Ramsey:1992:IEN

[RM92] James B. Ramsey and Alvaro Montenegro. Identification and estimation of noninvertible non-Gaussian MA(q) processes. Journal of Econometrics, 54(1–3):301–320, October/December 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290110D.

Robinson:1991:TSS

[Rob91] P. M. Robinson. Testing for strong serial correlation and dynamic conditional heteroskedasticity in multiple regression. Journal of Econometrics, 47(1):67–84, January 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190078R.

Romeo:1997:MILa

[Rom97a] Charles J. Romeo. Measuring information loss due to inconsistencies in duration data from longitudinal surveys. Journal of Econometrics, 78(1):159–177, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800085.

Romeo:1997:MILb

[Rom97b] Charles J. Romeo. Measuring information loss due to inconsistencies in duration data from longitudinal surveys. Journal of Econometrics, 78(2):159–177, June 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696000061.

Rothenberg:1990:SEC

[RR90] Thomas J. Rothenberg and Paul A. Ruud. Simultaneous equations with covariance restrictions. Journal of Econometrics, 44(1–2):25– 39, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090071Z.

Raynauld:1993:SBM

[RS93] Jacques Raynauld and Jean-Guy Simonato. Seasonal BVAR models: A search along some time domain priors. Journal of Econometrics, 55(1–2):203–229, January/February 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390012T.

Ryu:1996:TFF

[RS96] Hang K. Ryu and Daniel J. Slottje. Two flexible functional form approaches for approximating the Lorenz curve. Journal of Econometrics, 72(1–2):251–274, May/June 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017220.

Rothenberg:1997:INI

[RS97] Thomas J. Rothenberg and James H. Stock. Inference in a nearly integrated autoregressive model with nonnormal innovations. Journal of Econometrics, 80(2):269–286, October 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000407.

Rilstone:1996:SOB

[RSU96] Paul Rilstone, V. K. Srivastava, and Aman Ullah. The secondorder bias and mean squared error of nonlinear estimators. Journal of Econometrics, 75(2):369–395, December 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407696894577. See corrigendum [RU05].

Ridder:1999:SPL

[RT99] Geert Ridder and Insan Tunali. Stratified partial likelihood estimation. Journal of Econometrics, 92(2):193–232, October 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000906.

Rilstone:2005:CSO

[RU05] Paul Rilstone and Aman Ullah. Corrigendum to “The secondorder bias and mean squared error of nonlinear estimators”: [Journal of Econometrics 75(2) (1996) 369–395]. Journal of Econometrics, 124(1):203–204, January 2005. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407604000521. See [RSU96].

Ruggiero:1994:BSE

[Rug94] Michèle Ruggiero. Bayesian semiparametric estimation of proportional hazards models. Journal of Econometrics, 62(2):277– 300, June 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900256.

Ruiz:1994:QML

[Rui94] Esther Ruiz. Quasi-maximum likelihood estimation of stochastic volatility models. Journal of Econometrics, 63(1):289–306, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693015698. See [AS97b, Rui97].

Ruiz:1997:QGE

[Rui97] Esther Ruiz. QML and GMM estimators of stochastic volatility models: Response to Andersen and Sørensen. Journal of Econometrics, 76(1–2):405, January/February 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

030440769501800X. See [Rui94, AS97b].

Runde:1997:ANDa

[Run97a] Ralf Runde. The asymptotic null distribution of the Box–Pierce Q-statistic for random variables with infinite variance: An application to German stock returns. Journal of Econometrics, 78(1): 205–216, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697800103.

Runde:1997:ANDb

[Run97b] Ralf Runde. The asymptotic null distribution of the Box–Pierce Q-statistic for random variables with infinite variance an application to German stock returns. Journal of Econometrics, 78(2): 205–216, June 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696000085.

Ruud:1991:EEM

[Ruu91] Paul A. Ruud. Extensions of estimation methods using the EM algorithm. Journal of Econometrics, 49(3):305–341, September 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190001T.

Rosenzweig:1991:IBS

[RW91] Mark R. Rosenzweig and Kenneth I. Wolpin. Inequality at birth: The scope for policy intervention. Journal of Econometrics, 50(1–2):205–228, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190096V.

Ryu:1993:MEE

[Ryu93a] Hang K. Ryu. Maximum entropy estimation of density and regression functions. Journal of Econometrics, 56(3):397–440, April 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390128R.

Ryu:1993:SDA

[Ryu93b] Keunkwan Ryu. Structural duration analysis of management data. Journal of Econometrics, 57(1–3):91–115, May/June 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390060I.

Said:1991:URT

[Saï91] Saïd E. Saïd. Unit-roots test for time-series data with a linear time trend. Journal of Econometrics, 47(2–3):285–303, February 3, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190104L.

Sepanski:1993:SQV

[SC93] J. H. Sepanski and R. J. Carroll. Semiparametric quasilikelihood and variance function estimation in measurement error models. Journal of Econometrics, 58(1–2):223–256, July 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769390120T.

Schmidt:1990:TSL

[Sch90] Peter Schmidt. Three-stage least squares with different instruments for different equations. Journal of Econometrics, 43(3): 389–394, March 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090127F.

Schmertmann:1994:SBC

[Sch94] Carl P. Schmertmann. Selectivity bias correction methods in polychotomous sample selection models. Journal of Econometrics, 60(1–2):101–132, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900396.

Schotman:1996:BAE

[Sch96] Peter Schotman. A Bayesian approach to the empirical valuation of bond options. Journal of Econometrics, 75(1):183–215, November 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017763.

Schultz:1997:APB

[Sch97] T. Paul Schultz. Assessing the productive benefits of nutrition and health: An integrated human capital approach. Journal of Econometrics, 77(1):141–158, March 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696018106.

Sarabia:1999:OFL

[SCS99] J.-M. Sarabia, Enrique Castillo, and Daniel J. Slottje. An ordered family of Lorenz curves. Journal of Econometrics, 91(1): 43–60, July 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000487.

Sengupta:1990:TSD

[Sen90] Jati K. Sengupta. Transformations in stochastic DEA models. Journal of Econometrics, 46(1–2):109–123, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407690900504.

Sen:1991:WPF

[Sen91] Amartya Sen. Welfare, preference and freedom. Journal of Econometrics, 50(1–2):15–29, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190087T.

Seo:1998:SIC

[Seo98] Byeongseon Seo. Statistical inference on cointegration rank in error correction models with stationary covariates. Journal of Econometrics, 85(2):339–385, August 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769700105X.

Seo:1999:DTU

[Seo99] Byeongseon Seo. Distribution theory for unit root tests with conditional heteroskedasticity 1. Journal of Econometrics, 91(1):

113–144, July 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000517.

Sharma:1991:PRJ

[SG91] Subhash C. Sharma and Carmelo Giaccotto. Power and robustness of jackknife and likelihood-ratio tests for grouped heteroscedasticity. Journal of Econometrics, 49(3):343–372, September 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190002U.

Shanken:1990:IAP

[Sha90] Jay Shanken. Intertemporal asset pricing: An empirical investigation. Journal of Econometrics, 45(1–2):99–120, July/August 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090095B.

Shafer:1996:SJB

[Sha96] Glenn Shafer. The significance of Jacob Bernoulli’s ars conjectandi for the philosophy of probability today. Journal of Econometrics, 75(1):15–32, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017666.

Shephard:1994:LSM

[She94] Neil Shephard. Local scale models: State space alternative to integrated GARCH processes. Journal of Econometrics, 60(1–2):181–202, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900434.

Slottje:1994:NMD

[SHHS94] Daniel J. Slottje, Joseph G. Hirschberg, Kathy J. Hayes, and Gerald W. Scully. A new method for detecting individual and group labor market discrimination. Journal of Econometrics, 61(1):43– 64, March 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900760.

Shively:1993:TAD

[Shi93] Thomas S. Shively. Testing for autoregressive disturbances in a time series regression with missing observations. Journal of Econometrics, 57(1–3):233–255, May/June 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390066E.

Sims:1993:REM

[Sim93] Christopher A. Sims. Rational expectations modeling with seasonally adjusted data. Journal of Econometrics, 55(1–2):9–19, January/February 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390002M.

Sims:1998:EIG

[Sim98] Christopher A. Sims. Econometric implications of the government budget constraint. Journal of Econometrics, 83(1–2):9–19, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000638.

Silvapulle:1993:NTA

[SK93] Paramsothy Silvapulle and Maxwell L. King. Nonnested testing for autocorrelation in the linear regression model. Journal of Econometrics, 58(3):295–314, August 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407693900479.

Smith:1996:NRU

[SK96] Michael Smith and Robert Kohn. Nonparametric regression using Bayesian variable selection. Journal of Econometrics, 75(2):317– 343, December 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017631.

Shively:1997:BAM

[SK97] Thomas S. Shively and Robert Kohn. A Bayesian approach to model selection in stochastic coefficient regression models and structural time series models. Journal of Econometrics, 76(1–2):39–52, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769501781X.

Sandmann:1998:ESV

[SK98] Gleb Sandmann and Siem Jan Koopman. Estimation of stochastic volatility models via Monte Carlo maximum likelihood. Journal of Econometrics, 87(2):271–301, December 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000165.

Shively:1994:TLS

[SKA94] Thomas S. Shively, Robert Kohn, and Craig F. Ansley. Testing for linearity in a semiparametric regression model. Journal of Econometrics, 64(1–2):77–96, September/October 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900582.

Saikkonen:1997:TCI

[SL97] Pentti Saikkonen and Ritva Luukkonen. Testing cointegration in infinite order vector autoregressive processes. Journal of Econometrics, 81(1):93–126, November 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000365.

Slesnick:1990:IRP

[Sle90] Daniel T. Slesnick. Inflation, relative price variation, and inequality. Journal of Econometrics, 43(1–2):135–151, January/ February 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690901127.

Slesnick:1991:NIN

[Sle91] Daniel T. Slesnick. Normative index numbers. Journal of Econometrics, 50(1–2):107–130, October 11, 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190092R.

Slottje:1990:EIS

[Slo90] Daniel Slottje. Editor’s introduction: The state of empirical work on economic inequality. Journal of Econometrics, 43 (1–2):1–3, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769090103Z.

Srivastava:1995:EPF

[SM95] V. K. Srivastava and Koichi Maekawa. Efficiency properties of feasible generalized least squares estimators in SURE models under non-normal disturbances. Journal of Econometrics, 66(1–2):99–121, March/April 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016094.

Small:1994:EPS

[Sma94a] John P. Small. The exact powers of some autocorrelation tests when the disturbances are heteroscedastic. Journal of Econometrics, 61(2):383–394, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900914.

Small:1994:AGE

[Sma94b] Kenneth A. Small. Approximate generalized extreme value models of discrete choice. Journal of Econometrics, 62(2):351–382, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900280.

Smith:1994:EDV

[Smi94] Murray D. Smith. Exact densities for variance estimators of the structural disturbances in simultaneous equations models. Journal of Econometrics, 60(1–2):157–180, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900426.

Smith:1999:FSP

[Smi99] David C. Smith. Finite sample properties of tests of the Epsteinzin asset pricing model. Journal of Econometrics, 93(1):113–148, November 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407699000068.

Snell:1999:TVC

[Sne99] Andy Snell. Testing for r versus r − 1 cointegrating vectors. Journal of Econometrics, 88(1):151–191, January 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000293.

Soofi:1990:ECI

[Soo90] Ehsan S. Soofi. Effects of collinearity on information about regression coefficients. Journal of Econometrics, 43(3):255–274, March 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090120I.

Sowell:1992:MLE

[Sow92] Fallaw Sowell. Maximum likelihood estimation of stationary univariate fractionally integrated time series models. Journal of Econometrics, 53(1–3):165–188, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900845.

Spanos:1990:SEM

[Spa90] Aris Spanos. The simultaneous-equations model revisited: Statistical adequacy and identification. Journal of Econometrics, 44(1–2):87–105, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407690900744.

Spanos:1995:TTE

[Spa95] Aris Spanos. On theory testing in econometrics: Modeling with nonexperimental data. Journal of Econometrics, 67(1):189–226, May 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401633B.

Squires:1994:FBU

[Squ94] Dale Squires. Firm behavior under input rationing. Journal of Econometrics, 61(2):235–257, April 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769490085X.

Steel:1991:BME

[SR91] Mark F. J. Steel and Jean-François Richard. Bayesian multivariate exogeneity analysis: An application to a UK money demand equation. Journal of Econometrics, 49(1–2):239–274, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407691900156.

Selvanathan:1992:EAC

[SR92] E. A. Selvanathan and D. S. Prasada Rao. An econometric approach to the construction of generalized Theil– Tornqvist indices for multilateral comparisons. Journal of Econometrics, 54(1–3):335–346, October/December 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692901125.

Segerson:1990:MEC

[SS90] Kathleen Segerson and Dale Squires. On the measurement of economic capacity utilization for multi-product industries. Journal of Econometrics, 44(3):347–361, June 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090063Y.

Salemi:1992:SSM

[SS92] Michael K. Salemi and Jaeyeong Song. Saddlepath solutions for multivariate linear rational expectations models. Journal of Econometrics, 53(1–3):245–269, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900878.

Srivastava:1997:IESa

[SS97a] Anil K. Srivastava and Shalabh. Improved estimation of the slope parameter in a linear ultrastructural model when measurement errors are not necessarily normal. Journal of Econometrics, 78(1): 153–157, ???? 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697800073.

Srivastava:1997:IESb

[SS97b] Anil K. Srivastava and Shalabh. Improved estimation of the slope parameter in a linear ultrastructural model when measurement errors are not necessarily normal. Journal of Econometrics, 78(2): 153–157, June 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407696018222.

Seiford:1990:RDD

[ST90] Lawrence M. Seiford and Robert M. Thrall. Recent developments in DEA: The mathematical programming approach to frontier analysis. Journal of Econometrics, 46(1–2):7–38, October/ November 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769090045U.

Skeels:1995:SEP

[ST95] Christopher L. Skeels and Larry W. Taylor. On a simultaneous equations pre-test estimator. Journal of Econometrics, 68(2):269– 286, August 1995. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769401638G.

Smith:1998:ACV

[ST98] Richard J. Smith and A. M. Robert Taylor. Additional critical values and asymptotic representations for seasonal unit root tests. Journal of Econometrics, 85(2):269–288, August 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697001024.

Steel:1991:BAS

[Ste91] Mark F. J. Steel. A Bayesian analysis of simultaneous equation models by combining recursive analytical and numerical approaches. Journal of Econometrics, 48(1–2):83–117, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190033A.

Steigerwald:1992:AET

[Ste92a] Douglas G. Steigerwald. Adaptive estimation in time series regression models. Journal of Econometrics, 54(1–3):251–275, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692901084. See comment [Pöt95]

and reply [Ste95].

Steigerwald:1992:FSB

[Ste92b] Douglas G. Steigerwald. On the finite sample behavior of adaptive estimators. Journal of Econometrics, 54(1–3):371–400, October/December 1992. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407692901147.

Steigerwald:1995:RBM

[Ste95] Douglas G. Steigerwald. Reply to B. M. Pötscher’s comment on ‘adaptive estimation in time series regression models’. Journal of Econometrics, 66(1–2):131–132, March/April 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0304407694016113. See [Pöt95, Ste92a].

Stern:1996:SES

[Ste96] Steven Stern. Semiparametric estimates of the supply and demand effects of disability on labor force participation. Journal of Econometrics, 71(1–2):49–70, March/April 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016941.

Stigler:1996:BB

[Sti96] Stephen M. Stigler. The Bernoullis of Basel. Journal of Econometrics, 75(1):7–13, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017658.

Stock:1994:DB

[Sto94] James H. Stock. Deciding between i(1) and i(0). Journal of Econometrics, 63(1):105–131, July 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769301562Z.

Stoker:1996:SBM

[Sto96] Thomas M. Stoker. Smoothing bias in the measurement of marginal effects. Journal of Econometrics, 72(1–2):49–84, May/ June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694017158.

Stutzer:1995:BAD

[Stu95] Michael Stutzer. A Bayesian approach to diagnosis of asset pricing models. Journal of Econometrics, 68(2):367–397, August 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769401656K.

Sueyoshi:1992:SPH

[Sue92] Glenn T. Sueyoshi. Semiparametric proportional hazards estimation of competing risks models with time-varying covariates. Journal of Econometrics, 51(1–2):25–58, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769290028P.

Skeels:1999:MCI

[SV99] Christopher L. Skeels and Francis Vella. A Monte Carlo investigation of the sampling behavior of conditional moment tests in Tobit and Probit models. Journal of Econometrics, 92(2):275– 294, October 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440769800092X.

Schotman:1991:BAU

[SvD91] Peter Schotman and Herman K. van Dijk. A Bayesian analysis of the unit root in real exchange rates. Journal of Econometrics, 49(1–2):195–238, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407691900145.

Swofford:1994:RPT

[SW94] James L. Swofford and Gerald A. Whitney. A revealed preference test for weakly separable utility maximization with incomplete adjustment. Journal of Econometrics, 60(1–2):235–249, January/February 1994. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407694900450.

Sin:1996:ICS

[SW96] Chor-Yiu Sin and Halbert White. Information criteria for selecting possibly misspecified parametric models. Journal of Econometrics, 71(1–2):207–225, March/April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017018.

Szpiro:1997:NUNa

[Szp97a] George G. Szpiro. Noise in unspecified, non-linear time series. Journal of Econometrics, 78(1):229–255, ???? 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697800127.

Szpiro:1997:NUNb

[Szp97b] George G. Szpiro. Noise in unspecified, non-linear time series. Journal of Econometrics, 78(2):229–255, June 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407696000103.

Szroeter:1994:EFS

[Szr94] Jerzy Szroeter. Exact finite-sample relative efficiency of suboptimally weighted least squares estimators in models with ordered heteroscedasticity. Journal of Econometrics, 64(1–2): 29–43, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900566.

Terza:1998:ECD

[Ter98] Joseph V. Terza. Estimating count data models with endogenous switching: Sample selection and endogenous treatment effects. Journal of Econometrics, 84(1):129–154, May 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000821.

Thompson:1993:SEB

[Tho93] T. Scott Thompson. Some efficiency bounds for semiparametric discrete choice models. Journal of Econometrics, 58(1–2):257– 274, July 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390121K.

Thoma:1994:SIA

[Tho94] Mark A. Thoma. Subsample instability and asymmetries in money-income causality. Journal of Econometrics, 64(1–2): 279–306, September/October 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694900663.

Thursby:1992:CSE

[Thu92] Jerry G. Thursby. A comparison of several exact and approximate tests for structural shift under heteroscedasticity. Journal of Econometrics, 53(1–3):363–386, July/September 1992. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900926.

Tinbergen:1991:MW

[Tin91] Jan Tinbergen. On the measurement of welfare. Journal of Econometrics, 50(1–2):7–13, October 11, 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190086S.

Thompson:1990:RMB

[TLL^{+}90] Russell G. Thompson, Larry N. Langemeier, Chih-Tah Lee, Euntaik Lee, and Robert M. Thrall. The role of multiplier bounds in efficiency analysis with application to Kansas farming. Journal of Econometrics, 46(1–2):93–108, October/November 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769090049Y.

Tsurumi:1993:ETT

[TM93] Hiroki Tsurumi and Peter Mehr. Exogeneity tests in a truncated structural equation. Journal of Econometrics, 56(3):371– 396, April 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390127Q.

Tanizaki:1998:NNG

[TM98] Hisashi Tanizaki and Roberto S. Mariano. Nonlinear and non-Gaussian state-space modeling with Monte Carlo simulations. Journal of Econometrics, 83(1–2):263–290, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697802266.

Toda:1993:SEU

[TP93] Hiro Y. Toda and Peter C. B. Phillips. The spurious effect of unit roots on vector autoregressions: An analytical study. Journal of Econometrics, 59(3):229–255, October 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390024Y.

Thomas:1997:HWE

[TS97] Duncan Thomas and John Strauss. Health and wages: Evidence on men and women in urban Brazil. Journal of Econometrics, 77(1):159–185, March 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407696018118.

Tieslau:1996:MDE

[TSB96] Margie A. Tieslau, Peter Schmidt, and Richard T. Baillie. A minimum distance estimator for long-memory processes. Journal of Econometrics, 71(1–2):249–264, March/April 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694017034.

Tsionas:1999:MCI

[Tsi99] Efthymios G. Tsionas. Monte Carlo inference in econometric models with symmetric stable disturbances. Journal of Econometrics, 88(2):365–401, February 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000396.

Torelli:1993:MIJ

[TT93] Nicola Torelli and Ugo Trivellato. Modelling inaccuracies in jobsearch duration data. Journal of Econometrics, 59(1–2):187–211, September 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407693900457.

Turkington:1998:EEL

[Tur98] Darrell A. Turkington. Efficient estimation in the linear simultaneous equations model with vector autoregressive disturbances. Journal of Econometrics, 85(1):51–74, July 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000948.

Tsurumi:1991:MSE

[TW91] Hiroki Tsurumi and Hajime Wago. Mean squared errors of forecast for selecting nonnested linear models and comparison with other criteria. Journal of Econometrics, 48(1–2):215–240, April/ May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190039G.

Toda:1995:SIV

[TY95] Hiro Y. Toda and Taku Yamamoto. Statistical inference in vector autoregressions with possibly integrated processes. Journal of Econometrics, 66(1–2):225–250, March/April 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694016168.

Tybout:1992:MND

[Tyb92] James R. Tybout. Making noisy data sing: Estimating production technologies in developing countries. Journal of Econometrics, 53(1–3):25–44, July/September 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407692900786.

Tauchen:1996:VVL

[TZL96] George Tauchen, Harold Zhang, and Ming Liu. Volume, volatility, and leverage: A dynamic analysis. Journal of Econometrics, 74(1):177–208, September 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017550.

Urbain:1995:PVF

[Urb95] Jean-Pierre Urbain. Partial versus full system modelling of cointegrated systems an empirical illustration. Journal of Econometrics, 69(1):177–210, September 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401668P.

Ullah:1994:MRQ

[US94] Aman Ullah and Virendra K. Srivastava. Moments of the ratio of quadratic forms in non-normal variables with econometric examples. Journal of Econometrics, 62(2):129–141, June 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900191.

VanDijk:1992:ICE

[Van92] Herman K. Van Dijk. International conference on econometric inference using simulation techniques. Journal of Econometrics, 51(1–2):287, January/February 1992. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769290040X.

Varian:1990:GFO

[Var90] Hal R. Varian. Goodness-of-fit in optimizing models. Journal of Econometrics, 46(1–2):125–140, October/November 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090051T.

VanPraag:1990:PCW

[VB90] Bernard M. S. Van Praag and Michael R. Baye. The poverty concept when prices are income-dependent. Journal of Econometrics, 43(1–2):153–166, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690901138.

vandenBerg:1997:AMD

[vdB97] Gerard J. van den Berg. Association measures for durations in bivariate hazard rate models. Journal of Econometrics, 79(2):221– 245, August 1997. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407697000213.

vandenBroeck:1994:SFM

[vdBKOS94] Julien van den Broeck, Gary Koop, Jacek Osiewalski, and Mark F. J. Steel. Stochastic frontier models: A Bayesian perspective. Journal of Econometrics, 61(2):273–303, April 1994. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407694900876.

vanderLeeuw:1994:CMA

[vdL94] Jan van der Leeuw. The covariance matrix of ARMA errors in closed form. Journal of Econometrics, 63(2):397–405, August 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900329.

Vahid:1997:CC

[VE97] Farshid Vahid and Robert F. Engle. Codependent cycles. Journal of Econometrics, 80(2):199–221, October 1997. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000328.

Velasco:1999:NSL

[Vel99] Carlos Velasco. Non-stationary log-periodogram regression. Journal of Econometrics, 91(2):325–371, August 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000803.

Vijverberg:1993:MUP

[Vij93] Wim P. M. Vijverberg. Measuring the unidentified parameter of the extended Roy model of selectivity. Journal of Econometrics, 57(1–3):69–89, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390059E.

Vijverberg:1997:MCE

[Vij97] Wim P. M. Vijverberg. Monte Carlo evaluation of multivariate normal probabilities. Journal of Econometrics, 76(1–2): 281–307, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407695017925.

Vinod:1995:DBS

[Vin95] H. D. Vinod. Double bootstrap for shrinkage estimators. Journal of Econometrics, 68(2):287–302, August 1995. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401639H.

Vinod:1998:FCF

[Vin98] H. D. Vinod. FELLOW’S CORNER foundations of statistical inference based on numerical roots of robust pivot functions. Journal of Econometrics, 86(2):387–396, October 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S030440769700119X.

VanSoest:1990:CIT

[VK90] Arthur Van Soest and Peter Kooreman. Coherency of the indirect translog demand system with binding nonnegativity constraints. Journal of Econometrics, 44(3):391–400, June 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407690900663.

Verbeek:1993:MME

[VN93] Marno Verbeek and Theo Nijman. Minimum MSE estimation of a regression model with fixed effects from a series of crosssections. Journal of Econometrics, 59(1–2):125–136, September 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407693900424.

Vogelsang:1999:SNP

[Vog99] Timothy J. Vogelsang. Sources of nonmonotonic power when testing for a shift in mean of a dynamic time series. Journal of Econometrics, 88(2):283–299, February 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000347.

vanPraag:1991:OCU

[vP91] Bernard M. S. van Praag. Ordinal and cardinal utility: An integration of the two dimensions of the welfare concept. Journal of Econometrics, 50(1–2):69–89, October 11, 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190090Z.

vanSoest:1993:CRD

[vSKK93] Arthur van Soest, Arie Kapteyn, and Peter Kooreman. Coherency and regularity of demand systems with equality and inequality constraints. Journal of Econometrics, 57(1–3):161–188, May/June 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390063B.

Vella:1999:TSE

[VV99] Francis Vella and Marno Verbeek. Two-step estimation of panel data models with censored endogenous variables and selection bias. Journal of Econometrics, 90(2):239–263, June 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000438.

Vuong:1993:MCS

[VW93] Quang H. Vuong and Weiren Wang. Minimum chi-square estimation and tests for model selection. Journal of Econometrics, 56(1–2):141–168, March 1993. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769390104D.

Whang:1993:TSP

[WA93] Yoon-Jae Whang and Donald W. K. Andrews. Tests of specification for parametric and semiparametric models. Journal of Econometrics, 57(1–3):277–318, May/June 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390068G.

Wang:1998:ECL

[Wan98] Liqun Wang. Estimation of censored linear errors-in-variables models. Journal of Econometrics, 84(2):383–400, June 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000936.

West:1995:PAS

[WC95] Kenneth D. West and Dongchul Cho. The predictive ability of several models of exchange rate volatility. Journal of Econometrics, 69(2):367–391, October 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401654I.

Wegge:1996:LIF

[Weg96] Leon L. Wegge. Local identifiability of the factor analysis and measurement error model parameter. Journal of Econometrics, 70(2):351–382, February 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407694016763.

Weiss:1990:LAE

[Wei90] Andrew A. Weiss. Least absolute error estimation in the presence of serial correlation. Journal of Econometrics, 44(1–2):127–158, April/May 1990. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407690900766.

Weiss:1991:MSE

[Wei91] Andrew A. Weiss. Multi-step estimation and forecasting in dynamic models. Journal of Econometrics, 48(1–2):135–149, April/ May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190035C.

Weiss:1993:SAM

[Wei93] Andrew A. Weiss. Some aspects of measurement error in a censored regression model. Journal of Econometrics, 56(1–2):169– 188, March 1993. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440769390105E.

West:1996:ISS

[Wes96] Mike West. Inference in successive sampling discovery models. Journal of Econometrics, 75(1):217–238, November 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017771.

West:1997:AHA

[Wes97] Kenneth D. West. Another heteroskedasticity- and autocorrelationconsistent covariance matrix estimator. Journal of Econometrics, 76(1–2):171–191, January/February 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017887.

Wickens:1996:ICV

[Wic96] Michael R. Wickens. Interpreting cointegrating vectors and common stochastic trends. Journal of Econometrics, 74(2):255–271, October 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017461.

Wilfling:1996:LOG

[Wil96] Bernd Wilfling. Lorenz ordering of generalized beta-II income distributions. Journal of Econometrics, 71(1–2):381–388, March/ April 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694017107.

Whang:1999:ADN

[WL99] Yoon-Jae Whang and Oliver Linton. The asymptotic distribution of nonparametric estimates of the Lyapunov exponent for stochastic time series. Journal of Econometrics, 91(1):1–42, July 1999. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407698000475.

Wolff:1990:MIE

[Wol90] Edward N. Wolff. Methodological issues in the estimation of the size distribution of household wealth. Journal of Econometrics, 43(1–2):179–195, January/February 1990. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090115A.

Wooldridge:1990:EAC

[Woo90] Jeffrey M. Wooldridge. An encompassing approach to conditional mean tests with applications to testing nonnested hypotheses. Journal of Econometrics, 45(3):331–350, ???? 1990. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769090003C.

Wooldridge:1991:ARR

[Woo91a] Jeffrey M. Wooldridge. On the application of robust, regressionbased diagnostics to models of conditional means and conditional variances. Journal of Econometrics, 47(1):5–46, January 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440769190076P.

Wooldridge:1991:STQ

[Woo91b] Jeffrey M. Wooldridge. Specification testing and quasi-maximum-likelihood estimation. Journal of Econometrics, 48(1–2):29–55, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407691900318.

Wooldridge:1995:SCP

[Woo95] Jeffrey M. Wooldridge. Selection corrections for panel data models under conditional mean independence assumptions. Journal of Econometrics, 68(1):115–132, July 1995. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769401645G.

Wooldridge:1996:ESE

[Woo96] Jeffrey M. Wooldridge. Estimating systems of equations with different instruments for different equations. Journal of Econometrics, 74(2):387–405, October 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017623.

Wooldridge:1999:DFE

[Woo99] Jeffrey M. Wooldridge. Distribution-free estimation of some nonlinear panel data models. Journal of Econometrics, 90(1):77– 97, May 1999. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000335.

Waterman:1996:EVI

[WW96] David Waterman and Andrew A. Weiss. The effects of vertical integration between cable television systems and pay cable networks. Journal of Econometrics, 72(1–2):357–395, May/June 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694017263.

Wansbeek:1998:MEE

[WW98] Tom Wansbeek and Michel Wedel. Marketing and econometrics: Editors’ introduction. Journal of Econometrics, 89(1–2):1–14, November 26, 1998. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407698000529.

Wang:1997:ECB

[WWC97] C. Y. Wang, Suojin Wang, and R. J. Carroll. Estimation in choice-based sampling with measurement error and bootstrap analysis. Journal of Econometrics, 77(1):65–86, March 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407696018064.

Wei:1997:APC

[WY97] Quanling Wei and Gang Yu. Analyzing properties of K-cones in the generalized data envelopment analysis model. Journal of Econometrics, 80(1):63–84, September 1997. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697000031.

Xiao:1998:HOA

[XP98] Zhijie Xiao and Peter C. B. Phillips. Higher-order approximations for frequency domain time series regression. Journal of Econometrics, 86(2):297–336, October 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697001188.

Xu:2010:NPC

[XTZ10] Jianjun Xu, Xianming Tan, and Runchu Zhang. A note on Phillips (1991): “A constrained maximum likelihood approach to estimating switching regressions”. Journal of Econometrics, 154(1):35– 41, January 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001481. See [Phi91].

Yang:1998:SEC

[Yan98] Minxian Yang. System estimators of cointegrating matrix in absence of normalising information. Journal of Econometrics, 85(2):317–337, August 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407697001048.

Yikang:1998:LPF

[Yik98] Li Yikang. Low-pass filtered least squares estimators of cointegrating vectors. Journal of Econometrics, 85(2):289–316, August 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697001036.

Yitzhaki:1994:EDO

[Yit94] Shlomo Yitzhaki. Economic distance and overlapping of distributions. Journal of Econometrics, 61(1):147–159, March 1994. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407694900817.

Young:1996:PBF

[YP96] Karen D. S. Young and Lawrence I. Pettit. On priors and Bayes factors. Journal of Econometrics, 75(1):113–119, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017720.

Yamada:1998:IPI

[YT98] Hiroshi Yamada and Hiro Y. Toda. Inference in possibly integrated vector autoregressive models: some finite sample evidence. Journal of Econometrics, 86(1):55–95, September 1998. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407697001097.

Zabel:1993:CNT

[Zab93] Jeffrey E. Zabel. A comparison of nonnested tests for misspecified models using the method of approximate slopes. Journal of Econometrics, 57(1–3):205–232, May/June 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390065D.

Zellner:1991:TDJ

[Zel91] Arnold Zellner. Tribute to Dennis j.aigner. Journal of Econometrics, 50(3):231, December 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407691900189.

Zellner:1993:DSB

[Zel93] Arnold Zellner. Discussion: Seasonal BVAR models. Journal of Econometrics, 55(1–2):231–234, January/February 1993. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769390013U.

Zellner:1996:MPI

[Zel96] Arnold Zellner. Models, prior information, and Bayesian analysis. Journal of Econometrics, 75(1):51–68, November 1996. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407695017682.

Zellner:1998:FSP

[Zel98] Arnold Zellner. The finite sample properties of simultaneous equations’ estimates and estimators Bayesian and non-Bayesian approaches. Journal of Econometrics, 83(1–2):185–212, March/April 1998. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407697000699.

Zha:1999:BRS

[Zha99] Tao Zha. Block recursion and structural vector autoregressions. Journal of Econometrics, 90(2):291–316, June 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000451.

Zheng:1996:CTF

[Zhe96] John Xu Zheng. A consistent test of functional form via nonparametric estimation techniques. Journal of Econometrics, 75(2):263– 289, December 1996. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407695017607.

Zellner:1991:FTP

[ZHkM91] Arnold Zellner, Chansik Hong, and Chung ki Min. Forecasting turning points in international output growth rates using Bayesian exponentially weighted autoregression, time-varying parameter, and pooling techniques. Journal of Econometrics, 49 (1–2):275–304, July/August 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407691900167.

Zellner:1999:FTP

[ZkM99] Arnold Zellner and Chung ki Min. Forecasting turning points in countries’ output growth rates: A response to Milton Friedman. Journal of Econometrics, 88(2):203–206, February 1999. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407698000177.

Zinde-Walsh:1991:ELR

[ZWG91] Victoria Zinde-Walsh and John W. Galbraith. Estimation of a linear regression model with stationary ARMA(p, q) errors. Journal of Econometrics, 47(2–3):333–357, February 3, 1991. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440769190106N.